A Complete Bibliography of the Journal of Number Theory (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

24 January 2024 Version 1.01

Title word cross-reference

(2, 2, 2) [33]. (2, 4) [651]. (2, 5) [442]. (3X + 1) [568]. (a mod n) [615]. (nα) [640]. (q; q)_{n} [1176]. (|ℑ(z^{n})|) [1220]. {−1, −1} [859]. −3 [121]. 0 [134, 212]. 1 [380, 444, 517, 709, 776, 834]. 12 [206]. 1 ≤ s ≤ q − 2 [598]. 2 [30, 41, 142, 195, 248, 274, 295, 338, 484, 547, 585, 590, 605, 657, 841, 850, 966, 970, 973, 984, 1086, 1094, 1168, 1172]. 2^{l} [410]. 2^{N} [590]. 2^{n} + 2 [326]. 2n [400]. 3 [338, 858, 888, 1168, 1172]. 3u^{2} − 2 = v^{6} [753]. 3x + 1 [635]. 4 [627, 774]. 4/n = 1/x + 1/y + 1/z [396]. 4p [1178]. 5 [589, 1130, 1142, 1200]. 6 [1237]. 7 [589, 984]. 8 [1177]. 9 [206]. ?(x) [1097]. [n^{c}] [521]. [nα + γ] [319]. ^{2} [39]. 2F_{1}(13, 23; 12; x) [1014]. A [781]. a^{x} + b^{y} = c^{z} [339]. a^{x} = R(x, y) [433]. α_{κ} [969]. andaquestionofErdős[278]. a p [377]. a p^{x} + bq^{y} = c + dp^{z}q^{w} [51]. ax + by = 1 [1095]. aY^{2l} = bX^{2l} + cZ^{2l} [1203]. aY^{l} = bX^{l} + cZ^{l} [1203]. B_{2k} [455, 488, 655]. B_{3} [727]. B_{h}[g] [656, 674]. β_{κ} [969]. \binom{2n}{n} [167]. mod [289]. mod1 [916]. mod3 [1161]. mod5 [506]. modp [394, 699]. C_{p} [998]. D_{1}x^{2} + D_{2} = 4p^{n} [829]. D_{1}x^{2} − D_{2}y^{2} = λk^{z} [649]. Δ [154]. Δ(x)

[477]. e [611, 950, 991]. E(T) [477]. E_{8} [438]. ℓ [932]. f(n, x) = 0 [1057]. f(x, y) [1189]. f(x, y) = f(u, v) [1140]. F_{2^{h}}[T] [549]. F_{p}[T] [761]. F_{q} [1179]. F_{q}(T) [845]. F_{q}(x) [598]. F_{q}[T] [38, 242, 803, 806, 951, 1044, 1180]. F_{q}[x] [13]. F_{r}[U] [1107]. G [778]. G_{2} [1076]. G_{K} [389, 745]. Γ [234]. Γ(2) [530]. ≥ 1 [1221]. ≥ 3 [1040]. H [498, 994]. H(d, q) [417]. j [67, 857]. J(q) [949]. J_{0}(p) [393]. J_{k}(n) [19]. K [336, 348, 491, 636, 658, 708, 734, 783, 931, 981, 1162, 1211, 1247]. k((t)) [258, 867]. K3 [1049]. K_{2} [272, 1177]. K_{2}(O_{F}) [30]. K_{2}O_{F} [405]. κ [969]. L [39, 111, 255, 256, 283, 460, 539, 569, 710, 740, 781, 797, 845, 865, 869, 888, 932, 973, 977, 1008, 1085, 1150, 1174, 1203, 1206]. l = 3, 5 [1226]. L^{2}(R^{n}) [698, 1182]. l^{n} [1226]. λ [234]. ⌊nα + s⌋ [508]. M [184, 664]. F_{2} [227]. Q [1058]. Cl_{2}(k^{1}) [981]. GL(2) [1046]. GL(3) [209]. GL(n) [996]. GL_{2} [634, 978]. GL_{2}(F_{3}) [162]. GL_{n} [503, 1144]. Sp(2n, Q(i)) [556]. n! [892]. n [554, 615, 794, 816, 1122]. nα + γ [915]. nα + s [508]. Ω [90]. P [24, 29, 53, 66, 93, 103, 134, 135, 172, 178, 226, 240, 250, 351, 357, 398, 436, 564, 569, 601, 650, 675, 676, 696, 703, 718, 721, 732, 751, 764, 781, 783, 793, 817, 872, 880, 882, 897, 898, 963, 978, 987, 1008, 1060, 1095, 1127, 1143, 1146, 1150, 1181, 1189, 1191, 1216, 1251, 1254]. p(n) [528, 1029]. P (t^{m}, Y) [276]. P (x) [477]. P (x) = [278]. p − 1 [701]. P^{1} [759]. p^{2} [963]. p^{3} [52]. P^{n} [166]. p^{x} − b^{y} = c [339]. Φ_{T} (X, Y) [559]. π [91, 300]. π_{1} [1009]. pq [524]. Q [8, 182, 349, 469, 484, 657, 711, 762, 897, 898, 912, 949, 1126, 1132, 1151, 1152, 1176, 1181]. Q(p, q, r) [993]. Q(ζ_{p}) [50]. Q ∗ (modp) [701]. Q_{ab}(t) [437]. r [1145]. S [58, 839, 858, 1050]. s(N) [517]. s = 0 [704]. S_{4} [711, 1151]. ∑(1/(q^{n} + r)) [120]. ∑_{i=1}^{N} 1/x_{i} [756]. ∑_{j=1}^{s}(1/x_{j}) + 1/(x_{1} ⋯ x_{s}) = 1 [327]. ∑_{p≤n} 1(p log p) [163]. ∑ t^{n}Aα^{n} + Bβ^{n} [1092]. T_{2} [695]. θ [333]. Ã_{n} [516]. φ [779, 869, 931]. φ_{k} [931]. φ_{k}(n) [931]. X^{2} + 1 = 2Y^{4} [107, 445]. X^{2} + 1 = dY^{4} [828]. x^{2} + D = 4p^{n} [233]. x^{2} + D = p^{z} [1229]. X^{2} + Y^{2} + Z^{2} = aXY Z [43]. x^{2} − D = 4p^{n} [252]. x^{2} − Dy^{2} = nz^{2} [199]. x^{4} + mx^{2}y^{2} + y^{4} = z^{2} [522]. x^{n} + y^{m} + 1 = 0 [1064]. X_{1}(m) [1003]. x_{1}x_{2} ≡ x_{3}x_{4} [766]. x_{1}x_{2} ≡ x_{3}x_{4}(modp) [766]. xyz = x + y + z [216]. y^{2} = x^{3} − ax [32]. y^{2} = x^{5} + 1/4 [1200]. Y^{2} = x^{q} − 1 [896]. Y^{m} = f(X) [597]. Z [439, 715, 1155]. Z_{p} [1007, 1150]. Z_{p}^{r} [134]. ζ [898, 1143]. ζ(2k + 1) [1227]. ζ(k) [373]. ζ(s) [622]. ζ(s)/Π^{s} [598]. ζ′(s) [622]. |η(z)| [209].

-additive [778, 1132]. -Adic [29, 53, 66, 93, 134, 135, 172, 178, 240, 250, 357, 436, 564, 569, 601, 650, 675, 696, 703, 751, 764, 793, 817, 872, 888, 898, 912, 963, 973, 978, 987, 1008, 1086, 1127, 1143, 1146, 1150, 1181, 1251]. -adique [817, 912, 1254]. -adiques [696]. -admissibility [783]. -Analogue [762, 1181]. -analogues [182]. -bases [498]. -Bernoulli [182, 1126].

-body [932]. -calculation [469]. -class [585, 627, 1094, 1172]. -class-group [484]. -Component [590]. -cores [774]. -curves [1058]. -cyclic [794]. -decompositions [498]. -Descent [547]. -dimensional [657, 1122]. -estimates [90]. -extension [134]. -extensions [516, 711, 721, 932, 1007, 1150]. -fold [931]. -free [336, 636]. -function [1155]. -Functions [111, 255, 256, 283, 333, 460, 539, 569, 710, 740, 797, 845, 865, 869, 888, 898, 973, 977, 1008, 1143, 1150, 1174, 1206]. -Funktion [24]. -group [1060]. -groups [195, 783]. -harmonic [781]. -ième [882]. -invariant [67]. -invariants [234, 857]. -lattices [715]. -linear [1179]. -modules [779]. -order [184, 664]. -periodic [912]. -périodiques [912]. -power [966]. -powers [295]. -prim [30]. -problem [568]. -rank [676, 850, 1145, 1152]. -Ranks [1177, 1216]. -rational [932]. -rationnels [932]. -regular [491]. -Sequences [455, 488, 674, 727]. -series [8, 349, 781]. -sheaves [869]. -sums [1211]. -th [658, 882, 994, 1162, 1247]. -th-powers [348]. -theory [734]. -torsion [1200, 1237]. -transforms [234]. -type [1049]. -unit [839]. -unités [1050]. -units [58, 398, 858, 1050]. -value [1085].

144 [1251].

216 [1253].

45 [768]. 450 [1252].

56 [1255]. 58 [768].

73 [1252].

A. [777]. abbreviated [1117]. ABC [1123]. Abelian [4, 30, 33, 72, 85, 87, 112, 176, 195, 200, 247, 326, 407, 457, 523, 561, 596, 678, 725, 775, 818, 887, 905, 906, 917, 926, 1094, 1106, 1150, 1252]. abéliennes [852, 917]. abéliens [1117]. absolument [852]. absolute [1075, 1191]. achievable [917]. Addendum [90, 105, 237, 292, 822, 918]. additif [952]. Addition [13, 456, 595, 596, 668, 723, 907, 918]. Additive [3, 60, 115, 174, 205, 280, 348, 370, 410, 539, 541, 588, 625, 663, 665, 682, 754, 778, 866, 912, 987, 1054, 1101, 1132, 1247]. Additively [429]. adelic [548]. Adic [29, 53, 66, 93, 134, 135, 172, 178, 240, 250, 357, 436, 564, 569, 601, 650, 675, 696, 703, 751, 764, 793, 817, 872, 888, 898, 912, 963, 973, 978, 987, 1008, 1086, 1127, 1143, 1146, 1150, 1181, 1251]. adique [817, 912, 1254]. adiques [696]. adjoint [1085, 1223]. admissibility [783]. Advances [722]. affine [692, 767, 869, 1236]. Airy [937]. al [1065]. aleatoric [414]. algebra [246, 1119]. Algebraic [15, 23, 46, 68, 117, 202, 210, 212, 222, 227, 272, 279, 297, 345, 368, 399, 402, 415, 419, 470, 494, 501, 502, 519, 529, 563, 594, 602, 609, 698, 705, 706, 739, 776, 786, 817, 847, 897, 909, 930, 953, 976, 1012, 1048, 1076, 1079, 1089, 1154, 1182, 1217, 1246]. algebraic-integers [609]. Algebraicity [781]. algebras [135, 308, 661, 1174, 1185, 1188]. algébrique [279, 909, 953]. algébriques [279, 817, 1254]. Algorithm [310, 485, 716, 768, 1115, 1212]. Algorithmic [1160]. algorithmique [1160]. Algorithms [113, 328]. Almost [179, 707, 1036, 1193]. alphabet [188]. Alternate [1180]. alternating [498]. amazing [774]. ambiguous [1120]. among [402]. analog [209, 570]. Analogue [762, 908, 1181]. analogues [182, 950]. analysis [221, 414, 426, 476]. Analytic [333, 364, 793, 1107]. analytical [909]. analyticity [464]. analytiques [909]. Anderson [934, 983]. Ankeny [449]. Apéry [250]. Application [8, 505, 691, 952, 1038, 1168]. Applications [68, 266, 324, 367, 473, 542, 566, 573, 771, 880, 996, 1001]. Approach [588, 909, 1160, 1180]. Approche [909, 1160]. approximable [31, 808, 1236]. approximants [1128]. approximate [382]. Approximation [9, 40, 70, 117, 305, 397, 423, 446, 462, 479, 508, 610, 696, 738, 808, 810, 835, 930, 949, 991, 1047, 1246]. approximations [114, 169, 300, 311, 374, 988]. April [412, 543, 700, 871, 1013, 1148]. arbitrary [995]. arguments [789]. arising [1047]. Arithmetic [126, 207, 271, 280, 342, 365, 389, 402, 404, 406, 584, 599, 728, 741, 745, 747, 798, 923, 946, 958, 1002, 1015, 1028, 1088, 1129, 1144, 1215]. Arithmetical [90, 269, 385]. arithmetically [56]. arithmetics [370]. arithmétique [1002]. arithmétiques [207, 1015]. arithmetische [741]. article [1252, 1253]. Artin [164, 172, 363, 408, 449, 1145]. aspects [872]. assertion [138]. Associate [288]. Associated [38, 134, 385, 419, 444, 487, 548, 712, 741, 840, 857, 977, 1193, 1221]. association [417]. associées [1221]. assoziierte [741]. assumes [232]. Astérisque [287]. asymmetric [397, 724, 933]. Asymptotic [91, 100, 141, 210, 228, 373, 730, 738, 812, 826, 922, 1023, 1103, 1162]. Atkin [151, 214]. attached [354, 402, 545, 858]. Auflösung [936]. August [451, 593, 755, 927, 1045, 1196]. Author [35, 63, 96, 128, 158, 191, 223, 261, 293, 322, 360, 386, 411, 440, 471, 499, 532, 552, 579, 603, 632, 653, 685, 713, 743, 769, 795, 823, 855, 889, 910, 942, 964, 985, 1004, 1021, 1034, 1055, 1080, 1112, 1135, 1158, 1183, 1208, 1250]. autodécrite [944]. Automata [719]. Automata-style [719]. Automorphic [537, 797, 1174, 1248]. Automorphism [486, 672, 814]. Automorphismengruppe [672]. Automorphisms [6, 258, 591, 867, 1064, 1071, 1114]. Average [364, 680]. Averages [1199]. axiomatization [473]. ayant [1237].

Badly [31, 1236]. Baker [313]. Banal [436]. Barrucand [378]. Base [623, 885, 996, 1110, 1232]. based [1014]. Bases [7, 20, 50, 85, 100, 147, 177, 217, 218, 228, 236, 253, 257, 288, 328, 407, 498, 538, 619, 698, 720, 794, 818, 973, 1011, 1167, 1182]. basis [606, 1019, 1054, 1212]. be [1012]. beam [7]. behaviour [812]. Bell [145]. Berechnung [7]. Bernoulli [49, 182, 566, 600, 764, 772, 1086, 1126, 1173, 1224]. Besicovitch [1039]. Bessel [978]. best [479]. between [34, 49, 271, 336, 554, 636, 817, 830, 843, 879, 1030, 1105]. Bianchi [219, 610]. bicyclic [585, 586, 1187]. Bielliptic [1170]. binary [138, 343, 365, 489, 548, 849, 938, 1012, 1099, 1123]. binomial [62, 226, 296, 573, 730, 884, 1051]. biquadratic [132, 133, 291, 510, 558, 585, 586, 1187]. biquadratiques [291]. Birch [1161]. Bloch [710]. blocks [1217]. Board [2, 37, 65, 98, 130, 160, 193, 225, 263]. bodies [153, 291, 752, 785]. body [672, 932, 1190]. Bombieri [266, 642]. Borel [1223]. Bornes [1050]. Boston [721]. bound [164, 266, 300, 597, 928, 957]. boundary [371, 431]. bounded [31, 61, 316, 345]. boundedness [465]. Bounds [198, 409, 959, 1050, 1141, 1172]. branched [153, 752]. Brauer [935]. Bruhat [661]. Brun [679, 1235]. Bugeaud [1254]. buildings [661]. Burnside [961].

C. [388]. Calculating [1029]. calculation [7, 469, 505, 548]. Calculations [862]. Calculs [862]. calculus [1179]. cannot [1012]. Canonical [81, 93, 464, 465, 874]. Cantor [84, 221, 716, 768]. capitulation [518]. caractères [903]. caractéristique [953]. Carlitz [76, 164, 183, 351, 789, 836, 886, 989, 1225]. Carmichael [1052]. carrés [1105]. cas [49]. case [49, 253, 363, 586, 634, 950, 1223]. cases [284, 712, 1149]. castling [1124]. Catalan [16, 421, 575]. Cauchy [302]. central [72, 518, 555, 661, 1174, 1185]. Certain [4, 66, 84, 112, 117, 122, 150, 171, 178, 207, 269, 282, 291, 295, 320, 505, 523, 545, 590, 687, 760, 765, 788, 862, 875, 876, 912, 1030, 1041, 1060, 1103, 1115, 1142, 1150, 1168, 1221, 1244]. certaines [207, 862, 912, 1168, 1221]. certains [291, 876]. chains [275, 1243]. Chalk [921]. Champernowne [1099]. change [885, 976, 996, 1110, 1232]. Changes [477]. Character [247, 399, 406, 417, 481, 605, 641, 766, 813, 896, 1127, 1199]. characteristic [134, 212, 351, 436, 612, 718, 732, 781, 815, 880, 881, 930, 953, 1010, 1030, 1095, 1187, 1207, 1245]. characteristics [338, 354]. characterization [525, 599]. characterizations [434, 587]. characterized [798]. Characterizing [486]. characters [142, 318, 903, 923]. Chowla [449, 893–895, 1065]. Circle [1106, 1176, 1252]. circles [312]. circulaires [1007]. circular [54, 218, 497, 662, 818, 887, 1007, 1164]. Class [4, 7, 10, 18, 34, 92, 94, 95, 102, 122, 153, 161, 197, 277, 286, 290, 295, 305, 329, 340, 347, 370, 378, 380, 401, 449, 453, 458, 484, 510, 520, 558, 585, 621, 627, 631, 659, 660, 676, 709, 773, 841, 842, 905, 906, 924, 959, 961, 966, 993, 1003, 1017, 1020, 1023, 1031, 1066, 1075, 1094, 1102, 1108, 1130, 1133, 1172, 1186, 1214, 1216, 1234, 1241, 1244, 1255]. Classes [153, 248, 385, 389, 525, 546, 590, 684, 711, 917, 1018, 1028, 1033, 1083, 1121, 1132, 1160, 1169, 1242]. classical [5, 570]. classification [338]. classified [1036]. close [776]. Closed [998]. closure [683, 811]. CM [157, 493, 659, 676, 896, 966, 1241]. CM-fields [676, 966, 1241]. CM-types [493]. Coates [618]. coatings [1190]. cocycle [1143]. cocycles [1143]. Codages [1053]. Codecà [90]. codes [164, 406, 644, 746, 1074]. codimension [1040]. coding [266, 1053]. coefficient [296]. coefficients [62, 89, 141, 226, 229, 408, 528, 564, 875, 884, 948, 960, 1073, 1204]. coendlicher [805]. Cogalois [577]. Cohen [1000]. Cohn [378, 1198]. cohomologous [1143]. cohomology [219, 689, 1033]. coincident [805]. colored [664]. colors [184]. combinations [598]. combinatorial [301, 628, 725]. comme [1117]. common [447, 802]. commutantes [870]. commutative [404, 463, 923]. compact [53]. comparaison [757]. comparison [757]. Complements [1202]. Complete [25, 155, 595, 645, 828, 869]. completion [205, 348, 541, 1247]. Completions [515, 1107]. Complex [470, 574, 627, 641, 670, 813, 903]. complexe [903]. complexity [400, 974]. Component [590]. composita [161]. composite [945]. compositum [662]. Computation [88, 104, 286, 302, 405, 650, 704, 1020]. computations [520, 1089]. Computing [219, 418, 840]. concerning [21, 138, 343, 478, 606, 721]. condition [799]. conditional [220]. Conditions [448, 509, 1009]. conductor [11, 561, 983, 1108]. conductors [928]. confidence [1059]. congruence [34, 296, 443, 448, 509, 684, 766, 831]. Congruences [49, 145, 298, 379, 449, 524, 697, 737, 764, 866, 979, 1041, 1043, 1086, 1204, 1224]. congurences [244]. Conjecture [16, 21, 57, 114, 163, 169, 284, 311, 332, 341, 388, 408, 416, 578, 682, 710, 762, 885, 913, 921, 1065, 1123, 1145, 1149, 1161, 1219]. conjectures [99, 287, 666, 1200]. connected [845]. connections [435]. consecutive [44, 271, 827, 879, 994]. consequences [825]. considerations [884]. constant [940, 1012]. constants [739, 821]. Constraints [372]. construct [328]. Constructing [681, 1118]. Construction [177, 407, 425, 526, 557, 848, 1099, 1212, 1228]. contained [731]. Containing [1057]. continuants [31]. Continuation [1185]. Continued [27, 150, 175, 197, 213, 221, 242, 374, 397, 529, 581, 630, 716, 758, 760, 768, 835, 930, 950, 988, 1089, 1246]. continues [952, 1053, 1072, 1096]. Continuous [952, 1053, 1096, 1139]. contour [1129]. contraintes [1072]. contribution [60, 421]. Contributions [78, 763]. controlling [11]. convergence [988]. converse [363, 807, 851]. convolutions [90]. coordinates [1115]. Coquet [825]. cores [774]. corps [153, 291, 752, 785, 882, 932, 1190]. Correction [926, 1133]. correlation [762, 860]. correspondence [577, 1231]. corresponding [93]. Corrigendum [237, 1067, 1068, 1251, 1252, 1255]. Cosine [656].

Counting [58, 345, 708, 999, 1249]. courbe [1237]. Courbes [154, 711, 785, 899, 903, 970, 1168, 1221, 1237]. Covering [583, 903]. coverings [594, 767]. Critère [882]. criteria [565, 709, 841]. Criterion [69, 304, 496, 575, 847, 882, 1202]. critical [977]. cross [457]. Crossed [404, 787]. crystals [919]. cubes [569, 1109, 1122, 1205]. Cubic [25, 67, 88, 126, 363, 486, 507, 548, 614, 638, 699, 704, 814, 844, 850, 888, 924, 1109, 1221]. cubics [165]. cubiques [1221]. curve [23, 235, 305, 475, 591, 602, 750, 936, 1003, 1064, 1130, 1237]. Curves [67, 154, 164, 227, 338, 350, 364, 401, 539, 547, 551, 574, 594, 595, 641, 644, 659, 670, 681, 711, 740, 746, 759, 785, 813, 830, 833, 850, 857, 858, 883, 896, 899, 900, 903, 955, 968, 970, 976, 984, 1009, 1048, 1058, 1074, 1115, 1118, 1161, 1168, 1170, 1197, 1203, 1221, 1237]. cusp [487, 977]. cuspidal [1003]. cycles [291]. Cyclic [20, 236, 328, 422, 486, 585, 590, 644, 783, 794, 814, 966, 981, 1060, 1102]. Cyclotomic [92, 137, 208, 218, 286, 380, 398, 693, 780, 800, 880, 905, 906, 983, 1011, 1051, 1066, 1075, 1157, 1210, 1245]. cyclotomie [735]. cyclotomy [735].

d’addition [13]. dans [13, 49, 882, 1007]. d’approximation [949]. Davenport [644, 717]. decay [831, 1046]. December [490, 643, 809, 972, 1091, 1240]. decimal [117]. decompositions [179, 498]. Dedekind [89, 434, 669, 702, 929, 1061, 1157]. Defined [112, 1076, 1203, 1217]. defining [1190]. definite [423, 448, 509, 526, 848, 849, 864, 1040, 1178, 1193]. definition [887, 1190]. Deformation [1232]. Deformations [1024, 1223]. Deforming [11, 1063]. degeneracy [608]. degenerate [493]. degenerations [891]. Degert [709, 841]. Degree [15, 17, 41, 46, 200, 326, 410, 590, 792, 963, 987, 1079]. Delange [389, 745]. delay [371, 431]. Deligne [863]. Dem’janenko [18, 561, 775, 1168]. Dennis [511]. denominators [48, 706]. Dense [656, 834, 1169]. denses [834]. densité [876]. Densities [270, 545, 627, 860, 904, 1038, 1077]. Density [363, 606, 646, 656, 674, 876, 907]. d’entiers [147, 1105]. derivatives [1155]. Descent [547, 1190]. Descente [1190]. described [944]. describing [582]. designs [428]. detecting [253, 1082]. determinant [712, 775]. Determinantal [1003, 1157]. Determinants [4, 302, 652, 673, 1166]. determine [1115]. Deterministic [414]. deux [652, 953]. deviations [207]. diagonal [66, 282, 747]. Diagonalising [196]. Dickson [148, 482, 749]. d’idéaux [291, 852]. Différence [554, 1105]. differences [459, 1162]. different [124, 664]. Differential [371, 431, 1042]. differentials [900]. Digit [78, 929]. Digital [188, 573]. digits [31, 61, 798, 826, 938, 1028, 1121, 1217, 1242]. dihedral [17, 1167, 1241]. Dimension [221, 446, 601, 605, 716, 742, 768, 805, 1047, 1205]. dimensional [396, 531, 569, 657, 842, 853, 919, 1122]. d’indépendance [279]. d’interpolation [652, 1166]. diophantienne [696]. Diophantine [25, 40, 51, 70, 114, 115, 169, 171, 199, 233, 252, 265, 277, 311, 355, 374, 397, 445, 462, 466, 610, 645, 649, 696, 738, 739, 753, 828, 835, 930, 959, 1019, 1047, 1140, 1213, 1246]. DiPippo [1252]. Dirichlet [70, 111, 182, 255, 256, 469, 487, 592, 860]. discrepancy [267, 640]. discrete [1139]. discriminant [77, 444, 1131, 1178, 1204]. discriminants [815, 1221]. Discriminator [1090]. discriminators [749]. dispersion [497]. distance [817, 922, 1254]. distances [168]. distinct [619, 681, 1169]. distinguished [1243]. distributed [352]. Distribution [31, 39, 110, 124, 189, 226, 260, 312, 377, 384, 394, 478, 517, 570, 680, 727, 750, 778, 803, 816, 839, 912, 916, 922, 946, 1015, 1018, 1028, 1121, 1132, 1147, 1169, 1242]. distributions [217, 644, 978]. divides [447, 802]. diviseurs [690, 834, 1165]. Divisibility [295, 560, 799, 875]. Division [32, 135, 1119]. Divisor [198, 331, 615, 933, 1103]. Divisors [167, 356, 408, 519, 544, 690, 701, 834, 931, 1061, 1137, 1165]. d’Iwasawa [757]. domain [1061]. domains [434]. donnée [1237]. Drinfeld [274, 313, 354, 444, 559, 616, 647, 680, 857, 948, 980, 1010, 1073, 1131, 1133, 1156, 1197, 1204]. dual [576]. duality [1116]. duals [468]. d’un [1015]. duplication [756]. Dyadic [156, 366]. Dyer [151, 1161]. dynamical [868]. dynamics [84, 843]. Dynamique [1072].

E. [341]. easier [367]. editor [203, 901]. Editorial [2, 37, 65, 98, 130, 160, 193, 225, 263]. effectif [882]. effective [745, 817, 882, 1050, 1166, 1190, 1254]. effectives [1050]. Egyptian [48, 310, 356]. Ehrenfeucht [69]. Eigenvalues [219, 285, 695]. eighth [170]. einem [7]. Eisenstein [301, 318, 503, 1026]. Elasticity [82, 536]. element [125, 859]. élémentaire [1002]. elementary [30, 72, 325, 435, 648, 884, 1002, 1172]. elements [245, 297, 335, 381, 840, 917]. eleventh [1164]. Elliptic [67, 81, 154, 185, 338, 350, 364, 401, 464, 465, 475, 539, 551, 571, 574, 595, 641, 659, 670, 696, 711, 718, 813, 833, 850, 883, 900, 903, 908, 955, 1030, 1058, 1085, 1115, 1161, 1192, 1221]. elliptic-curves [641]. elliptic-surfaces [464]. elliptical [785, 899, 1237]. elliptique [1237]. elliptiques [154, 696, 711, 785, 899, 903, 1221, 1237]. elliptische [185]. embedding [437, 555]. embeddings [715, 1188]. endomorphism [53]. Ensembles [870, 876]. entier [1015]. Entiers [154, 690, 834, 1165]. Entireness [869]. entre [49, 817, 1105, 1254]. entropy [588]. enumeration [370]. Equal [62, 179, 338, 557, 884]. equality [73, 847, 1154]. Equation [43, 51, 107, 187, 199, 216, 233, 252, 274, 278, 382, 421, 445, 514, 522, 529, 575, 649, 753, 766, 828, 829, 948, 1025, 1095, 1140, 1229]. Equations [25, 61, 66, 132, 133, 171, 265, 277, 282, 316, 320, 339, 340, 355, 371, 431, 433, 474, 491, 586, 613, 645, 649, 691, 692, 792, 801, 828, 839, 854, 959, 982, 1050, 1163]. Equidistant [406]. equivalence [546, 587, 599, 1079]. equivalent [56, 468, 1216]. equivariant [1006]. Erdos [21, 163, 186, 287, 478, 639, 754, 851, 873, 1219, 1238]. ergodic [952, 1032]. ergodique [952]. Erratum [190, 238, 768, 837, 925, 941, 1134, 1253]. error [19, 194, 893–895, 1047, 1162]. errors [446]. estimates [90, 1023]. Estimating [850]. estimation [177]. étale [767]. Euclidean [434, 485, 620, 1100]. Euclidean-like [434]. Euler [127, 331, 693, 734, 786, 1224]. Evaluation [42, 469]. even [217, 526, 987, 1063, 1178]. Everett [1252]. everywhere [955]. Exact [141, 228, 1082]. exceptional [601]. exclusion [373]. Existence [721, 794, 900]. expansion [529, 974, 1131]. expansions [78, 240, 581, 760, 835, 930]. Explicit [7, 162, 389, 878, 1077, 1117, 1225]. explicite [1117]. expliziten [7]. exponent [235, 484]. Exponential [164, 171, 242, 460, 649, 758, 784, 854, 996, 1128]. exponents [275, 892, 1246]. Expressible [1175, 1253]. extending [1019]. Extension [28, 56, 134, 201, 326, 1084, 1231]. Extensions [15, 30, 72, 123, 200, 202, 258, 259, 407, 422, 484, 486, 501, 516, 518, 577, 590, 627, 665, 711, 721, 752, 773, 787, 794, 814, 833, 852, 897, 917, 920, 932, 971, 973, 1007, 1142, 1150, 1151, 1201]. Extra [518]. Extremal [41, 236, 788]. extreme [267].

facteur [690]. facteurs [1015]. factor [127, 286, 331, 690]. factorable [1062]. factorial [434]. factorials [1061]. Factorisatio [372]. factorization [82, 414, 892, 1234]. factorizations [297, 536]. factors [303, 335, 963, 1015]. facts [774]. Families [113, 320, 626, 638, 883, 970, 982, 1163, 1168]. familles [970, 1168]. family [25, 53, 80, 165, 364, 551, 581, 614, 650, 767, 828, 888, 947, 1036]. Farey [637, 922]. February [14, 106, 204, 307, 395, 512, 667, 838, 997, 1125]. Fermat [187, 235, 253, 591, 672, 786, 936, 945, 1049]. Fermatkörpers [672]. Fermatkurve [936]. few [168, 259]. Fibonacci [355]. fibre [79]. Field [7, 92, 112, 134, 155, 201, 202, 218, 222, 290, 407, 558, 577, 634, 704, 750, 751, 787, 842, 882, 887, 905, 906, 935, 950, 971, 978, 993, 1057, 1094, 1119, 1157, 1172, 1243]. field-extensions [407]. Fields [4, 12, 20, 30, 32, 33, 56, 66, 67, 72, 80, 88, 93, 104, 123, 126, 132, 133, 156, 187, 197, 216, 241, 248, 286, 306, 308, 318, 326, 329, 347, 358, 363, 366, 380, 402, 407, 409, 449, 482, 501, 515, 523, 527, 538, 586, 614, 626, 627, 662, 691, 693, 709, 734, 736, 747, 748, 765, 786, 791, 800, 812, 818, 841, 849, 859, 861, 865, 913, 924, 955, 981, 983, 995, 1020, 1031, 1037, 1064, 1066, 1079, 1084, 1094, 1102, 1106, 1108, 1129, 1147, 1150, 1167, 1172, 1179, 1187, 1191, 1203, 1235, 1239, 1244, 1248, 1252]. fields [6, 7, 34, 43, 68, 73, 76, 77, 82, 95, 102, 110, 118, 136, 146, 157, 177, 208, 210, 282, 328, 354, 381, 399, 401, 414, 453, 484, 494, 496, 502, 510, 561– 563, 585, 588, 616, 621, 631, 638, 676, 678, 687, 703, 729, 739, 773, 775, 794, 847, 848, 850, 863, 878, 888, 963, 966, 973, 1017, 1024, 1042, 1071, 1118, 1121, 1127, 1154, 1177, 1186, 1188, 1207, 1216, 1241]. Fine [456]. finis [1117]. Finite [6, 85, 155, 188, 195, 231, 236, 241, 257, 282, 334, 414, 457, 463, 482, 498, 596, 616, 725, 739, 747, 750, 787, 812, 853, 1037, 1076, 1099, 1106, 1117, 1118, 1144, 1147, 1203, 1252]. Finite-dimensional [853]. Finitely [15, 701, 1191]. Finiteness [1133]. first [286, 1026]. Fitting [1108]. five [235]. fixed [44, 1012, 1061]. Flächensingularitäten [936]. fold [931]. Folded [213]. fonctions [207, 909, 912, 1015]. fondamentale [291]. Fontaine [284, 1149]. form [132, 133, 202, 268, 521, 586, 613, 691, 1085]. Formal [111, 255, 256, 298, 357, 545, 735, 843]. formations [842]. formelles [735]. Formes [652, 1156]. Forms [8, 89, 103, 109, 121, 138, 156, 196, 201, 229, 260, 343, 359, 365, 380, 410, 415, 423, 424, 429, 442, 448, 460, 487, 489, 509, 526, 537, 540, 545, 546, 548, 549, 651, 652, 671, 691, 703, 746, 806, 807, 821, 848, 864, 872, 977, 980, 987, 1000, 1029, 1030, 1038, 1040, 1076, 1077, 1109, 1156, 1174, 1178, 1193, 1233, 1236, 1248]. Formula [29, 68, 91, 92, 140, 382, 528, 551, 676, 730, 844, 904, 1003, 1006, 1077, 1103, 1118, 1131, 1157, 1162, 1227, 1234]. formulae [738]. formulas [141, 308, 373, 389, 453, 745, 780, 1116]. four [1194]. Fourier [89, 793]. fourth [535]. fractal [264]. Fraction [175, 221, 242, 374, 529, 581, 716, 760, 768, 835, 846, 930, 988]. fractional [124, 330, 504, 513, 915]. fractions [27, 48, 117, 150, 197, 213, 310, 356, 397, 630, 758, 950, 952, 975, 1053, 1072, 1089, 1096, 1246]. free [303, 306, 336, 636, 956, 999]. Freiman [60]. French [13, 49, 99, 116, 125, 153, 154, 207, 279, 291, 652, 690, 696, 711, 735, 752, 757, 785, 817, 834, 862, 876, 882, 899, 903, 909, 912, 917, 932, 944, 949, 952, 953, 960, 962, 970, 1002, 1007, 1015, 1050, 1053, 1093, 1096, 1105, 1117, 1156, 1160, 1165, 1166, 1168, 1190, 1221, 1228, 1237]. Frobenius [21, 184, 466, 664, 697, 723, 1002, 1019]. Fröhlich [1006]. Fuchs [851]. Fuchsian [958]. Fueter [299]. fully [752]. Function [12, 19, 24, 76, 80, 112, 118, 146, 177, 183, 187, 197, 222, 232, 254, 318, 319, 358, 381–383, 449, 452, 453, 458, 479, 496, 523, 531, 535, 544, 548, 631, 693, 712, 729, 732, 736, 739, 789, 812, 826, 863, 865, 881, 886, 916, 922, 935, 950, 1017, 1024, 1031, 1044, 1047, 1064, 1066, 1097, 1121, 1131, 1155, 1157, 1164, 1179, 1187, 1204, 1207, 1235, 1248, 1249]. functional [221, 382, 491, 692]. Functions [42, 88, 111, 122, 141, 152, 174, 185, 198, 207, 210, 212, 243, 249, 250, 255, 256, 277, 280, 283, 333, 342, 375, 385, 389, 444, 460, 467, 473, 515, 534, 539, 563, 564, 569, 571, 682, 689, 710, 718, 728, 740, 745, 751, 778, 784, 793, 797, 831, 845, 865, 868, 869, 888, 898, 909, 912, 946, 957, 973, 977, 1001, 1008, 1015, 1023, 1038, 1046, 1048, 1088, 1101, 1107, 1124, 1128, 1129, 1132, 1139, 1143, 1146, 1150, 1174, 1181, 1197, 1206, 1242, 1251]. fundamental [23, 291, 409, 418, 507, 602, 638, 878]. Funktion [24]. Funktionen [185]. further [1099].

Galois [11, 17, 23, 126, 148, 162, 185, 195, 243, 245, 329, 354, 398, 422, 435, 484, 501, 571, 574, 594, 602, 626, 657, 665, 791, 836, 840, 861, 882, 897, 989, 1007, 1033, 1060, 1114, 1117, 1119, 1142, 1190, 1191, 1201, 1223]. galoisien [882]. galoisienne [1007]. galoisiens [1190]. Galoismodulstruktur [185]. gamma [650, 789, 886, 1181]. Ganzheitsbasen [7]. Gap [504]. gaps [271, 336, 636, 879]. Gauss [29, 118, 443, 644, 844, 862, 934, 990]. Gaussian [705, 826]. Genera [95, 546, 788]. General [645, 763, 933, 1039, 1047]. généralisation [116]. Generalization [69, 116, 333, 378, 405, 531, 568, 635, 637, 1116]. generalizations [480]. Generalized [184, 187, 250, 384, 514, 516, 581, 664, 688, 746, 829, 999, 1061, 1074, 1195, 1229, 1242]. generate [28]. generated [15, 391, 701, 1191]. Generates [1057]. generating [319, 751, 840]. generation [511]. generators [241]. genre [970, 1168]. Genus [126, 222, 858, 900, 923, 970, 976, 984, 1168]. geometric [164]. géométrique [757]. geometry [235, 584, 757]. German [7, 10, 24, 179, 185, 672, 741, 805, 936]. Gilmer [181]. Gittern [805]. Given [268, 1057, 1237]. gleichen [179]. global [73, 453, 465, 496, 527, 563, 1048, 1186, 1191]. going [113, 581]. Goldbach [179, 688, 940]. Goldbach-Zerlegungen [179]. Göllnitz [1001, 1016, 1153]. Golomb [194, 582, 944]. Good [267, 955]. Gordon [1001]. Goss [531, 789, 886, 1044]. Graham [21, 186]. grand [690]. Grandes [207]. Granville [777]. graphs [247]. gratings [805]. Green [1197]. Greenberg [913]. Gröbner [253]. Groessencharacters [957]. großer [10]. Group [18, 23, 112, 161, 172, 218, 367, 391, 401, 405, 430, 484, 585, 590, 591, 602, 659, 672, 805, 819, 905, 906, 935, 961, 1031, 1060, 1064, 1075, 1076, 1119, 1160, 1172, 1186]. groupe [1160]. groupes [279, 1117]. Groups [17, 46, 67, 73, 85, 94, 111, 162, 176, 195, 219, 231, 236, 255– 257, 279, 298, 347, 357, 380, 392, 401, 436, 457, 476, 498, 520, 527, 596, 610, 660, 668, 676, 725, 783, 791, 818, 843, 847, 850, 897, 913, 920, 926, 958, 966, 967, 1060, 1098, 1108, 1117, 1152, 1154, 1191, 1216, 1255]. Growth [1204].

H. [1198]. Haar [698, 1182]. half [22, 196, 396]. half-integral [22, 196]. Hamming [417, 746]. Hans [413]. Hardy [140, 528, 629, 1155]. Harmonic [476, 781, 790]. Hasse [80, 172, 289, 644, 1154, 1206]. Hausdorff [221, 446, 601, 716, 768, 1047]. hauteur [899]. Hauteurs [1228]. having [859, 1083, 1215, 1237]. Hecke [142, 143, 219, 246, 308, 318, 333, 530, 583, 695, 710, 957, 977, 1111]. height [81, 316, 345, 415, 464, 465, 868, 899, 1082]. heights [874, 1228]. Heisenberg [476]. Hensel [560]. hereditary [1185]. Hermite [1128].

Hermitian [200, 423, 526, 848, 849, 1038]. high [792, 1059]. Higher [333, 531, 569, 734, 842, 1079, 1110, 1155]. higher-dimensional [531, 569]. highly [1062]. Hilbert [126, 212, 290, 505, 510, 861, 1020, 1079, 1172, 1205, 1239]. Hirzebruch [143]. Hisashi [1253]. Hodge [18, 891]. holistic [7]. holomorphic [487, 537]. homogeneous [121, 424, 442, 651, 702]. homology [53]. Homomorphisms [401, 787]. Hooley [1129]. Howe [1252]. Hurwitz [49, 474, 676]. Huxley [1129]. hyper [1141]. hyper-Kloosterman [1141]. Hyperbolas [877]. hyperelliptic [80, 547, 857, 858, 903]. hyperelliptiques [903]. hypergeometric [250, 564, 1014]. Hypergraphs [771]. hyperplanes [1207]. hypothesis [732, 761, 939, 1044, 1202].

Ideal [94, 218, 253, 380, 520, 662, 676, 731, 905, 906, 966, 1186]. ideals [449, 748, 1108, 1110, 1120]. Idele [847]. Idempotent [402]. idempotents [961]. identities [253, 301, 458, 589, 1026]. identity [463, 702, 881, 1014, 1153]. ième [882]. Igusa [712, 740, 1124]. Ihara [607]. ii [31, 46, 114, 133, 157, 205, 256, 330, 340, 371, 456, 464, 476, 478, 509, 524, 602, 637, 763, 802, 878, 894, 906, 975, 989, 1063, 1066, 1079, 1099, 1143, 1145, 1149]. III [48, 103, 115, 169, 176, 431, 465, 586, 876, 895, 926, 971, 1032]. imaginär [7]. imaginär-quadratischen [7]. Imaginary [4, 7, 33, 43, 104, 110, 216, 329, 407, 484, 585, 773, 775, 818, 848, 849, 981, 1102, 1172, 1187]. imperfect [347]. improved [266]. improvement [382]. inclusion [373]. incomplete [1141]. incongruence [44]. indecomposable [109, 429, 526, 848]. indefinite [121, 424, 442, 651]. Indépendance [909, 953]. independence [178, 279, 358, 909, 953, 1217]. Index [35, 63, 92, 96, 128, 132, 133, 158, 191, 223, 261, 293, 322, 360, 386, 391, 411, 440, 471, 499, 532, 552, 579, 586, 603, 624, 632, 653, 685, 691, 713, 743, 769, 780, 795, 823, 855, 880, 889, 910, 942, 964, 985, 1004, 1006, 1021, 1034, 1055, 1080, 1112, 1135, 1158, 1183, 1208, 1250]. indices [33]. Indivisible [883]. inequalities [115, 578, 646, 860, 947, 1213]. inequality [249, 303, 478, 492, 736, 1087, 1128]. infinies [153]. Infinite [153, 240, 354, 430, 956, 992]. Infinitely [621]. infinitesimal [891]. infinity [113, 581]. inhomogeneous [462, 760, 835, 1039]. Integer [285, 314, 319, 446, 504, 513, 554, 601, 683, 706, 764, 914, 916, 1011, 1015, 1060, 1082, 1084, 1087, 1165, 1175, 1253]. integer-detecting [1082]. Integer-valued [314, 683, 914, 1084]. Integers [44, 61, 95, 200, 272, 299, 303, 422, 469, 557, 574, 609, 628, 690, 691, 708, 820, 849, 907, 963, 975, 1028, 1051, 1075, 1105, 1169, 1175, 1177, 1238, 1249, 1253]. Integral [9, 20, 22, 50, 67, 196, 232, 368, 407, 429, 474, 538, 546, 597, 794, 800, 864, 973, 1076, 1115, 1201, 1207]. integrals [650, 1181]. interesting [984]. interpolation [652, 1042, 1143, 1166]. intersection [155]. intersections [869]. interval [47]. intervals [383, 629, 707]. invariance [417]. Invariant [67, 80, 152, 175, 665, 963]. invariants [93, 104, 234, 402, 555, 647, 678, 752, 857, 1244]. inverse [390, 554, 742]. inverses [816]. irrational [40, 304, 479, 513, 1012]. Irrationality [108, 120, 171, 351, 565, 612, 1092, 1195]. irreducibility [69, 276, 1111]. Irreducible [482, 562, 644, 750, 803, 1147]. irregularity [189]. Isogenies [616, 843]. isogeny [1133]. isomorphic [681]. Iterate [931]. iteration [690]. IV [311]. Iwaniec [396]. Iwasawa [104, 134, 234, 539, 678, 757, 1244].

J [1255]. Jack [605]. Jacobi [113, 142, 166, 250, 344, 392, 419, 450, 463, 481, 564, 671, 807, 872, 1000, 1043, 1233]. Jacobian [235, 1200]. Jacobians [547, 681, 830]. January [1, 97, 192, 294, 387, 500, 654, 824, 986, 1113]. Jarnik [1039]. Jean [825]. joint [1132]. Jordan [19]. July [55, 149, 251, 346, 441, 580, 744, 911, 1035, 1184]. June [45, 139, 239, 337, 427, 567, 733, 902, 1027, 1171].

K. [777]. Kátai [682]. Kato [710]. kernel [511, 859]. key [1153]. kind [1086]. Klassenzahl [10]. Kleitman [332]. Kloosterman [550, 592, 937, 1098, 1141]. Kneser [324, 577]. Knopp [702]. Kodaira [1228]. Kolyvagin [693]. Kronecker [570, 587, 1214]. Kronecker-sequences [570]. Krull [434]. Kubilius [303, 1087]. Kubota [556]. Kummer [76, 146, 208, 259, 298, 496, 920, 963, 979, 1225]. Kummer-type [979].

L [388]. Laguerre [516]. lambda [104]. Lang [57]. Large [10, 87, 207, 271, 303, 340, 428, 628, 674, 690, 736, 866, 879, 969]. Largest [1175, 1253]. Lattice [312, 390, 624, 877]. Lattices [41, 95, 168, 425, 468, 576, 620, 715, 788, 819, 840, 849, 1078, 1104]. Laurent [1254]. laws [333, 595, 1225]. Least [101, 121, 447, 802, 970]. Lee [460]. Legendre [151, 1025]. Lehmer [91]. Lehner [214]. Lemke [332]. lemma [208, 560, 962]. lemme [962]. Length [48, 581, 609, 846, 1012]. lengths [113, 291, 297]. Letter [203, 901]. level [380, 1063, 1130]. Li [1202]. Lifted [53]. lifting [550, 885]. liftings [657, 891]. Light [1097]. like [434, 560]. Limit [29, 1227]. limits [751]. Lindemann [1166]. line [767]. linéaires [652, 1156]. Linear [34, 42, 61, 78, 108, 178, 229, 316, 355, 415, 466, 598, 652, 1019, 1109, 1156, 1179]. linked [641]. Littlewood [353, 629]. Ljunggren [107]. Local [134, 152, 308, 347, 464, 545, 562, 563, 634, 712, 752, 842, 904, 990, 1038, 1048, 1071, 1077, 1119, 1124, 1185, 1188, 1210, 1225, 1243]. local-field [634]. Locally [1107]. locaux [752]. Log [781, 1181]. Log-Algebraicity [781]. logarithm [415]. logarithmes [652, 696, 1156]. logarithmic [1160]. logarithmiques [1160]. logarithms [229, 652, 696, 1156]. longueurs [291]. look [780]. Lorentz [805]. Lorentz-Gittern [805]. low [974]. lower [300, 1172]. Lubin [1201]. Luca [52, 325]. Lucas [726, 1083].

M. [324]. même [1237]. Maass [283, 671]. magic [569]. magnitude [101]. Mahler [473, 491, 565, 611, 776, 853, 1050, 1128]. Mahler-type [491]. Maillet [775]. main [729]. manifolds [70, 738]. Manin [1168]. Many [15, 83, 227, 335, 621, 642, 746, 1074, 1118, 1137]. Many-Variable [642]. Maps [175, 608, 830]. March [26, 119, 215, 317, 403, 533, 686, 856, 1005, 1136]. Markoff [43, 116]. Markov [474, 967, 1036]. Matrices [9, 211, 229, 285, 385, 601, 673, 870, 1087]. matrix [18, 135, 561, 650, 775]. Maximal [642, 661, 773, 968, 1066, 1076]. maximale [805]. maximum [805]. May [36, 129, 224, 323, 420, 553, 714, 890, 1022, 1159]. Mazur [284, 1149]. McCulloh [1006]. Mean [283, 382, 383, 487, 534, 592, 669, 766, 797, 865, 1103, 1162]. mean-square [1162]. Mean-value [383]. measure [38, 178, 189, 611, 776, 853, 949]. measures [108, 175, 210, 351, 357, 573, 612, 765]. Memoriam [413]. Mesure [949]. metabelian [102]. metacyclic [917]. métacycliques [917]. method [174, 183, 473, 537, 684, 831, 1046, 1129, 1168]. méthode [1168]. methodical [909]. méthodique [909]. methods [39, 414]. metric [39, 810, 952]. Metrical [27, 330, 619, 765]. métrique [952]. Meuser [152]. mid [62]. mid-coefficients [62]. middle [342]. Mignosi [568]. Minimal [362, 576, 1054]. minimization [817, 1156]. Minimum [77, 656, 722]. Minkowski [102, 1097]. Minoration [817, 1254]. Minorations [1156]. Mirimanoff [244, 618]. missing [1028]. mixed [115]. mod [1226]. mod- [1226]. model [299]. modulaire [962]. Modular [89, 126, 196, 274, 350, 380, 381, 417, 559, 584, 620, 634, 718, 788, 821, 830, 861, 948, 962, 980, 1001, 1003, 1030, 1073, 1085, 1104, 1130, 1152, 1170, 1174, 1197]. module [134, 185, 329, 422, 571, 836, 983, 989, 1201]. modules [274, 313, 354, 444, 542, 616, 647, 680, 779, 836, 857, 989, 1010, 1073, 1133, 1156, 1239]. moduli [94, 206, 342, 589, 945, 1010]. Modulo [103, 166, 226, 296, 377, 478, 517, 524, 554, 816, 844, 990, 1029, 1189, 1210, 1211]. modulus [268, 866, 979]. moins [970]. moment [535]. Moments [174, 336]. monogeneity [299]. monoids [923]. Mordell [112, 659, 742, 819, 1221]. morphisms [362, 759, 1012]. Morse [264, 1096, 1099]. motives [1049]. Muller [1074]. multidimensional [42]. Multigrade [275]. multinomial [408, 564, 875]. multiple [79, 447, 723, 802, 918, 1116]. multiples [513, 646, 876]. Multiplication [574, 641, 670, 813, 903]. Multiplicative [122, 131, 249, 265, 342, 358, 394, 534, 665, 1088]. multiplicatively [625]. multiplier [572]. n [278]. Nagell [514, 829, 1229]. nahezu [179]. natural [956, 994]. nenner [24]. Neron [338]. Nesterenko [473]. newforms [22, 59]. Niemeier [425]. nilpotent [257]. No [62, 708]. nombre [690, 711, 1015]. nombres [49, 99, 817, 882, 932, 1254]. Non [121, 126, 404, 424, 442, 484, 537, 623, 626, 651, 859, 870, 917, 1012, 1146, 1206, 1241, 1251]. non-abelian [917]. non-commutantes [870]. non-commutative [404]. non-holomorphic [537]. non-homogeneous [121, 424, 442, 651]. non-normal [623, 1241]. Non-trivial [859, 1012, 1206]. Non-Urs [1251]. nonabelian [920]. Nonexistence [788, 955, 1142]. Nonholomorphic [392]. nonmodular [141]. Nonmonotonicity [1220]. nonnegative [1088]. Nontotients [309]. Nonvanishing [977]. Nörlund [1086]. Norm [73, 260, 289, 291, 505, 527, 642, 847, 857, 1154]. norm-forms [260]. Normal [28, 50, 241, 328, 623, 720, 973, 1212, 1241]. normales [147]. normality [619]. normalized [62, 1193]. normaux [870]. Norme [291]. normierter [24]. norms [419]. Note [23, 44, 54, 74, 161, 182, 254, 271, 393, 419, 514, 536, 568, 600, 602, 655, 737, 777, 829, 887, 898, 924, 1078, 1090, 1111, 1219, 1254]. notebooks [458]. Notes [1198]. nouveaux [752]. November [86, 180, 281, 376, 483, 633, 796, 965, 1081, 1230]. Number [4, 7, 18, 30, 33, 40, 60, 68, 82, 92, 95, 123, 127, 131– 133, 155, 156, 201, 202, 210, 286, 297, 306, 316, 326– 329, 340, 366, 399, 401, 402, 407, 435, 449, 453, 457, 485, 495, 501, 502, 551, 558, 586, 615, 621, 663, 678, 690, 691, 698, 709, 711, 715, 734, 748, 768, 775, 780, 786, 791, 818, 829, 841, 847, 859, 860, 905, 906, 931, 963, 993, 1003, 1015, 1017, 1057, 1079, 1102, 1121, 1129, 1130, 1150, 1154, 1172, 1177, 1182, 1203, 1211, 1216, 1234, 1239, 1244, 1251–1253, 1255]. number [282, 494, 631, 754, 773, 794, 975, 1241]. number-fields [407, 449, 586]. Numbers [4, 10, 15, 34, 47, 49, 91, 99, 117, 124, 145, 157, 171, 178, 182, 197, 211, 240, 250, 277, 295, 331, 336, 355, 368, 378, 398, 415, 419, 470, 479, 525, 560, 565, 566, 600, 623, 629, 636, 666, 706, 764, 772, 776, 799, 808, 810, 817, 834, 845, 882, 924, 932, 934, 939, 950, 956, 974, 994, 1012, 1062, 1066, 1086, 1126, 1137, 1173, 1193, 1214, 1224, 1244]. numerical [555, 752]. numériques [752]. Numerorum [372].

Observations [99, 1049]. obstruction [763, 1122, 1154]. Obtained [516]. octic [77, 1241]. October [75, 273, 369, 472, 617, 782, 954, 1069, 1222]. Odd [17, 127, 200, 217, 275, 331, 1083, 1203]. old [1233]. Omega [933]. omitting [1207]. One [377, 415, 478, 621, 623, 919, 980, 1130, 1151, 1152]. One-dimensional [919]. only [142, 1151]. Ono [777]. operator [695, 1218]. operators [143, 308, 530, 1000, 1042]. Optimal [27, 832, 1078]. order [100, 184, 228, 248, 362, 664, 668, 701, 764, 1054, 1204]. Orders [5, 10, 79, 176, 404, 507, 661, 741, 840, 926, 963, 966, 1185]. ordinariness [1093]. ordinarité [1093]. ordinary [217, 246, 634]. Ordnungen [10, 741]. originating [458]. Orthogonal [995, 1245]. overlap [722].

Padé [1128]. Pages [1, 14, 26, 36, 45, 55, 64, 75, 86, 97, 106, 119, 129, 139, 149, 159, 180, 192, 204, 215, 224, 239, 251, 262, 273, 281, 294, 307, 317, 323, 337, 346, 361, 369, 376, 387, 395, 403, 412, 420, 427, 441, 451, 461, 472, 483, 490, 500, 512, 533, 543, 553, 567, 580, 593, 604, 617, 633, 643, 654, 667, 686, 700, 714, 733, 744, 755, 770, 782, 796, 809, 824, 838, 856, 871, 890, 902, 911, 927, 943, 954, 965, 972, 986, 997, 1013, 1027, 1045, 1091]. Pages [768, 1005, 1022, 1035, 1056, 1069, 1081, 1113, 1125, 1136, 1148, 1159, 1171, 1184, 1196, 1209, 1222, 1230, 1240, 1251–1253]. pair [371, 431, 691, 762]. Pairs [576, 642, 866, 928]. Palindromy [1120]. Paolo [90]. paper [677, 1133, 1254]. parabolic [689, 1124]. parameterized [614]. parameters [969, 1083]. parametrization [572]. parametrizations [1058]. parametrized [1163]. Parité [291, 711]. parity [220, 291, 558, 711, 892, 993, 1101]. part [319, 504, 916, 1123]. partial [24, 704, 1143]. Partition [232, 254, 375, 454, 589, 663, 673, 1016, 1153]. Partitions [140, 184, 664, 697, 1001]. parts [124, 330, 513, 915]. Patterns [950]. Peano [305]. Pell [801, 982]. perfect [127, 331, 935]. period [113, 277, 443, 581, 689]. Periodic [113, 675, 912, 1036]. periodicity [1193]. périodiques [912, 1072]. periods [705]. Permutation [439]. permutations [267]. Perron [211]. perturbed [625]. Petersson [702]. petite [899]. pfaffian [712]. Picard [1060]. Pjateckij [321]. plane [470]. Plücker [572]. Plünnecke [606]. Poincaré [392]. Poincaré-type [392]. point [624]. Points [154, 155, 227, 260, 312, 390, 401, 446, 523, 583, 675, 740, 746, 877, 883, 899, 922, 970, 976, 1012, 1036, 1074, 1115, 1118, 1203, 1207]. Pollard [1070]. Pòlya [951]. polygons [305]. polynômes [960, 1221]. Polynomial [253, 414, 541, 559, 618, 684, 688, 716, 749, 768, 811, 815, 853, 854, 959, 960, 1147, 1213]. Polynomials [17, 28, 83, 148, 151, 243, 245, 276, 314, 384, 482, 486, 516, 562, 592, 642, 683, 699, 706, 784, 803, 814, 815, 914, 995, 1052, 1073, 1084, 1106, 1111, 1146, 1147, 1173, 1189, 1221, 1251, 1252]. Positive [121, 423, 429, 442, 448, 469, 509, 526, 606, 609, 612, 651, 848, 849, 864, 907, 930, 1178, 1193, 1207, 1245]. positivity [939]. Posner [416]. postage [288]. pour [207, 279, 752, 909, 912, 1050, 1156]. Power [44, 167, 172, 208, 269, 368, 408, 481, 491, 529, 535, 545, 557, 561, 573, 607, 760, 882, 930, 938, 966, 1011, 1030, 1042, 1051, 1108, 1167, 1217, 1246]. Powerfree [489]. powerful [1123]. powers [115, 124, 170, 198, 295, 296, 348, 658, 817, 846, 994, 1008, 1105, 1162, 1164, 1247]. practical [99, 666]. pratiques [99]. Prehomogeneous [971, 1032, 1124]. premier [690]. premiers [1015]. prescribed [202, 1075, 1215]. previous [1133]. Prielipp [658]. prim [30]. Prime [101, 124, 167, 208, 260, 296, 303, 321, 335, 408, 561, 668, 690, 799, 866, 931, 979, 983, 1015, 1059, 1108, 1151, 1203, 1223]. prime-to-adjoint [1223]. Primes [12, 110, 271, 314, 315, 365, 521, 560, 707, 724, 844, 879, 1029, 1215]. Primitive [101, 245, 359, 423, 448, 509, 1012, 1215, 1249]. Principalisation [852]. principle [373, 1154, 1223]. Pro [897, 932, 1191]. Pro- [897, 932, 1191]. probability [210]. probable [1059]. Problem [21, 186, 212, 220, 231, 288, 325, 367, 371, 431, 454, 466, 484, 506, 568, 635, 639, 658, 688, 722, 723, 725, 728, 742, 754, 873, 908, 933, 940, 1019, 1031, 1102, 1103, 1122, 1130, 1186, 1198]. problems [3, 188, 370, 437, 504, 555, 663, 763, 1129]. Prodihedral [791]. Product [404, 446, 844, 1131]. Products [79, 83, 137, 426, 642, 772, 783, 1126]. progression [365, 1215]. progressions [271, 342, 1088]. Projective [657, 874, 1009, 1114, 1203, 1207]. Proof [21, 301, 333, 388, 648, 719]. proofs [589]. proper [615]. properties [269, 352, 423, 464, 465, 625, 778, 798, 820, 826, 875, 1028]. providing [615]. proximité [1165]. proximity [519, 1165]. pseudorandom [1099]. puissance [882]. puissances [817, 1105, 1254]. pure [67, 638].

Quadratic [8, 10, 30, 34, 43, 95, 103, 104, 109, 110, 121, 125, 138, 197, 248, 268, 295, 329, 340, 343, 359, 365, 401, 407, 409, 424, 429, 442, 448, 449, 484, 509, 524, 538, 545, 546, 549, 614, 621, 627, 638, 651, 662, 691, 709, 715, 746, 751, 773, 785, 797, 806, 833, 841, 848, 849, 859, 864, 878, 913, 919, 955, 973, 981, 1029, 1040, 1077, 1094, 1172, 1177, 1178, 1193, 1206, 1231]. quadratique [125]. quadratiques [147, 785]. Quadratische [10]. quadratischen [7]. quantitatifs [279]. Quantitative [279]. Quartic [20, 136, 216, 248, 366, 484, 529, 538, 626, 687, 691, 828, 861, 947, 1016, 1140, 1153, 1167, 1221, 1241]. quartiques [1221]. quasicrystals [438]. Quaternary [1178]. quaternion [308, 1174]. Quelques [1093]. question [717, 721]. quick [301]. Quintic [703]. quotient [476]. quotients [62, 945]. qu’une [1237].

R. [181]. r.a.t [515]. r.a.t.-valued [515]. race [506]. Rademacher [528]. radical [56, 787]. raising [1063]. Ramanujan [138, 140, 247, 343, 374, 458, 467, 514, 528, 589, 645, 829, 1014, 1164, 1229]. Ramanujan-type [645]. Ramification [23, 347, 366, 602, 867, 920, 1071]. ramified [135, 142, 422, 665, 1151]. ramifiées [153, 752]. Random [820, 1138]. rang [1221]. rank [30, 206, 274, 364, 444, 676, 850, 1145, 1152, 1172, 1221]. Rankin [537, 831, 1000, 1046]. Ranks [627, 1177, 1216]. rapid [831, 1046]. rarefied [264]. rate [988]. Rational [87, 112, 219, 227, 277, 300, 401, 446, 515, 523, 546, 689, 736, 740, 784, 789, 846, 932, 935, 970, 976, 991, 1058, 1118]. rationals [31, 56, 810]. rationnels [932, 970]. Ratios [123]. ray [329]. Real [9, 77, 88, 95, 123, 171, 330, 409, 449, 510, 621, 627, 638, 678, 704, 709, 751, 810, 841, 878, 905, 906, 913, 955, 973, 1066, 1094, 1106, 1108, 1150, 1199, 1244, 1252]. réalisables [917]. Réalisation [1117]. realisations [162]. realizability [195]. realization [1117]. Reciprocal [1175, 1253]. Reciprocity [318, 333, 344, 863, 1051, 1225]. recurrence [566]. recurrences [78]. Recurrent [108, 692, 960]. récurrentes [960]. recurring [1123]. recursion [904]. Reduced [470]. Reducibility [699, 1189]. Reducible [71]. reduction [475, 539, 936, 955, 1152]. reductions [291]. reductive [436, 1098]. réduits [291]. Reduktion [936]. Reed [1074]. Refined [1172]. reflection [805]. regular [491, 1117]. regularity [12]. regulators [123, 687]. régulière [1117]. related [19, 137, 150, 331, 371, 431, 525, 555, 639, 828, 950, 979, 1009, 1131, 1181]. relatifs [870]. relation [631, 912]. Relations [34, 206, 398, 402, 540, 588, 961, 1001, 1017, 1030]. Relative [4, 20, 50, 92, 291, 391, 794, 913, 1241]. relatives [147]. Remark [214, 459, 488, 720, 721]. Remarks [611, 677, 1088, 1093, 1119, 1226]. remarques [1093]. remove [756]. Rényi [506]. répartition [495, 1015]. representable [975]. Representation [23, 365, 367, 602, 628, 691, 849, 878, 1063, 1101, 1193]. Representations [11, 135, 308, 354, 359, 423, 448, 509, 634, 657, 740, 800, 832, 864, 1033, 1040, 1114, 1119, 1226]. residual [338]. residue [93, 347, 481, 505, 563, 1018, 1028, 1121, 1132, 1169, 1242]. residues [394, 502, 548, 919, 1173]. resolution [132, 133, 220, 586, 936]. resolvent [1192]. respect [78, 1231]. restricted [140, 375, 684, 804, 1070]. result [8, 378, 467, 719, 765, 825, 1070]. Résultats [279]. results [47, 279, 619, 639, 1195]. Revêtement [903]. revêtements [1190]. reverse [728]. revisited [694, 1120]. Rham [891]. Richaud [709, 841]. Riemann [382, 383, 452, 535, 676, 732, 761, 939, 1044, 1202]. Rigidity [437, 647]. ring [95, 297, 961, 1060]. Rings [148, 172, 200, 212, 243, 245, 272, 299, 430, 463, 574, 623, 759, 963, 1011, 1139, 1177, 1213, 1232]. Rogers [374]. Rohrbach [231]. Root [101, 551, 845, 934, 1057, 1215]. Roots [259, 1106, 1252]. Rosenfeld [1219]. rotations [1053]. Roth [74]. Rubin [1200]. Rudin [1099]. Rumsey [416].

Salem [525]. Salié [937]. same [28, 1237]. sans [690]. Satisfying [529, 799, 1161]. saturated [1243]. scaled [201]. scheme [417]. Schmidt [578]. Schneider [173]. Schreier [164]. Schur [820]. second [729, 1086]. sectors [110]. Selberg [353, 389, 537, 605, 745, 831, 1046]. self [582, 944]. self-described [944]. Selmer [660, 850, 1255]. Semi [354, 647, 785, 1210]. semi-invariants [647]. Semi-local [1210]. Semi-simplicity [354]. Semi-stable [785]. semi-stables [785]. semigroup [389, 745]. Sen [675]. senary [864]. September [64, 159, 262, 361, 461, 604, 770, 943, 1056, 1209]. sequel [944]. sequence [144, 194, 264, 315, 582, 640, 939, 992]. Sequences [108, 270, 304, 330, 334, 352, 373, 400, 446, 455, 488, 502, 504, 570, 625, 637, 655, 656, 674, 692, 726, 727, 891, 960, 995, 1082, 1083, 1099, 1123, 1138, 1147]. Sequential [425]. series [8, 41, 150, 172, 182, 240, 269, 301, 349, 392, 426, 459, 469, 487, 491, 503, 529, 545, 607, 618, 630, 735, 760, 781, 790, 930, 1014, 1026, 1041, 1042, 1195, 1217, 1218, 1246]. Serre [885]. set [428, 606, 624]. Sets [71, 84, 148, 221, 280, 390, 394, 432, 446, 541, 557, 601, 646, 683, 708, 716, 768, 798, 799, 876, 907, 956, 999, 1018, 1047, 1098, 1169, 1249]. Several [480, 545, 671]. shadow [1104]. Shanks [506]. Shapiro [321, 1099]. sharpening [745]. sheaves [869]. Shelling [438]. shifts [1036]. Shimura [393, 608]. Shintani [751, 1143]. Short [47, 383, 629, 707, 1214]. Sidon [144, 334, 432, 992, 1018, 1138]. Sidon-sequence [144]. Siegel [41, 388, 584, 1000, 1130]. sieve [220, 306, 371, 396, 431, 736, 969]. sieves [628]. Sign [477]. signature [121]. signed [658]. Simple [548, 661, 1185, 1188]. simplest [924]. simplicity [354]. Simplifying [107]. simultanée [949]. Simultaneous [66, 124, 446, 691, 737, 738, 801, 949, 982]. single [40]. singular [712, 1002, 1010, 1048]. singularities [936]. singulier [1002]. Sinnott [887]. size [597, 804, 1176]. Small [12, 46, 103, 268, 379, 415, 421, 446, 524, 687, 899, 1025, 1029]. smallest [30]. smooth [748, 1194]. Solely [176, 926]. Solitary [501]. Solitons [1180]. solution [107, 325, 445, 645, 828]. Solutions [25, 216, 268, 282, 316, 320, 327, 379, 474, 597, 829, 839, 959, 1025, 1050]. solvable [498]. Solving [792, 982]. Some [39, 145, 206, 270, 284, 287, 324, 352, 353, 375, 389, 435, 516, 520, 555, 584, 601, 639, 649, 659, 745, 774, 810, 812, 818, 825, 878, 884, 984, 1071, 1088, 1093, 1119, 1149, 1201]. sommes [862, 903]. sort [664]. sous [952]. sous-additif [952]. space [548, 671, 874]. spaces [428, 476, 620, 971, 1032, 1037, 1124, 1207]. sparse [315, 784]. Special [250, 731, 781, 793, 845, 1157]. Spectra [771, 967]. Spectral [283, 426, 891]. spectrum [497, 808]. Speiser [1239]. Spiegelungsgruppe [805]. Split [314]. splittable [71]. splitting [290, 310]. Square [7, 306, 382, 625, 726, 1083, 1162]. Square-Classes [1083]. square-free [306]. squarefree [47]. Squares [205, 289, 349, 827, 864, 1105, 1174, 1194, 1248]. stabile [936]. stable [785, 936]. stables [785]. stamp [288]. standardized [24]. Stark [1200]. Stein [511]. Steinitz [248, 659]. Stephen [1252]. steps [485]. Stickelberger [125, 218, 540, 662, 731, 917, 979, 1110]. Stirling [145, 560, 1224]. Stirling-like [560]. Strahlklassenkörpern [7]. Stretching [109]. strip [977]. Structure [134, 185, 200, 264, 329, 422, 562, 571, 574, 670, 705, 723, 798, 836, 918, 989, 1007, 1201]. study [174, 253]. stufe [136]. style [719]. subadditive [952]. Subfields [538, 998, 1066]. Subgroup [393]. subgroups [94, 162, 530, 608, 701, 831, 833, 1076, 1144]. subideals [979]. sublattices [819]. subsequences [804]. subset [832, 914]. subsets [100, 228, 820, 1061]. Substitution [683]. Substitutions [188, 735]. subword [400]. suite [944]. suites [960]. Sum [61, 142, 163, 188, 432, 544, 606, 708, 798, 804, 813, 826, 844, 864, 956, 975, 999, 1109, 1116, 1121, 1175, 1242, 1253]. sum-free [956, 999]. sum-of-digits [826, 1242]. Summanden [179]. summands [179, 439]. summation [68, 1164]. summatory [1023]. Sums [29, 58, 71, 89, 118, 143, 144, 164, 166, 170, 198, 247, 250, 264, 344, 349, 356, 390, 399, 406, 419, 452, 460, 463, 480, 481, 513, 550, 557, 564, 573, 592, 605, 641, 644, 656, 658, 669, 684, 702, 730, 766, 772, 784, 790, 827, 832, 862, 903, 915, 929, 934, 937, 990, 994, 996, 1008, 1030, 1043, 1070, 1126, 1127, 1137, 1141, 1169, 1194, 1199, 1211, 1214, 1248]. Sumsets [335, 832, 1037]. Supercongruences [151]. supersingular [49, 680]. supersingulier [49]. support [908]. supports [53]. surface [936]. surfaces [81, 464, 465, 747, 1221]. Swan [1239]. Swinnerton [151, 1161]. Swinnerton-Dyer [151, 1161]. Sylow [94]. Symbol [450, 556, 1079]. symbols [505, 511, 563, 1152, 1214, 1225]. Symmetric [150, 285, 724, 740]. symmetrical [397]. Symplectic [819]. system [316]. Systems [93, 495, 507, 595, 693, 734, 1100, 1236]. Szego [853].

table [417]. taken [58]. tame [511]. tamely [135]. Tate [1201]. Teichmüller [821]. Teilwerte [24]. tenth [212]. term [19, 194, 893–895, 1026, 1162]. terms [726, 1123]. Ternary [460, 613, 691, 940, 1193]. test [1059]. th [348, 658, 882, 994, 1162, 1247]. thêta [1228]. Thaine [54]. Thakur [934]. their [425, 468, 645, 859, 996]. Theodor [173]. theorem [13, 52, 70, 74, 90, 131, 181, 289, 313, 324, 343, 363, 383, 390, 405, 606, 648, 675, 679, 723, 729, 757, 777, 786, 807, 851, 853, 918, 933, 951, 952, 1002, 1016, 1039, 1153, 1166, 1235]. théorème [13, 757, 952, 1002, 1166]. theorems [5, 39, 330, 353, 456, 596, 668, 810]. theoretic [860]. Théorie [99, 116, 757, 952]. Theory [5, 22, 27, 60, 99, 116, 131, 146, 220, 266, 367, 371, 431, 435, 663, 698, 734, 754, 757, 763, 780, 842, 952, 1032, 1104, 1182, 1255]. Theta [41, 426, 458, 459, 467, 572, 821, 1041, 1164, 1218, 1228, 1231]. theta-function [458]. Thin [85, 257]. Thomas [341, 648]. three [349, 742]. threefolds [584]. Thue [264, 320, 405, 578, 792, 828, 947, 1050, 1096, 1099, 1163]. Thue-Mahler [1050]. time [716]. Times [1057]. Titchmarsh [679, 1235]. Tits [661]. Torsion [67, 87, 430, 833, 1033, 1200, 1237]. totalement [752]. totally [77, 88, 123, 216, 609, 627, 704, 1244]. totient [19]. Tours [153]. tower [1094]. towers [812]. Trace [156, 201, 202, 308, 540, 857, 1218]. Traces [368, 958]. Transcendance [909, 1096]. Transcendence [46, 183, 611, 612, 718, 719, 765, 789, 886, 909, 974, 1096]. transcendental [117]. transfer [201]. transformation [1124]. transformations [1016]. transforms [234, 793]. translates [394]. triangle [967]. trigonometric [492]. Triple [452, 790]. trivial [859, 1012, 1206]. Trotter [57]. Truncated [230, 729]. Tschakaloff [1195]. Turán [303, 478, 639, 1087]. turns [153]. twins [321]. Twisted [781, 1085]. Twists [59, 475, 797, 1114, 1206]. Two [3, 5, 52, 62, 113, 525, 595, 652, 666, 668, 773, 854, 900, 953, 1010, 1030, 1126, 1162, 1187, 1234, 1241]. type [33, 151, 175, 244, 373, 389, 392, 442, 491, 645, 651, 709, 745, 841, 979, 1049, 1070]. types [493].

Uchiyama [164]. undecidability [739]. Uniform [39, 384, 478, 517, 808, 916, 1169]. Uniformly [352, 808]. unimodular [95, 359, 526, 1078, 1104]. Unique [195]. uniqueness [280, 425]. Unit [33, 137, 216, 291, 839, 840, 905, 906, 913, 975, 1106, 1176, 1252]. unité [291]. unités [1007, 1050]. Units [54, 58, 102, 218, 230, 345, 391, 398, 409, 418, 430, 494, 507, 519, 638, 662, 818, 858, 878, 880, 887, 990, 1007, 1050, 1200, 1210, 1245]. unity [259]. universal [217]. unobstructed [1223]. Unramified [326, 594, 634, 721, 773, 973, 1142]. Upper [198, 597, 674, 928, 957]. Urs [1146, 1251]. Ursim [1146, 1251]. use [174]. Using [39, 253, 1029].

valeurs [909, 1015]. Valuation [759, 912, 1139]. Value [121, 148, 371, 382, 383, 431, 487, 534, 592, 946, 1075, 1085, 1103]. valued [314, 515, 683, 914, 1084]. Values [24, 58, 88, 183, 232, 250, 283, 351, 424, 442, 449, 489, 625, 651, 669, 704, 766, 781, 797, 815, 845, 865, 909, 1015, 1116, 1157, 1174, 1195]. Vandermonde [302]. Vandiver [694]. Variable [268, 642]. variables [83, 545, 671, 854]. variance [929]. variants [776]. Variation [464, 465]. variations [52]. Varieties [87, 112, 126, 402, 523, 861, 1106, 1252]. Vasquez [648]. vector [428, 971, 1032, 1037, 1124]. vectors [446, 576]. Verification [341]. version [1006, 1166]. Very [879]. VI [342]. via [1002, 1173]. view [763, 1122]. view-obstruction [763, 1122]. Vinogradov [492, 940]. Voloch [719]. Volume [35, 63, 96, 128, 158, 191, 223, 261, 293, 322, 360, 386, 411, 440, 471, 499, 532, 552, 579, 603, 632, 653, 685, 713, 743, 768, 769, 795, 823, 855, 889, 910, 942, 964, 985, 1004, 1021, 1034, 1055, 1080, 1112, 1135, 1158, 1183, 1208, 1250–1253]. Voronoı [1152].

W [1252]. Wade [183]. Wagner [677]. Waring [367]. waveform [283]. way [664, 756]. Web [542]. Weierstrass [53, 1166]. Weierstrasse [24]. Weierstraßschen [24]. Weight [22, 196, 644, 980]. weighted [747]. weights [746]. Weil [112, 157, 266, 480, 659, 819, 1206]. Weintraub [460]. welding [109]. Well [1018]. where [338, 598, 931]. which [28, 232, 260, 623, 708]. Whole [154, 834]. whose [229, 446]. wild [859]. wildly [422, 665]. Wiles [618]. without [574, 690]. witnesses [615]. Witt [80, 1216]. Witt-invariant [80]. wreath [783].

Yokota [1253].

Zahlkörper [7]. Zassenhaus [413]. Zerlegungen [179]. Zero [80, 460, 508, 804]. zero-sum [804]. zeroes [1206]. zeros [103, 524, 622, 750, 888, 962, 987, 1155]. Zeta [38, 42, 53, 88, 152, 183, 210, 351, 382, 383, 449, 452, 531, 535, 548, 650, 704, 712, 732, 868, 957, 1038, 1044, 1048, 1116, 1124, 1157]. zeta-function [382, 383, 449, 535]. zeta-functions [210, 957, 1048]. Zur [7].

References

Anonymous:1990:PJa

[1] Anonymous. Pages 1–131 (January 1990). Journal of Number Theory, 34(1):??, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1990:EBa

[2] Anonymous. Editorial Board. Journal of Number Theory, 34(1):i, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

      1658 (electronic). URL http://www.sciencedirect.com/science/
      article/pii/0022314X9090046T.

Erdos:1990:TAP

[3] Paul Erdős and Gregory Freiman. On two additive problems. Journal of Number Theory, 34(1):1–12, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

      sciencedirect.com/science/article/pii/0022314X9090047U.

Endo:1990:RCN

[4] Akira Endô. The relative class numbers of certain imaginary abelian number fields and determinants. Journal of Number Theory, 34(1):13– 20, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

      1658 (electronic). URL http://www.sciencedirect.com/science/
      article/pii/0022314X9090048V.

Brzezinski:1990:TCT

[5] J. Brzezinski. On two classical theorems in the theory of orders. Journal of Number Theory, 34(1):21–32, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090049W.

Lenstra:1990:AFF

[6] H. W. Lenstra. Automorphisms of finite fields. Journal of Number Theory, 34(1):33–40, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X90900502.

Schertz:1990:EBG

[7] Reinhard Schertz. Zur expliziten Berechnung von Ganzheitsbasen in Strahlklassenkörpern über einem imaginär-quadratischen Zahlkörper. (German) [On the explicit calculation of holistic bases in beam class fields over an imaginary square number field]. Journal of Number Theory, 34(1):41–53, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090051R.

Andrews:1990:RSA

[8] George E. Andrews and Yoshiyuki Kitaoka. A result on q-series and its application to quadratic forms. Journal of Number Theory, 34(1): 54–62, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090052S.

Harman:1990:ARM

[9] Glyn Harman. Approximation of real matrices by integral matrices. Journal of Number Theory, 34(1):63–81, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090053T.

Halter-Koch:1990:QOG

[10] Franz Halter-Koch. Quadratische Ordnungen mit großer Klassenzahl. (German) [Quadratic orders with large class numbers]. Journal of Number Theory, 34(1):82–94, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9090054U.

Gouvea:1990:DGR

[11] Fernando Q. Gouvêa. Deforming Galois representations: controlling the conductor. Journal of Number Theory, 34(1):95–113, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090055V.

Gekeler:1990:RSP

[12] Ernst-Ulrich Gekeler. On regularity of small primes in function fields. Journal of Number Theory, 34(1):114–127, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9090056W.

Cherly:1990:TAF

[13] Jørgen Cherly and Jean-Marc Deshouillers. Un théorème d’addition dans F_{q}[x]. (French) [An addition theorem in F_{q}[x]]. Journal of Number Theory, 34(1):128–131, January 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090057X.

Anonymous:1990:PF

[14] Anonymous. Pages 133–250 (February 1990). Journal of Number Theory, 34(2):??, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jian-Ping:1990:DEG

[15] Zhou Jian-Ping. On the degree of extensions generated by finitely many algebraic numbers. Journal of Number Theory, 34(2):133–141, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090144G.

Inkeri:1990:CC

[16] K. Inkeri. On Catalan’s conjecture. Journal of Number Theory, 34 (2):142–152, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090145H.

Williamson:1990:ODP

[17] Clifton J. Williamson. Odd degree polynomials with dihedral Galois groups. Journal of Number Theory, 34(2):153–173, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9090146I.

Hazama:1990:DMC

[18] Fumio Hazama. Demjanenko matrix, class number, and Hodge group. Journal of Number Theory, 34(2):174–177, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9090147J.

Adhikari:1990:ETR

[19] Sukumar Das Adhikari and A. Sankaranarayanan. On an error term related to the Jordan totient function J_{k}(n). Journal of Number Theory, 34(2):178–188, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090148K.

Hymo:1990:RIB

[20] John A. Hymo and Charles J. Parry. On relative integral bases for cyclic quartic fields. Journal of Number Theory, 34(2):189–197, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090149L.

Dixmier:1990:PCE

[21] Jacques Dixmier. Proof of a conjecture by Erdős and Graham concerning the problem of Frobenius. Journal of Number Theory, 34 (2):198–209, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090150P.

Manickam:1990:TNH

[22] M. Manickam, B. Ramakrishnan, and T. C. Vasudevan. On the theory of newforms of half-integral weight. Journal of Number Theory, 34(2):210–224, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090151G.

Oda:1990:NRG

[23] Takayuki Oda. A note on ramification of the Galois representation on the fundamental group of an algebraic curve. Journal of Number Theory, 34(2):225–228, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090152H.

Schertz:1990:NNT

[24] Reinhard Schertz. Über die nenner normierter Teilwerte der Weierstraßschen p-Funktion. (German) [On the standardized partial values of the Weierstrasse p function]. Journal of Number Theory, 34(2):229– 234, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090153I.

Thomas:1990:CSF

[25] Emery Thomas. Complete solutions to a family of cubic Diophantine equations. Journal of Number Theory, 34(2):235–250, February 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9090154J.

Anonymous:1990:PMa

[26] Anonymous. Pages 251–394 (March 1990). Journal of Number Theory, 34(3):??, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bosma:1990:MTO

[27] Wieb Bosma and Cor Kraaikamp. Metrical theory for optimal continued fractions. Journal of Number Theory, 34(3):251–270, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9090135E.

Dragos:1990:NPP

[28] Chirica Dragos. On normal polynomials and polynomials which generate the same extension. Journal of Number Theory, 34(3):271– 275, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090136F.

Suwa:1990:ALF

[29] Noriyuki Suwa. A p-adic limit formula for Gauss sums. Journal of Number Theory, 34(3):276–283, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9090137G.

Berger:1990:QEN

[30] Ruth I. Berger. Quadratic extensions of number fields with elementary abelian 2-prim K_{2}(O_{F}) of smallest rank. Journal of Number Theory, 34(3):284–292, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090138H.

Hensley:1990:DBA

[31] Doug Hensley. The distribution of badly approximable rationals and continuants with bounded digits, II. Journal of Number Theory, 34(3): 293–334, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090139I.

Gupta:1990:DF

[32] Rajiv Gupta. Division fields of y^{2} = x^{3} − ax. Journal of Number Theory, 34(3):335–345, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090140M.

Hirabayashi:1990:UII

[33] Mikihito Hirabayashi and Ken ichi Yoshino. Unit indices of imaginary abelian number fields of type (2, 2, 2). Journal of Number Theory, 34(3):346–361, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090141D.

Uehara:1990:LCR

[34] Tsuyoshi Uehara. On linear congruence relations between class numbers of quadratic fields. Journal of Number Theory, 34(3):362–392, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090142E.

Anonymous:1990:AIVa

[35] Anonymous. Author index for volume 34. Journal of Number Theory, 34(3):393–394, March 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090143F.

Anonymous:1990:PMb

[36] Anonymous. Pages 1–108 (May 1990). Journal of Number Theory, 35(1): ??, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1990:EBb

[37] Anonymous. Editorial Board. Journal of Number Theory, 35(1): i, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090099D.

Thakur:1990:ZMA

[38] Dinesh S. Thakur. Zeta measure associated to F_{q}[T]. Journal of Number Theory, 35(1):1–17, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X90901006.

Nair:1990:STM

[39] R. Nair. Some theorems on metric uniform distribution using L ^{2} methods. Journal of Number Theory, 35(1):18–52, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9090101V.

Tong:1990:DAS

[40] Jingcheng Tong. Diophantine approximation of a single irrational number. Journal of Number Theory, 35(1):53–57, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9090102W.

Peters:1990:STS

[41] Meinhard Peters. Siegel theta series of degree 2 of extremal lattices. Journal of Number Theory, 35(1):58–61, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090103X.

Actor:1990:EML

[42] Alfred Actor. Evaluation of multidimensional linear zeta functions. Journal of Number Theory, 35(1):62–71, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090104Y.

Silverman:1990:MEQ

[43] Joseph H. Silverman. The Markoff equation X^{2} + Y^{2} + Z^{2} = aXY Z over quadratic imaginary fields. Journal of Number Theory, 35(1):72– 104, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9090105Z.

Bremser:1990:NIC

[44] Priscilla S. Bremser, Peter D. Schumer, and Lawrence C. Washington. A note on the incongruence of consecutive integers to a fixed power. Journal of Number Theory, 35(1):105–108, May 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X90901062.

Anonymous:1990:PJb

[45] Anonymous. Pages 109–226 (June 1990). Journal of Number Theory, 35 (2):??, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Tubbs:1990:AGS

[46] Robert Tubbs. Algebraic groups and small transcendence degree, II. Journal of Number Theory, 35(2):109–127, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X90901073.

Filaseta:1990:SIR

[47] Michael Filaseta. Short interval results for squarefree numbers. Journal of Number Theory, 35(2):128–149, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X90901084.

Tenenbaum:1990:LDE

[48] Gérald Tenenbaum and Hisashi Yokota. Length and denominators of Egyptian fractions, III. Journal of Number Theory, 35(2):150– 156, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90901095.

Chellali:1990:CNB

[49] Mustapha Chellali. Congruences entre nombres de Bernoulli–Hurwitz dans le cas supersingulier. (French) [Congruences between Bernoulli– Hurwitz numbers in the supersingular case]. Journal of Number Theory, 35(2):157–179, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090110D.

Greither:1990:RIN

[50] Cornelius Greither. Relative integral normal bases in Q(ζ_{p}). Journal of Number Theory, 35(2):180–193, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X90901114.

Skinner:1990:DE

[51] Christopher M. Skinner. On the diophantine equation ap^{x} + bq^{y} = c + dp^{z}q^{w}. Journal of Number Theory, 35(2):194–207, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X90901125.

Bailey:1990:TVL

[52] D. F. Bailey. Two p^{3} variations of Lucas’ Theorem. Journal of Number Theory, 35(2):208–215, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X90901136.

Kato:1990:LAH

[53] Goro Kato. Lifted p-adic homology with compact supports of the Weierstrass family and its zeta endomorphism. Journal of Number Theory, 35(2):216–223, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X90901147.

Lettl:1990:NTC

[54] Günter Lettl. A note on Thaine’s circular units. Journal of Number Theory, 35(2):224–226, June 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X90901158.

Anonymous:1990:PJc

[55] Anonymous. Pages 227–350 (July 1990). Journal of Number Theory, 35 (3):??, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jacobson:1990:FAE

[56] Eliot Jacobson and William Yslas Vélez. Fields arithmetically equivalent to a radical extension of the rationals. Journal of Number Theory, 35(3):227–246, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X90901169.

Wan:1990:LTC

[57] Daqing Wan. On the Lang–Trotter conjecture. Journal of Number Theory, 35(3):247–268, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090117A.

Everest:1990:CVT

[58] G. R. Everest. Counting the values taken by sums of S-units. Journal of Number Theory, 35(3):269–286, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090118B.

Hijikata:1990:TN

[59] Hiroaki Hijikata, Arnold K. Pizer, and Thomas R. Shemanske. Twists of newforms. Journal of Number Theory, 35(3):287–324, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9090119C.

Fishburn:1990:CFA

[60] Peter C. Fishburn. On a contribution of Freiman to additive number theory. Journal of Number Theory, 35(3):325–334, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9090120G.

Schlickewei:1990:LEI

[61] Hans Peter Schlickewei. Linear equations in integers with bounded sum of digits. Journal of Number Theory, 35(3):335–344, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X90901217.

Bang:1990:NTQ

[62] Thøger Bang and Bent Fuglede. No two quotients of normalized binomial mid-coefficients are equal. Journal of Number Theory, 35(3): 345–349, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90901228.

Anonymous:1990:AIVb

[63] Anonymous. Author index for volume 35. Journal of Number Theory, 35(3):350, July 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90901239.

Anonymous:1990:PS

[64] Anonymous. Pages 1–132 (September 1990). Journal of Number Theory, 36(1):??, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1990:EBc

[65] Anonymous. Editorial Board. Journal of Number Theory, 36(1):i, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90900018.

Dorner:1990:SDE

[66] Eberhard Dörner. Simultaneous diagonal equations over certain p-adic fields. Journal of Number Theory, 36(1):1–11, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X90900029.

Fung:1990:TGE

[67] G. W. Fung, H. Ströher, H. C. Williams, and H. G. Zimmer. Torsion groups of elliptic curves with integral j-invariant over pure cubic fields. Journal of Number Theory, 36(1):12–45, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9090003A.

Rausch:1990:SFA

[68] Ulrich Rausch. A summation formula in algebraic number fields and applications, I. Journal of Number Theory, 36(1):46–79, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9090004B.

Angermuller:1990:GEI

[69] G. Angermüller. A generalization of Ehrenfeucht’s irreducibility criterion. Journal of Number Theory, 36(1):80–84, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9090005C.

Dodson:1990:DTD

[70] M. M. Dodson, B. P. Rynne, and J. A. G. Vickers. Dirichlet’s theorem and diophantine approximation on manifolds. Journal of Number Theory, 36(1):85–88, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090006D.

Erdos:1990:RSS

[71] Paul Erdős and Abraham Zaks. Reducible sums and splittable sets. Journal of Number Theory, 36(1):89–94, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090007E.

Horie:1990:CEE

[72] Mitsuko Horie. On central extensions of elementary abelian fields. Journal of Number Theory, 36(1):95–107, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090008F.

Stern:1990:ENG

[73] Leonid Stern. On the equality of norm groups of global fields. Journal of Number Theory, 36(1):108–126, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090009G.

Gross:1990:NRT

[74] Robert Gross. A note on Roth’s Theorem. Journal of Number Theory, 36(1):127–132, September 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090010O.

Anonymous:1990:PO

[75] Anonymous. Pages 133–256 (October 1990). Journal of Number Theory, 36(2):??, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Schultheis:1990:CKF

[76] Fred Schultheis. Carlitz–Kummer function fields. Journal of Number Theory, 36(2):133–144, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X90900683.

Pohst:1990:MDT

[77] M. Pohst, J. Martinet, and F. Diaz y Diaz. The minimum discriminant of totally real octic fields. Journal of Number Theory, 36(2):145– 159, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90900694.

Grabner:1990:CDE

[78] P. J. Grabner and R. F. Tichy. Contributions to digit expansions with respect to linear recurrences. Journal of Number Theory, 36(2):160– 169, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90900708.

Kleinert:1990:OMF

[79] Ernst Kleinert. On orders and multiple fibre products. Journal of Number Theory, 36(2):170–186, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9090071X.

Kodama:1990:FHF

[80] Tetsuo Kodama and Tadashi Washio. A family of hyperelliptic function fields with Hasse–Witt-invariant zero. Journal of Number Theory, 36(2):187–200, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090072Y.

Kuwata:1990:CHE

[81] Masato Kuwata. The canonical height and elliptic surfaces. Journal of Number Theory, 36(2):201–211, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090073Z.

Valenza:1990:EFN

[82] R. J. Valenza. Elasticity of factorization in number fields. Journal of Number Theory, 36(2):212–218, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X90900742.

Beauzamy:1990:PPM

[83] Bernard Beauzamy, Enrico Bombieri, Per Enflo, and Hugh L. Montgomery. Products of polynomials in many variables. Journal of Number Theory, 36(2):219–245, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X90900753.

El-Zanati:1990:DCC

[84] Saad I. El-Zanati and William R. R. Transue. On dynamics of certain Cantor sets. Journal of Number Theory, 36(2):246–253, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90900764.

Jia:1990:TBF

[85] Xing-De Jia. Thin bases for finite abelian groups. Journal of Number Theory, 36(2):254–256, October 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X90900775.

Anonymous:1990:PN

[86] Anonymous. Pages 257–381 (November 1990). Journal of Number Theory, 36(3):??, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Flynn:1990:LRT

[87] E. V. Flynn. Large rational torsion on Abelian varieties. Journal of Number Theory, 36(3):257–265, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9090089A.

Halbritter:1990:CVZ

[88] U. Halbritter and M. Pohst. On the computation of the values of zeta functions of totally real cubic fields. Journal of Number Theory, 36(3):266–288, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090090E.

Bruggeman:1990:DSF

[89] Roelof W. Bruggeman. Dedekind sums and Fourier coefficients of modular forms. Journal of Number Theory, 36(3):289–321, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X90900915.

Petermann:1990:ATP

[90] Y.-F. S. Pétermann. About a theorem of Paolo Codecà’s and Ω-estimates for arithmetical convolutions: Addendum. Journal of Number Theory, 36(3):322–327, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X90900926.

Akiyama:1990:LNA

[91] Shigeki Akiyama. Lehmer numbers and an asymptotic formula for π. Journal of Number Theory, 36(3):328–331, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X90900937.

Endo:1990:IFR

[92] Akira Endô. On an index formula for the relative class number of a cyclotomic number field. Journal of Number Theory, 36(3):332–338, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90900948.

Benson:1990:CIC

[93] Steven R. Benson. Canonical invariants for corresponding residue systems in P -adic fields. Journal of Number Theory, 36(3):339–353, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90900959.

Harnchoowong:1990:SSI

[94] Ajchara Harnchoowong and Wen-Ch’ing Winnie Li. Sylow subgroups of ideal class groups with moduli. Journal of Number Theory, 36 (3):354–372, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090096A.

Takada:1990:GCN

[95] Ichiro Takada. On the genera and the class number of unimodular lattices over the ring of integers in real quadratic fields. Journal of Number Theory, 36(3):373–379, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9090097B.

Anonymous:1990:AIVc

[96] Anonymous. Author index for volume 36. Journal of Number Theory, 36(3):380–381, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9090098C.

Anonymous:1991:PJa

[97] Anonymous. Pages 1–122 (January 1991). Journal of Number Theory, 37(1):??, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1991:EBa

[98] Anonymous. Editorial Board. Journal of Number Theory, 37(1):i, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05800216.

Margenstern:1991:NPT

[99] Maurice Margenstern. Les nombres pratiques: théorie, observations et conjectures. (French) [Practical numbers: theory, observations and conjectures]. Journal of Number Theory, 37(1):1–36, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05800228.

Jia:1991:OSA

[100] Xing-De Jia. On the order of subsets of asymptotic bases. Journal of Number Theory, 37(1):37–46, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0580023X.

Murata:1991:MLP

[101] Leo Murata. On the magnitude of the least prime primitive root. Journal of Number Theory, 37(1):47–66, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05800241.

Marszalek:1991:MUC

[102] Roman Marszałek. Minkowski units in a class of metabelian fields. Journal of Number Theory, 37(1):67–91, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05800253.

Cochrane:1991:SZQ

[103] Todd Cochrane. Small zeros of quadratic forms modulo p, III. Journal of Number Theory, 37(1):92–99, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05800265.

Dummit:1991:CIL

[104] D. S. Dummit, D. Ford, H. Kisilevsky, and J. W. Sands. Computation of Iwasawa lambda invariants for imaginary quadratic fields. Journal of Number Theory, 37(1):100–121, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05800277.

Anonymous:1991:A

[105] Anonymous. Addendum. Journal of Number Theory, 37(1):122, January 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05800289.

Anonymous:1991:PF

[106] Anonymous. Pages 123–252 (February 1991). Journal of Number Theory, 37(2):??, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Steiner:1991:SSL

[107] Ray Steiner and Nikos Tzanakis. Simplifying the solution of Ljunggren’s equation X^{2} + 1 = 2Y^{4}. Journal of Number Theory, 37(2):123–132, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05800290.

Habsieger:1991:LRS

[108] Laurent Habsieger. Linear recurrent sequences and irrationality measures. Journal of Number Theory, 37(2):133–145, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05800307.

Gerstein:1991:SWI

[109] Larry J. Gerstein. Stretching and welding indecomposable quadratic forms. Journal of Number Theory, 37(2):146–151, February 1991. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0022314X05800319.

Zarzycki:1991:DPI

[110] Piotr Zarzycki. Distribution of primes of imaginary quadratic fields in sectors. Journal of Number Theory, 37(2):152–160, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05800320.

Kim:1991:FGD

[111] Jae Moon Kim. Formal groups and Dirichlet L-functions. Journal of Number Theory, 37(2):161–167, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05800332.

Hazama:1991:MWG

[112] Fumio Hazama. On the Mordell–Weil group of certain abelian varieties defined over the rational function field. Journal of Number Theory, 37(2):168–172, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05800344.

Levesque:1991:TFP

[113] C. Levesque and G. Rhin. Two families of periodic Jacobi algorithms with period lengths going to infinity. Journal of Number Theory, 37(2):173–180, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05800356.

Chen:1991:CDAa

[114] Yong-Gao Chen. On a conjecture in diophantine approximations, II. Journal of Number Theory, 37(2):181–198, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05800368.

Brudern:1991:ADI

[115] Jörg Brüdern. Additive diophantine inequalities with mixed powers, III. Journal of Number Theory, 37(2):199–210, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0580037X.

Perrine:1991:GTM

[116] Serge Perrine. Sur une généralisation de la Théorie de Markoff. (French) [On a generalization of Markoff’s theory]. Journal of Number Theory, 37(2):211–230, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05800381.

Amou:1991:ACT

[117] Masaaki Amou. Approximation to certain transcendental decimal fractions by algebraic numbers. Journal of Number Theory, 37(2):231–241, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05800393.

Thakur:1991:GSF

[118] Dinesh S. Thakur. Gauss sums for function fields. Journal of Number Theory, 37(2):242–252, February 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0580040X.

Anonymous:1991:PMa

[119] Anonymous. Pages 253–370 (March 1991). Journal of Number Theory, 37(3):??, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Borwein:1991:I

[120] Peter B. Borwein. On the irrationality of ∑(1/(q^{n}+r)). Journal of Number Theory, 37(3):253–259, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05800411.

Aggarwal:1991:LPV

[121] Satish K. Aggarwal and D. P. Gupta. Least positive value of nonhomogeneous indefinite quadratic forms of signature −3. Journal of Number Theory, 37(3):260–278, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05800423.

Wolke:1991:CCM

[122] D. Wolke. On a certain class of multiplicative functions. Journal of Number Theory, 37(3):279–287, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05800435.

Costa:1991:RRT

[123] Antone Costa and Eduardo Friedman. Ratios of regulators in totally real extensions of number fields. Journal of Number Theory, 37(3):288– 297, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05800447.

Tolev:1991:SDF

[124] D. I. Tolev. On the simultaneous distribution of the fractional parts of different powers of prime numbers. Journal of Number Theory, 37(3): 298–306, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05800459.

Cassou-Nogues:1991:ESQ

[125] Ph. Cassou-Nogués and M. J. Taylor. Un élément de Stickelberger quadratique. (French) [A quadratic Stickelberger element]. Journal of Number Theory, 37(3):307–342, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05800460.

Grundman:1991:AGH

[126] H. G. Grundman. The arithmetic genus of Hilbert modular varieties over non–Galois cubic fields. Journal of Number Theory, 37(3):343– 365, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05800472.

Starni:1991:EFO

[127] Paolo Starni. On the Euler’s factor of an odd perfect number. Journal of Number Theory, 37(3):366–369, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05800484.

Anonymous:1991:AIVa

[128] Anonymous. Author index for volume 37. Journal of Number Theory, 37(3):370, March 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05800496.

Anonymous:1991:PMb

[129] Anonymous. Pages 1–133 (May 1991). Journal of Number Theory, 38(1): ??, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1991:EBb

[130] Anonymous. Editorial Board. Journal of Number Theory, 38(1): i, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190088S.

Kan:1991:TMN

[131] Jiahai Kan. A theorem in multiplicative number theory. Journal of Number Theory, 38(1):1–17, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9190089T.

Gaal:1991:RIFa

[132] I. Gaál, A. Pethö, and M. Pohst. On the resolution of index form equations in biquadratic number fields, I. Journal of Number Theory, 38(1):18–34, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190090X.

Gaal:1991:RIFb

[133] I. Gaál, A. Pethö, and M. Pohst. On the resolution of index form equations in biquadratic number fields, II. Journal of Number Theory, 38(1):35–51, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190091O.

Nagaraj:1991:SIM

[134] D. S. Nagaraj. The structure of the Iwasawa module associated with a Z_{p}^{r}-extension of a p-adic local field of characteristic 0. Journal of Number Theory, 38(1):52–57, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190092P.

Reimann:1991:RTR

[135] Harry Reimann. Representations of tamely ramified p-adic division and matrix algebras. Journal of Number Theory, 38(1):58– 105, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190093Q.

Parnami:1991:SQF

[136] J. C. Parnami, M. K. Agrawal, and A. R. Rajwade. On the stufe of quartic fields. Journal of Number Theory, 38(1):106–109, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9190094R.

Bressoud:1991:CUR

[137] David M. Bressoud. On a cyclotomic unit and related products. Journal of Number Theory, 38(1):110–117, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9190095S.

Williams:1991:ARC

[138] Kenneth S. Williams. On an assertion of Ramanujan concerning binary quadratic forms. Journal of Number Theory, 38(1):118–133, May 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9190096T.

Anonymous:1991:PJb

[139] Anonymous. Pages 135–250 (June 1991). Journal of Number Theory, 38 (2):??, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Almkvist:1991:HRF

[140] Gert Almkvist and George E. Andrews. A Hardy–Ramanujan formula for restricted partitions. Journal of Number Theory, 38(2):135– 144, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190079Q.

Almkvist:1991:EAF

[141] Gert Almkvist. Exact asymptotic formulas for the coefficients of nonmodular functions. Journal of Number Theory, 38(2):145–160, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9190080U.

Prapavessi:1991:JSH

[142] Despina T. Prapavessi. On Jacobi sum Hecke characters ramified only at 2. Journal of Number Theory, 38(2):161–184, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9190081L.

Lu:1991:HSH

[143] Hongwen Lu. Hirzebruch sums and Hecke operators. Journal of Number Theory, 38(2):185–195, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190082M.

Erdos:1991:SSS

[144] P. Erdős and R. Freud. On sums of a Sidon-sequence. Journal of Number Theory, 38(2):196–205, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190083N.

Tsumura:1991:SCB

[145] Hirofumi Tsumura. On some congruences for the Bell numbers and for the Stirling numbers. Journal of Number Theory, 38(2):206– 211, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190084O.

Okada:1991:KTF

[146] Shozo Okada. Kummer’s theory for function fields. Journal of Number Theory, 38(2):212–215, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190085P.

Massy:1991:BNE

[147] Richard Massy. Bases normales d’entiers relatives quadratiques. Journal of Number Theory, 38(2):216–239, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9190086Q.

Bremser:1991:VSD

[148] Priscilla S. Bremser and Javier Gomez-Calderon. Value sets of Dickson polynomials over Galois rings. Journal of Number Theory, 38(2): 240–250, June 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190087R.

Anonymous:1991:PJc

[149] Anonymous. Pages 251–378 (July 1991). Journal of Number Theory, 38 (3):??, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Tamura:1991:SCF

[150] Jun ichi Tamura. Symmetric continued fractions related to certain series. Journal of Number Theory, 38(3):251–264, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X91900187.

Coster:1991:SAS

[151] M. J. Coster and L. van Hamme. Supercongruences of Atkin and Swinnerton-Dyer type for Legendre polynomials. Journal of Number Theory, 38(3):265–286, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X91900198.

Pas:1991:LZF

[152] Johan Pas. Local zeta functions and Meuser’s invariant functions. Journal of Number Theory, 38(3):287–299, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9190020C.

Perret:1991:TRI

[153] Marc Perret. Tours ramifiées infinies de corps de classes. (French) [Infinite branched turns of class bodies]. Journal of Number Theory, 38(3):300–322, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X91900213.

Ayad:1991:PEC

[154] Mohamed Ayad. Points Δ entiers sur les courbes elliptiques. (French) [Whole Δ points on elliptic curves]. Journal of Number Theory, 38(3): 323–337, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X91900224.

Hooley:1991:NPC

[155] C. Hooley. On the number of points on a complete intersection over a finite field. Journal of Number Theory, 38(3):338–358, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X91900235.

Epkenhans:1991:TFD

[156] Martin Epkenhans. Trace forms of dyadic number fields. Journal of Number Theory, 38(3):359–365, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X91900246.

Odoni:1991:WNC

[157] R. W. K. Odoni. Weil numbers and CM fields, II. Journal of Number Theory, 38(3):366–377, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X91900257.

Anonymous:1991:AIVb

[158] Anonymous. Author index for volume 38. Journal of Number Theory, 38(3):378, July 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X91900268.

Anonymous:1991:PS

[159] Anonymous. Pages 1–128 (September 1991). Journal of Number Theory, 39(1):??, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1991:EBc

[160] Anonymous. Editorial Board. Journal of Number Theory, 39(1):i, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X91900279.

Cornell:1991:NCG

[161] Gary Cornell. A note on the class group of composita. Journal of Number Theory, 39(1):1–4, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190028A.

Schneps:1991:ERS

[162] Leila Schneps. Explicit realisations of subgroups of GL_{2}(F_{3}) as Galois groups. Journal of Number Theory, 39(1):5–13, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9190029B.

Zhang:1991:CES

[163] Zhenxiang Zhang. On a conjecture of Erdős on the sum ∑_{p≤n} 1(p log p). Journal of Number Theory, 39(1):14–17, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9190030F.

Lachaud:1991:ASC

[164] Gilles Lachaud. Artin–Schreier curves, exponential sums, and the Carlitz–Uchiyama bound for geometric codes. Journal of Number Theory, 39(1):18–40, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X91900316.

Mignotte:1991:FC

[165] M. Mignotte and N. Tzanakis. On a family of cubics. Journal of Number Theory, 39(1):41–49, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X91900327.

Wang:1991:JSM

[166] Jun Wang. On the Jacobi sums modulo P^{n}. Journal of Number Theory, 39(1):50–64, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X91900338.

Sander:1991:PPD

[167] J. W. Sander. Prime power divisors of \binom{2n}{n}. Journal of Number Theory, 39(1):65–74, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X91900349.

Conway:1991:LFD

[168] J. H. Conway and N. J. A. Sloane. Lattices with few distances. Journal of Number Theory, 39(1):75–90, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9190035A.

Chen:1991:CDAb

[169] Yong-Gao Chen. On a conjecture in Diophantine approximations III. Journal of Number Theory, 39(1):91–103, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9190036B.

Choudhry:1991:SEP

[170] Ajai Choudhry. On sums of eighth powers. Journal of Number Theory, 39(1):104–107, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190037C.

Becker:1991:EDE

[171] Paul-Georg Becker. Exponential diophantine equations and the irrationality of certain real numbers. Journal of Number Theory, 39 (1):108–116, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190038D.

Hoechsmann:1991:AHP

[172] Klaus Hoechsmann and Jürgen Ritter. The Artin–Hasse power series and p-adic group rings. Journal of Number Theory, 39(1):117–128, September 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190039E.

Bundschuh:1991:TS

[173] Peter Bundschuh and Hans Zassenhaus. Theodor Schneider (1911– 1988). Journal of Number Theory, 39(2):129–143, October 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9190040I.

Delange:1991:UMM

[174] Hubert Delange. On the use of the method of moments for the study of additive functions. Journal of Number Theory, 39(2):144– 161, October 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X91900419.

Schweiger:1991:IMM

[175] Fritz Schweiger. Invariant measures for maps of continued fraction type. Journal of Number Theory, 39(2):162–174, October 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9190042A.

Srinivasan:1991:OSA

[176] S. Srinivasan. On orders solely of abelian groups, III. Journal of Number Theory, 39(2):175–180, October 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9190043B. See correction [926].

Schmidt:1991:CEB

[177] Wolfgang M. Schmidt. Construction and estimation of bases in function fields. Journal of Number Theory, 39(2):181–224, October 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9190044C.

Vaananen:1991:LIM

[178] K. Väänänen and R. Wallisser. A linear independence measure for certain p-adic numbers. Journal of Number Theory, 39(2):225–236, October 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190045D.

Wolke:1991:GZN

[179] Dieter Wolke. Über Goldbach-Zerlegungen mit nahezu gleichen Summanden. (German) [On Goldbach decompositions with almost equal summands]. Journal of Number Theory, 39(2):237–244, October 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9190046E.

Anonymous:1991:PN

[180] Anonymous. Pages 245–375 (November 1991). Journal of Number Theory, 39(3):??, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

McQuillan:1991:TRG

[181] Donald L. McQuillan. On a theorem of R. Gilmer. Journal of Number Theory, 39(3):245–250, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9190047F.

Tsumura:1991:NAD

[182] Hirofumi Tsumura. A note on q-analogues of the Dirichlet series and q-Bernoulli numbers. Journal of Number Theory, 39(3):251–256, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190048G.

Damamme:1991:TVC

[183] G. Damamme and Y. Hellegouarch. Transcendence of the values of the Carlitz zeta function by Wade’s method. Journal of Number Theory, 39(3):257–278, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190049H.

Kolitsch:1991:OGF

[184] Louis Worthy Kolitsch. M-order generalized Frobenius partitions with M colors. Journal of Number Theory, 39(3):279–284, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9190050L.

Schertz:1991:GEF

[185] Reinhard Schertz. Galoismodulstruktur und elliptische Funktionen. (German) [Galois module structure and elliptic functions]. Journal of Number Theory, 39(3):285–326, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9190051C.

Yokota:1991:PEG

[186] Hisashi Yokota. On a problem of Erdős and Graham. Journal of Number Theory, 39(3):327–338, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9190052D.

Bombieri:1991:GFE

[187] E. Bombieri and J. Mueller. The generalized Fermat equation in function fields. Journal of Number Theory, 39(3):339–350, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9190053E.

Dumont:1991:DSP

[188] J. M. Dumont and A. Thomas. Digital sum problems and substitutions on a finite alphabet. Journal of Number Theory, 39(3):351–366, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190054F.

Sobol:1991:NMI

[189] Ilya M. Sobol’ and Oleg V. Nuzhdin. A new measure of irregularity of distribution. Journal of Number Theory, 39(3):367–373, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9190055G.

Anonymous:1991:E

[190] Anonymous. Erratum. Journal of Number Theory, 39(3):374, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9190056H.

Anonymous:1991:AIVc

[191] Anonymous. Author index for volume 39. Journal of Number Theory, 39(3):375, November 1991. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9190057I.

Anonymous:1992:PJa

[192] Anonymous. Pages 1–125 (January 1992). Journal of Number Theory, 40(1):??, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1992:EBa

[193] Anonymous. Editorial Board. Journal of Number Theory, 40(1):i, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290023I.

Vardi:1992:ETG

[194] Ilan Vardi. The error term in Golomb’s sequence. Journal of Number Theory, 40(1):1–11, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290024J.

Jensen:1992:URF

[195] C. U. Jensen and A. Prestel. Unique realizability of finite abelian 2-groups as Galois groups. Journal of Number Theory, 40(1):12–31, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290025K.

Manickam:1992:DMF

[196] M. Manickam, B. Ramakrishnan, and T. C. Vasudevan. Diagonalising modular forms of half-integral weight. Journal of Number Theory, 40(1):32–37, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290026L.

Gonzalez:1992:CNQ

[197] Cristian D. González. Class numbers of quadratic function fields and continued fractions. Journal of Number Theory, 40(1):38–59, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290027M.

Norton:1992:UBS

[198] Karl K. Norton. Upper bounds for sums of powers of divisor functions. Journal of Number Theory, 40(1):60–85, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9290028N.

Cohen:1992:DE

[199] Edward L. Cohen. On the diophantine equation x^{2} − Dy^{2} = nz^{2}. Journal of Number Theory, 40(1):86–91, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290029O.

Erez:1992:HSR

[200] Boas Erez and Jorge Morales. The Hermitian structure of rings of integers in odd degree abelian extensions. Journal of Number Theory, 40(1):92–104, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290030S.

Kruskemper:1992:STF

[201] Martin Krüskemper. On the scaled trace forms and the transfer of a number field extension. Journal of Number Theory, 40(1):105–119, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290031J.

Kruskemper:1992:ANF

[202] Martin Krüskemper. Algebraic number field extensions with prescribed trace form. Journal of Number Theory, 40(1):120–124, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290032K.

Moroz:1992:LE

[203] B. Z. Moroz. Letter to the editor. Journal of Number Theory, 40(1): 125, January 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290033L.

Anonymous:1992:PF

[204] Anonymous. Pages 127–253 (February 1992). Journal of Number Theory, 40(2):??, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Balasubramanian:1992:ACS

[205] R. Balasubramanian and K. Soundararajan. On the additive completion of squares, II. Journal of Number Theory, 40(2):127–129, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290034M.

Gadea:1992:SRR

[206] Nicolas Santa Gadea. On some relations for the rank moduli 9 and 12. Journal of Number Theory, 40(2):130–145, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9290035N.

Balazard:1992:GDP

[207] M. Balazard, J.-L Nicolas, C. Pomerance, and G. Tenenbaum. Grandes déviations pour certaines fonctions arithmétiques. (French) [Large deviations for certain arithmetic functions]. Journal of Number Theory, 40(2):146–164, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290036O.

Washington:1992:KLP

[208] Lawrence C. Washington. Kummer’s lemma for prime power cyclotomic fields. Journal of Number Theory, 40(2):165–173, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9290037P.

Efrat:1992:A

[209] Isaac Efrat. On a GL(3) analog of |η(z)|. Journal of Number Theory, 40(2):174–186, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290038Q.

Matsumoto:1992:APM

[210] Kohji Matsumoto. Asymptotic probability measures of zeta-functions of algebraic number fields. Journal of Number Theory, 40(2):187–210, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290039R.

Lind:1992:MPN

[211] Douglas Lind. Matrices of Perron numbers. Journal of Number Theory, 40(2):211–217, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290040V.

Shlapentokh:1992:HTP

[212] Alexandra Shlapentokh. Hilbert’s tenth problem for rings of algebraic functions of characteristic 0. Journal of Number Theory, 40(2):218– 236, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290041M.

vanderPoorten:1992:FCF

[213] A. J. van der Poorten and J. Shallit. Folded continued fractions. Journal of Number Theory, 40(2):237–250, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290042N.

Manickam:1992:RAL

[214] M. Manickam, B. Ramakrishnan, and T. C. Vasudevan. On a remark of Atkin–Lehner. Journal of Number Theory, 40(2):251–253, February 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290043O.

Anonymous:1992:PMa

[215] Anonymous. Pages 255–376 (March 1992). Journal of Number Theory, 40(3):??, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Edgar:1992:USE

[216] Hugh Edgar, Jonathan Gordon, and Liang-Cheng Zhang. On unit solutions of the equation xyz = x + y + z in totally imaginary quartic fields. Journal of Number Theory, 40(3):255–263, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X92900016.

Kucera:1992:BOE

[217] Radan Kučera. On bases of odd and even universal ordinary distributions. Journal of Number Theory, 40(3):264–283, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X92900027.

Kucera:1992:BSI

[218] Radan Kučera. On bases of the Stickelberger ideal and of the group of circular units of a cyclotomic field. Journal of Number Theory, 40(3): 284–316, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92900038.

Scheutzow:1992:CRC

[219] Alexander Scheutzow. Computing rational cohomology and Hecke eigenvalues for Bianchi groups. Journal of Number Theory, 40(3):317– 328, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92900049.

Ng:1992:CRP

[220] Eugene K.-S Ng. A conditional resolution of the parity problem in sieve theory. Journal of Number Theory, 40(3):329–335, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9290005A.

Hensley:1992:CFC

[221] Doug Hensley. Continued fraction Cantor sets, Hausdorff dimension, and functional analysis. Journal of Number Theory, 40(3):336–358, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290006B.

Clement:1992:GFA

[222] Rosario Clement. The genus field of an algebraic function field. Journal of Number Theory, 40(3):359–375, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290007C.

Anonymous:1992:AIVa

[223] Anonymous. Author index for volume 40. Journal of Number Theory, 40(3):376, March 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290008D.

Anonymous:1992:PMb

[224] Anonymous. Pages 1–129 (May 1992). Journal of Number Theory, 41(1): ??, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1992:EBb

[225] Anonymous. Editorial Board. Journal of Number Theory, 41(1): i, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92900773.

Garfield:1992:DBC

[226] Richard Garfield and Herbert S. Wilf. The distribution of the binomial coefficients modulo p. Journal of Number Theory, 41(1):1– 5, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92900784.

Schoof:1992:ACM

[227] René Schoof. Algebraic curves over F_{2} with many rational points. Journal of Number Theory, 41(1):6–14, May 1992. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X92900795.

Chen:1992:EOS

[228] Sheng Chen and Weizhen Gu. Exact order of subsets of asymptotic bases. Journal of Number Theory, 41(1):15–21, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X92900809.

Roy:1992:MWC

[229] Damien Roy. Matrices whose coefficients are linear forms in logarithms. Journal of Number Theory, 41(1):22–47, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290081Y.

Levesque:1992:TU

[230] Claude Levesque. Truncated units. Journal of Number Theory, 41(1): 48–68, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290082Z.

Nathanson:1992:PRF

[231] Melvyn B. Nathanson. On a problem of Rohrbach for finite groups. Journal of Number Theory, 41(1):69–76, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X92900832.

Tamba:1992:PFW

[232] Manvendra Tamba. On a partition function which assumes all integral values. Journal of Number Theory, 41(1):77–86, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X92900843.

Maohua:1992:DEa

[233] Le Maohua. On the diophantine equation x^{2} + D = 4p^{n}. Journal of Number Theory, 41(1):87–97, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X92900854.

Satoh:1992:IIT

[234] Junya Satoh. Iwasawa λ-invariants of Γ-transforms. Journal of Number Theory, 41(1):98–101, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X92900865.

Lim:1992:GJF

[235] Chong Hai Lim. The geometry of the Jacobian of the Fermat curve of exponent five. Journal of Number Theory, 41(1):102–115, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X92900876.

Jia:1992:EBF

[236] Xing-De Jia. Extremal bases for finite cyclic groups. Journal of Number Theory, 41(1):116–127, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X92900887.

Anonymous:1992:CA

[237] Anonymous. Corrigendum and addendum. Journal of Number Theory, 41(1):128, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92900898.

Anonymous:1992:E

[238] Anonymous. Erratum. Journal of Number Theory, 41(1):129, May 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9290090C.

Anonymous:1992:PJb

[239] Anonymous. Pages 131–255 (June 1992). Journal of Number Theory, 41 (2):??, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Knopfmacher:1992:ISE

[240] Arnold Knopfmacher and John Knopfmacher. Infinite series expansions for p-adic numbers. Journal of Number Theory, 41(2):131– 145, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92901134.

Akbik:1992:NGF

[241] Safwan Akbik. Normal generators of finite fields. Journal of Number Theory, 41(2):146–149, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X92901145.

Thakur:1992:CFE

[242] Dinesh S. Thakur. Continued fraction for the exponential for F_{q}[T]. Journal of Number Theory, 41(2):150–155, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X92901156.

Brawley:1992:FPG

[243] Joel V. Brawley and Gary L. Mullen. Functions and polynomials over Galois rings. Journal of Number Theory, 41(2):156–166, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X92901167.

Hahn:1992:MTC

[244] Sang G. Hahn. On Mirimanoff type congurences. Journal of Number Theory, 41(2):167–171, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X92901178.

Bremser:1992:PPE

[245] Priscilla S. Bremser and Javier Gomez-Calderon. Polynomials and primitive elements in Galois rings. Journal of Number Theory, 41(2):172– 177, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92901189.

Gouvea:1992:OHA

[246] Fernando Q. Gouveâ. On the ordinary Hecke algebra. Journal of Number Theory, 41(2):178–198, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290119A.

Li:1992:CSA

[247] Wen-Ch’ing Winnie Li and Keqin Feng. Character sums and abelian Ramanujan graphs. Journal of Number Theory, 41(2):199–217, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9290120E.

Hymo:1992:SCO

[248] John A. Hymo and Charles J. Parry. Steinitz classes of order 2 in quadratic and quartic fields. Journal of Number Theory, 41(2):218– 224, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92901215.

Soundararajan:1992:IMF

[249] K. Soundararajan. An inequality for multiplicative functions. Journal of Number Theory, 41(2):225–230, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X92901226.

Young:1992:ANJ

[250] Paul Thomas Young. Apéry numbers, Jacobi sums, and special values of generalized p-adic hypergeometric functions. Journal of Number Theory, 41(2):231–255, June 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X92901237.

Anonymous:1992:PJc

[251] Anonymous. Pages 257–379 (July 1992). Journal of Number Theory, 41 (3):??, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Maohua:1992:DEb

[252] Le Maohua. On the diophantine equation x^{2}−D = 4p^{n}. Journal of Number Theory, 41(3):257–271, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X92901248.

Buchberger:1992:UGB

[253] Bruno Buchberger and Juan Elias. Using Gröbner bases for detecting polynomial identities: a case study on Fermat’s ideal. Journal of Number Theory, 41(3):272–279, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X92901259.

Tamba:1992:NPF

[254] Manvendra Tamba. Note on a partition function. Journal of Number Theory, 41(3):280–282, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290126A.

Childress:1992:FGDa

[255] Nancy Childress and Jeffrey Stopple. Formal groups and Dirichlet L-functions, I. Journal of Number Theory, 41(3):283–294, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9290127B.

Childress:1992:FGDb

[256] Nancy Childress and Jeffrey Stopple. Formal groups and Dirichlet L-functions, II. Journal of Number Theory, 41(3):295–302, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9290128C.

Jia:1992:TBF

[257] Xing-De Jia. Thin bases for finite nilpotent groups. Journal of Number Theory, 41(3):303–313, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290129D.

Keating:1992:AE

[258] Kevin Keating. Automorphisms and extensions of k((t)). Journal of Number Theory, 41(3):314–321, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290130H.

Albu:1992:KEF

[259] Toma Albu. Kummer extensions with few roots of unity. Journal of Number Theory, 41(3):322–358, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X92901318.

Coleman:1992:DPW

[260] M. D. Coleman. The distribution of points at which norm-forms are prime. Journal of Number Theory, 41(3):359–378, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X92901329.

Anonymous:1992:AIVb

[261] Anonymous. Author index for volume 41. Journal of Number Theory, 41(3):379, July 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290133A.

Anonymous:1992:PS

[262] Anonymous. Pages 1–122 (September 1992). Journal of Number Theory, 42(1):??, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1992:EBc

[263] Anonymous. Editorial Board. Journal of Number Theory, 42(1):i, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290103V.

Goldstein:1992:FSR

[264] Sheldon Goldstein, Kevin A. Kelly, and Eugene R. Speer. The fractal structure of rarefied sums of the Thue–Morse sequence. Journal of Number Theory, 42(1):1–19, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9290104W.

Rao:1992:MDE

[265] K. Srinivasa Rao, T. S. Santhanam, and V. Rajeswari. Multiplicative Diophantine equations. Journal of Number Theory, 42(1):20–31, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290105X.

Moreno:1992:IBW

[266] Carlos J. Moreno and Oscar Moreno. An improved Bombieri–Weil bound and applications to coding theory. Journal of Number Theory, 42(1):32–46, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290106Y.

Faure:1992:GPE

[267] Henri Faure. Good permutations for extreme discrepancy. Journal of Number Theory, 42(1):47–56, September 1992. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290107Z.

Grant:1992:SSG

[268] David Grant. Small solutions to a given quadratic form with a variable modulus. Journal of Number Theory, 42(1):57–60, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X92901082.

Nishioka:1992:APC

[269] Kumiko Nishioka, Iekata Shiokawa, and Jun-Ichi Tamura. Arithmetical properties of a certain power series. Journal of Number Theory, 42(1):61–87, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X92901093.

Eswarathasan:1992:DSS

[270] Arulappah Eswarathasan and David Lubell. On the densities of some sequences. Journal of Number Theory, 42(1):88–99, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9290110B.

Aaccagnini:1992:NLG

[271] Alessandro Aaccagnini. A note on large gaps between consecutive primes in arithmetic progressions. Journal of Number Theory, 42(1):100–102, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92901112.

Kolster:1992:RAI

[272] Manfred Kolster. K_{2} of rings of algebraic integers. Journal of Number Theory, 42(1):103–122, September 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X92901123.

Anonymous:1992:PO

[273] Anonymous. Pages 123–246 (October 1992). Journal of Number Theory, 42(2):??, October 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bae:1992:MED

[274] Sunghan Bae. On the modular equation for Drinfeld modules of rank 2. Journal of Number Theory, 42(2):123–133, October 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9290016I.

Choudhry:1992:MCO

[275] Ajai Choudhry. Multigrade chains with odd exponents. Journal of Number Theory, 42(2):134–140, October 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X9290017J.

Debes:1992:IP

[276] Pierre Dèbes. On the irreducibility of the polynomials P (t^{m}, Y). Journal of Number Theory, 42(2):141–157, October 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290018K.

Choie:1992:RPF

[277] Young Ju Choie. Rational period functions, class numbers, and diophantine equations. Journal of Number Theory, 42(2):158–188, October 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X9290019L.

Berend:1992:EQE

[278] Daniel Berend and Charles F. Osgood. On the equation P (x) = n! andaquestionofErdős. JournalofNumberTheory, 42(2) : 189 − −193, October1992.CODENJNUT A9.ISSN0022−314X(print), 1096−

1658(electronic).URLhttpwwwsciencedirectcomsciencearticlepii0022314X9290020P.

Ably:1992:RQI

[279] Mohammed Ably. Résultats quantitatifs d’indépendance algébrique pour les groupes algébriques. (French) [Quantitative results of algebraic independence for algebraic groups]. Journal of Number Theory, 42(2):194–231, October 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290021G.

Spiro:1992:AUS

[280] Claudia A. Spiro. Additive uniqueness sets for arithmetic functions. Journal of Number Theory, 42(2):232–246, October 1992. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290022H.

Anonymous:1992:PN

[281] Anonymous. Pages 247–373 (November 1992). Journal of Number Theory, 42(3):??, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Wolfmann:1992:NSC

[282] Jacques Wolfmann. The number of solutions of certain diagonal equations over finite fields. Journal of Number Theory, 42(3):247–257, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X92900913.

Motohashi:1992:SMV

[283] Yoichi Motohashi. Spectral mean values of Maass waveform L-functions. Journal of Number Theory, 42(3):258–284, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X92900924.

Boston:1992:SCF

[284] Nigel Boston. Some cases of the Fontaine–Mazur conjecture. Journal of Number Theory, 42(3):285–291, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X92900935.

Estes:1992:ESI

[285] Dennis R. Estes. Eigenvalues of symmetric integer matrices. Journal of Number Theory, 42(3):292–296, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X92900946.

Fung:1992:CFF

[286] Gilbert Fung, Andrew Granville, and Hugh C. Williams. Computation of the first factor of the class number of cyclotomic fields. Journal of Number Theory, 42(3):297–312, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X92900957.

Hall:1992:SCE

[287] R. R. Hall. On some conjectures of Erdős in Astérisque, I. Journal of Number Theory, 42(3):313–319, November 1992. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X92900968.

Selmer:1992:ABP

[288] Ernst S. Selmer. Associate bases in the postage stamp problem. Journal of Number Theory, 42(3):320–336, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X92900979.

Leep:1992:HNT

[289] David B. Leep and A. R. Wadsworth. The Hasse norm theorem mod squares. Journal of Number Theory, 42(3):337–348, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X9290098A.

Bond:1992:SHC

[290] Robert J. Bond. On the splitting of the Hilbert class field. Journal of Number Theory, 42(3):349–360, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X9290099B.

Louboutin:1992:NRU

[291] Stéphane Louboutin. Norme relative de l’unité fondamentale de certains corps biquadratiques et parité des longueurs de cycles d’idéaux réduits. (French) [Relative norm of the fundamental unit of certain biquadratic bodies and parity of the lengths of cycles of reductions]. Journal of Number Theory, 42(3):361–368, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X92901004.

Manickam:1992:A

[292] M. Manickam, B. Ramakrishnan, and T. C. Vasudevan. Addendum. Journal of Number Theory, 42(3):369–371, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0022314X9290101T.

Anonymous:1992:AIVc

[293] Anonymous. Author index for volume 42. Journal of Number Theory, 42(3):372–373, November 1992. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X9290102U.

Anonymous:1993:PJa

[294] Anonymous. Pages 1–122 (January 1993). Journal of Number Theory, 43(1):??, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Stevenhagen:1993:DPC

[295] P. Stevenhagen. Divisibility by 2-powers of certain quadratic class numbers. Journal of Number Theory, 43(1):1–19, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710012.

Davis:1993:BCC

[296] K. Davis and W. Webb. A binomial coefficient congruence modulo prime powers. Journal of Number Theory, 43(1):20–23, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710024.

Chapman:1993:LFE

[297] S. T. Chapman and W. W. Smith. On the lengths of factorizations of elements in an algebraic number ring. Journal of Number Theory, 43(1):24–30, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710036.

Freije:1993:KCF

[298] M. N. Freije. Kummer congruences and formal groups. Journal of Number Theory, 43(1):31–42, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710048.

Srivastav:1993:FMM

[299] A. Srivastav and S. Venkataraman. On the Fueter model and monogeneity of rings of integers. Journal of Number Theory, 43(1):43–50, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X8371005X.

Hata:1993:LBR

[300] M. Hata. A lower bound for rational approximations to π. Journal of Number Theory, 43(1):51–67, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710061.

Skoruppa:1993:QCP

[301] N. P. Skoruppa. A quick combinatorial proof of Eisenstein series identities. Journal of Number Theory, 43(1):68–73, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710073.

Muhlbach:1993:CCV

[302] G. Muhlbach. Computation of Cauchy–Vandermonde determinants. Journal of Number Theory, 43(1):74–81, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710085.

Xuan:1993:TKI

[303] T. Z. Xuan. The Turán–Kubilius inequality for integers free of large prime factors. Journal of Number Theory, 43(1):82–87, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710097.

Hancl:1993:CIS

[304] J. Hancl. Criterion for irrational sequences. Journal of Number Theory, 43(1):88–92, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710103.

Hlawka:1993:CAP

[305] E. Hlawka. On a class of approximation polygons for the Peano curve. Journal of Number Theory, 43(1):93–108, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710115.

Gouvea:1993:SFS

[306] F. Q. Gouvea. The square-free sieve over number fields. Journal of Number Theory, 43(1):109–122, January 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710127.

Anonymous:1993:PF

[307] Anonymous. Pages 123–250 (February 1993). Journal of Number Theory, 43(2):??, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Hijikata:1993:RQA

[308] H. Hijikata, H. Saito, and M. Yamauchi. Representations of quaternion algebras over local fields and trace formulas of Hecke operators. Journal of Number Theory, 43(2):123–167, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710139.

Zhang:1993:N

[309] M. Z. Zhang. On Nontotients. Journal of Number Theory, 43(2):168– 172, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710140.

Beeckmans:1993:SAE

[310] L. Beeckmans. The splitting algorithm for Egyptian fractions. Journal of Number Theory, 43(2):173–185, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710152.

Chen:1993:CDA

[311] Y. G. Chen. On a conjecture in diophantine approximations, IV. Journal of Number Theory, 43(2):186–197, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710164.

Cilleruelo:1993:DLP

[312] J. Cilleruelo. The distribution of the lattice points on circles. Journal of Number Theory, 43(2):198–202, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710176.

Denis:1993:BTD

[313] L. Denis. Baker Theorem and Drinfeld modules. Journal of Number Theory, 43(2):203–215, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710188.

Mcquillan:1993:SPI

[314] D. Mcquillan. Split primes and integer-valued polynomials. Journal of Number Theory, 43(2):216–219, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X8371019X.

Soundararajan:1993:PSS

[315] K. Soundararajan. Primes in a sparse sequence. Journal of Number Theory, 43(2):220–227, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710206.

Thunder:1993:NSB

[316] J. L. Thunder. The number of solutions of bounded height to a system of linear equations. Journal of Number Theory, 43(2):228–250, February 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710218.

Anonymous:1993:PMa

[317] Anonymous. Pages 251–379 (March 1993). Journal of Number Theory, 43(3):??, March 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Hayes:1993:HCE

[318] D. R. Hayes. Hecke characters and Eisenstein reciprocity in function fields. Journal of Number Theory, 43(3):251–292, March 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X8371022X.

Borwein:1993:GFI

[319] J. M. Borwein and P. B. Borwein. On the generating function of the integer part: [nα + γ]. Journal of Number Theory, 43(3):293– 318, March 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710231.

Thomas:1993:SCF

[320] E. Thomas. Solutions to certain families of Thue equations. Journal of Number Theory, 43(3):319–369, March 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710243.

Dufner:1993:PSP

[321] G. Dufner. Pjateckij–Shapiro prime twins. Journal of Number Theory, 43(3):370–378, March 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710255.

Anonymous:1993:AIVa

[322] Anonymous. Author index for volume 43. Journal of Number Theory, 43(3):379, March 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710267.

Anonymous:1993:PMb

[323] Anonymous. Pages 1–117 (May 1993). Journal of Number Theory, 44(1): ??, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Nash:1993:SAT

[324] J. C. M. Nash. Some applications of a theorem of M. Kneser. Journal of Number Theory, 44(1):1–8, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710279.

Cucurezeanu:1993:ESL

[325] I. Cucurezeanu. An elementary solution of Lucas’ Problem. Journal of Number Theory, 44(1):9–12, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710280.

Lan:1993:NFU

[326] Y. Z. Lan. On number fields with an unramified abelian extension of degree 2^{n} +2. Journal of Number Theory, 44(1):13–24, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710292.

Brenton:1993:NS

[327] L. Brenton and D. Drucker. On the number of solutions of ∑_{j=1}^{s}(1/x_{j})+ 1/(x_{1} ⋯ x_{s}) = 1. Journal of Number Theory, 44(1):25–29, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X83710309.

Schlickewei:1993:ACN

[328] H. P. Schlickewei and S. A. Stepanov. Algorithms to construct normal bases of cyclic number fields. Journal of Number Theory, 44(1): 30–40, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710310.

Ayala:1993:GMS

[329] E. J. G. Ayala and R. Schertz. Galois module structure in ray class fields over quadratic imaginary number fields. Journal of Number Theory, 44(1):41–46, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710322.

Harman:1993:MTF

[330] G. Harman. Metrical theorems on fractional parts of real sequences, II. Journal of Number Theory, 44(1):47–57, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710334.

Starni:1993:OPN

[331] P. Starni. Odd perfect numbers: a divisor related to the Euler’s factor. Journal of Number Theory, 44(1):58–59, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710346.

Geroldinger:1993:CKL

[332] A. Geroldinger. On a conjecture of Kleitman and Lemke. Journal of Number Theory, 44(1):60–65, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710358.

Berg:1993:GHF

[333] M. C. Berg. On a generalization of Hecke θ-functions and the analytic proof of higher reciprocity laws. Journal of Number Theory, 44(1): 66–83, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X8371036X.

Jia:1993:FSS

[334] X. D. Jia. On finite Sidon sequences. Journal of Number Theory, 44(1): 84–92, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710371.

Erdos:1993:ESM

[335] P. Erdos, C. Pomerance, A. Sarkozy, and C. L. Stewart. On elements of sumsets with many prime factors. Journal of Number Theory, 44(1):93–104, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710383.

Graham:1993:MGB

[336] S. W. Graham. Moments of gaps between k-free numbers. Journal of Number Theory, 44(1):105–117, May 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710395.

Anonymous:1993:PJb

[337] Anonymous. Pages 119–227 (June 1993). Journal of Number Theory, 44 (2):??, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Papadopoulos:1993:NCE

[338] I. Papadopoulos. Neron classification of elliptic curves where the residual characteristics equal 2 or 3. Journal of Number Theory, 44(2):119– 152, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710401.

Scott:1993:E

[339] R. Scott. On the equations p^{x}−b^{y} = c and a^{x}+b^{y} = c^{z}. Journal of Number Theory, 44(2):153–165, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710413.

Halterkoch:1993:QEL

[340] F. Halterkoch. Quadratic equations with a large class number, II. Journal of Number Theory, 44(2):166–171, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710425.

Mignotte:1993:VCT

[341] M. Mignotte. Verification of a conjecture of E. Thomas. Journal of Number Theory, 44(2):172–177, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710437.

Elliott:1993:MFA

[342] P. D. T. A. Elliott. Multiplicative functions on arithmetic progressions. VI. More middle moduli. Journal of Number Theory, 44(2):178– 208, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710449.

Halterkoch:1993:TRC

[343] F. Halterkoch. A theorem of Ramanujan concerning binary quadratic forms. Journal of Number Theory, 44(2):209–213, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710450.

Helou:1993:RJS

[344] C. Helou. Reciprocity for Jacobi sums. Journal of Number Theory, 44(2):214–221, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710462.

Everest:1993:CAU

[345] G. R. Everest and J. H. Loxton. Counting algebraic units with bounded height. Journal of Number Theory, 44(2):222–227, June 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710474.

Anonymous:1993:PJc

[346] Anonymous. Pages 229–368 (July 1993). Journal of Number Theory, 44 (3):??, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Desmit:1993:RGL

[347] B. Desmit. Ramification groups of local fields with imperfect residue class fields. Journal of Number Theory, 44(3):229–236, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710486.

Cilleruelo:1993:ACT

[348] J. Cilleruelo. The additive completion of K-th-powers. Journal of Number Theory, 44(3):237–243, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710498.

Krammer:1993:STS

[349] D. Krammer. Sums of three squares and q-series. Journal of Number Theory, 44(3):244–254, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710504.

Lorenzini:1993:MEC

[350] D. J. Lorenzini. On modular elliptic curves. Journal of Number Theory, 44(3):255–259, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710516.

Cherif:1993:IMC

[351] H. Cherif and B. Demathan. Irrationality measures of Carlitz zeta values in characteristic p. Journal of Number Theory, 44(3):260– 272, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710528.

Sun:1993:SPU

[352] Y. N. Sun. Some properties of uniformly distributed sequences. Journal of Number Theory, 44(3):273–280, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X8371053X.

Ramachandra:1993:STL

[353] K. Ramachandra and A. Sankaranarayanan. On some theorems of Littlewood and Selberg, I. Journal of Number Theory, 44(3):281– 291, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710541.

Taguchi:1993:SSG

[354] Y. Taguchi. Semi-simplicity of the Galois representations attached to Drinfeld modules over fields of “infinite characteristics”. Journal of Number Theory, 44(3):292–314, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710553. See correction [1133].

Drmota:1993:LDE

[355] M. Drmota. On linear diophantine equations and Fibonacci numbers. Journal of Number Theory, 44(3):315–327, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710565.

Friedman:1993:SDE

[356] C. N. Friedman. Sums of divisors and Egyptian fractions. Journal of Number Theory, 44(3):328–339, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710577.

Childress:1993:AMF

[357] N. Childress. p-adic measures and formal groups. Journal of Number Theory, 44(3):340–351, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710589.

Kisilevsky:1993:MIF

[358] H. Kisilevsky. Multiplicative independence in function fields. Journal of Number Theory, 44(3):352–355, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710590.

James:1993:PRU

[359] D. G. James. Primitive representations by unimodular quadratic forms. Journal of Number Theory, 44(3):356–366, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710607.

Anonymous:1993:AIVb

[360] Anonymous. Author index for volume 44. Journal of Number Theory, 44(3):367–368, July 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710619.

Anonymous:1993:PS

[361] Anonymous. Pages 1–116 (September 1993). Journal of Number Theory, 45(1):??, September 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Lhermitte:1993:MOM

[362] M. P. Lhermitte. Minimal order morphisms. Journal of Number Theory, 45(1):1–27, September 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710620.

Delcorso:1993:CAD

[363] I. Delcorso and R. Dvornicich. A converse of Artin’s density theorem: the case of cubic fields. Journal of Number Theory, 45(1):28–44, September 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710632.

Mai:1993:AAR

[364] L. Mai. The average analytic rank of a family of elliptic curves. Journal of Number Theory, 45(1):45–60, September 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710644.

Kaplan:1993:RPA

[365] P. Kaplan and K. S. Williams. Representation of primes in arithmetic progression by binary quadratic forms. Journal of Number Theory, 45(1):61–67, September 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710656.

Ashford:1993:DRQ

[366] S. R. Ashford. Dyadic ramification and quartic number fields. Journal of Number Theory, 45(1):68–91, September 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710668.

Habsieger:1993:AGR

[367] L. Habsieger. Applications of group representation theory to the easier Waring problem. Journal of Number Theory, 45(1):92–111, September 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X8371067X.

Desmit:1993:ANI

[368] B. Desmit. Algebraic numbers with integral power traces. Journal of Number Theory, 45(1):112–116, September 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710681.

Anonymous:1993:PO

[369] Anonymous. Pages 117–240 (October 1993). Journal of Number Theory, 45(2):??, October 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Ugrinsparac:1993:CEP

[370] D. Ugrinsparac. On a class of enumeration problems in additive arithmetics. Journal of Number Theory, 45(2):117–128, October 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X83710693.

Diamond:1993:BVP

[371] H. G. Diamond, H. Halberstam, and H. E. Richert. A boundary value problem for a pair of differential delay equations related to sieve theory, II. Journal of Number Theory, 45(2):129–185, October 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X8371070X.

Warlimont:1993:FNC

[372] R. Warlimont. Factorisatio Numerorum with Constraints. Journal of Number Theory, 45(2):186–199, October 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710711.

Akiyama:1993:NTI

[373] S. Akiyama. A new type of inclusion exclusion principle for sequences and asymptotic formulas for ζ(k). Journal of Number Theory, 45(2):200– 214, October 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710723.

Matalaaho:1993:DAR

[374] T. Matalaaho. On diophantine approximations of the Rogers–Ramanujan continued fraction. Journal of Number Theory, 45(2):215–227, October 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710735.

Borwein:1993:SRP

[375] P. Borwein. Some restricted partition functions. Journal of Number Theory, 45(2):228–240, October 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710747.

Anonymous:1993:PN

[376] Anonymous. Pages 241–377 (November 1993). Journal of Number Theory, 45(3):??, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jia:1993:DMO

[377] C. H. Jia. On the distribution of ap modulo one. Journal of Number Theory, 45(3):241–253, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710759.

Costa:1993:GRB

[378] A. Costa. A generalization of a result of Barrucand and Cohn on class numbers. Journal of Number Theory, 45(3):254–260, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X83710760.

Wang:1993:SSC

[379] Y. Wang. Small Solutions of Congruences. Journal of Number Theory, 45(3):261–280, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710772.

Kurihara:1993:ICG

[380] M. Kurihara. Ideal class groups of cyclotomic fields and modular forms of level 1. Journal of Number Theory, 45(3):281–294, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X83710784.

Tan:1993:MEF

[381] K. S. Tan. Modular elements over function fields. Journal of Number Theory, 45(3):295–311, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X83710796.

Kiuchi:1993:IMV

[382] I. Kiuchi. An improvement on the mean value formula for the approximate functional equation of the square of the Riemann zeta-function. Journal of Number Theory, 45(3):312–319, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X83710802.

Sankaranarayanan:1993:MVT

[383] A. Sankaranarayanan and K. Srinivas. Mean-value theorem of the Riemann zeta-function over short intervals. Journal of Number Theory, 45(3):320–326, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X83710814.

Haland:1993:UDG

[384] I. J. Haland. Uniform distribution of generalized polynomials. Journal of Number Theory, 45(3):327–366, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X83710826.

Bourque:1993:MAC

[385] K. Bourque and S. Ligh. Matrices associated with classes of arithmetical functions. Journal of Number Theory, 45(3):367–376, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X83710838.

Anonymous:1993:AIVc

[386] Anonymous. Author index for volume 45. Journal of Number Theory, 45(3):377, November 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X8371084X.

Anonymous:1994:PJa

[387] Anonymous. Pages 1–122 (January 1994). Journal of Number Theory, 46(1):??, January 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Ding:1994:PCC

[388] J. T. Ding. A proof of a conjecture of C. L. Siegel. Journal of Number Theory, 46(1):1–11, January 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710018.

Grytczuk:1994:EFD

[389] A. Grytczuk. Explicit formulas of Delange–Selberg type for some classes of arithmetic functions on the semigroup G_{K}. Journal of Number Theory, 46(1):12–28, January 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X8471002X.

Nathanson:1994:ITS

[390] M. B. Nathanson. An inverse theorem for sums of sets of lattice points. Journal of Number Theory, 46(1):29–59, January 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710031.

Odai:1994:IGG

[391] Y. Odai. On the index of the group generated by relative units. Journal of Number Theory, 46(1):60–69, January 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710043.

Kohnen:1994:NPT

[392] W. Kohnen. Nonholomorphic Poincaré-type series on Jacobi groups. Journal of Number Theory, 46(1):70–99, January 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710055.

Deshalit:1994:NSS

[393] E. Deshalit. A note on the Shimura subgroup of J_{0}(p). Journal of Number Theory, 46(1):100–107, January 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710067.

Hastad:1994:DMT

[394] J. Hastad, J. C. Lagarias, and A. M. Odlyzko. On the distribution of multiplicative translates of sets of residues (modp). Journal of Number Theory, 46(1):108–122, January 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710079.

Anonymous:1994:PF

[395] Anonymous. Pages 123–254 (February 1994). Journal of Number Theory, 46(2):??, February 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Sander:1994:IHD

[396] J. W. Sander. On 4/n = 1/x + 1/y + 1/z and Iwaniec’ half dimensional sieve. Journal of Number Theory, 46(2):123–136, February 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710080.

Kraaikamp:1994:SAD

[397] C. Kraaikamp. On symmetrical and asymmetric diophantine approximation by continued fractions. Journal of Number Theory, 46(2):137– 157, February 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710092.

Solomon:1994:GRC

[398] D. Solomon. Galois relations for cyclotomic numbers and p-units. Journal of Number Theory, 46(2):158–178, February 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710109.

Rausch:1994:CSA

[399] U. Rausch. Character sums in algebraic number fields. Journal of Number Theory, 46(2):179–195, February 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710110.

Rote:1994:SSC

[400] G. Rote. Sequences with subword complexity 2n. Journal of Number Theory, 46(2):196–213, February 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710122.

Soleng:1994:HGR

[401] R. Soleng. Homomorphisms from the group of rational points on elliptic curves to class groups of quadratic number fields. Journal of Number Theory, 46(2):214–229, February 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710134.

Kani:1994:IRA

[402] E. Kani and M. Rosen. Idempotent relations among arithmetic invariants attached to number fields and algebraic varieties. Journal of Number Theory, 46(2):230–254, February 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710146.

Anonymous:1994:PMa

[403] Anonymous. Pages 255–409 (March 1994). Journal of Number Theory, 46(3):??, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kessler:1994:CPO

[404] V. Kessler. Crossed product orders and non-commutative arithmetic. Journal of Number Theory, 46(3):255–302, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710158.

Skalba:1994:GTT

[405] M. Skalba. Generalization of Thue’s Theorem and computation of the group K_{2}O_{F}. Journal of Number Theory, 46(3):303–322, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X8471016X.

Gordon:1994:EAC

[406] D. M. Gordon. Equidistant arithmetic codes and character sums. Journal of Number Theory, 46(3):323–333, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710171.

Bley:1994:CIB

[407] W. Bley. Construction of integral bases in abelian field-extensions of imaginary quadratic number-fields. Journal of Number Theory, 46(3): 334–371, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710183.

Sander:1994:PPD

[408] J. W. Sander. Prime power divisors of multinomial coefficients and Artin’s Conjecture. Journal of Number Theory, 46(3):372–384, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710195.

Katayama:1994:BFU

[409] S. I. Katayama and S. Katayama. On bounds for fundamental units of real quadratic fields. Journal of Number Theory, 46(3):385–390, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710201.

Godinho:1994:AFD

[410] H. Godinho. Additive forms of degree 2^{l}. Journal of Number Theory, 46(3):391–408, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710213.

Anonymous:1994:AIVa

[411] Anonymous. Author index for volume 46. Journal of Number Theory, 46(3):409, March 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710225.

Anonymous:1994:PAa

[412] Anonymous. Pages 1–129 (April 1994). Journal of Number Theory, 47(1): ??, April 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Pohst:1994:MHZ

[413] M. Pohst. In memoriam: Hans Zassenhaus (1912–1991). Journal of Number Theory, 47(1):1–19, April 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710237.

Rothstein:1994:DAA

[414] M. Rothstein and H. Zassenhaus. Deterministic analysis of aleatoric methods of polynomial factorization over finite fields. Journal of Number Theory, 47(1):20–42, April 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710249.

Mignotte:1994:ANS

[415] M. Mignotte and M. Waldschmidt. On algebraic numbers of small height: linear forms in one logarithm. Journal of Number Theory, 47(1):43– 62, April 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710250.

Gyory:1994:CPR

[416] K. Gyory and A. Schinzel. On a conjecture of Posner and Rumsey. Journal of Number Theory, 47(1):63–78, April 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710262.

Bannai:1994:MIC

[417] E. Bannai and E. Bannai. Modular invariance of the character table of the Hamming association scheme H(d, q). Journal of Number Theory, 47(1):79–92, April 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710274.

Pohst:1994:CFU

[418] M. Pohst. On computing fundamental units. Journal of Number Theory, 47(1):93–105, April 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710286.

Yui:1994:NNA

[419] N. Yui. A note on the norms of algebraic numbers associated to Jacobi sums. Journal of Number Theory, 47(1):106–129, April 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710298.

Anonymous:1994:PMb

[420] Anonymous. Pages 131–245 (May 1994). Journal of Number Theory, 47 (2):??, May 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Glass:1994:SCC

[421] A. M. W. Glass, D. B. Meronk, T. Okada, and R. P. Steiner. A small contribution to Catalan’s equation. Journal of Number Theory, 47(2):131–137, May 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710304.

Elder:1994:GMS

[422] G. G. Elder and M. L. Madan. Galois module structure of the integers in wildly ramified cyclic extensions. Journal of Number Theory, 47(2):138–174, May 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710316.

Hsia:1994:RPD

[423] J. S. Hsia and J. P. Prietocox. Representations of positive definite Hermitian forms with approximation and primitive properties. Journal of Number Theory, 47(2):175–189, May 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710328.

Dumir:1994:VNH

[424] V. C. Dumir, R. J. Hansgill, and A. C. Woods. Values of nonhomogeneous indefinite quadratic forms. Journal of Number Theory, 47(2):190–197, May 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X8471033X.

Harada:1994:SCN

[425] K. Harada, M. L. Lang, and M. Miyamoto. Sequential construction of Niemeier lattices and their uniqueness. Journal of Number Theory, 47(2):198–223, May 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710341.

Flicker:1994:PTS

[426] Y. Z. Flicker and J. G. M. Mars. Products of theta series and spectral analysis. Journal of Number Theory, 47(2):224–245, May 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710353.

Anonymous:1994:PJb

[427] Anonymous. Pages 247–436 (June 1994). Journal of Number Theory, 47 (3):??, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Raychaudhuri:1994:LSD

[428] D. K. Raychaudhuri and E. J. Schram. A large set of designs on vector spaces. Journal of Number Theory, 47(3):247–272, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710365.

Plesken:1994:AIP

[429] W. Plesken. Additively indecomposable positive integral quadratic forms. Journal of Number Theory, 47(3):273–283, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710377.

Bovdi:1994:TUI

[430] A. Bovdi, Z. Marciniak, and S. K. Sehgal. Torsion units in infinite group rings. Journal of Number Theory, 47(3):284–299, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710389.

Diamond:1994:BVP

[431] H. G. Diamond, H. Halberstam, and H. E. Richert. A boundary value problem for a pair of differential delay equations related to sieve theory. III. Journal of Number Theory, 47(3):300–328, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710390.

Erdos:1994:SSS

[432] P. Erdos, A. Sarkozy, and T. Sos. On sum sets of Sidon sets, 1. Journal of Number Theory, 47(3):329–347, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710407.

Schmidt:1994:E

[433] W. M. Schmidt. Equations a^{x} = R(x, y). Journal of Number Theory, 47(3):348–358, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710419.

Queen:1994:ELC

[434] C. S. Queen. Euclidean-like characterizations of Dedekind, Krull, and factorial domains. Journal of Number Theory, 47(3):359–370, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710420.

Lambek:1994:SGC

[435] J. Lambek. Some Galois connections in elementary number theory. Journal of Number Theory, 47(3):371–377, June 1994. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710432.

Vigneras:1994:BCR

[436] M. F. Vigneras. Banal characteristic for reductive p-adic groups. Journal of Number Theory, 47(3):378–397, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710444.

Sonn:1994:REP

[437] J. Sonn. Rigidity and embedding problems over Q_{ab}(t). Journal of Number Theory, 47(3):398–404, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710456.

Moody:1994:SQ

[438] R. V. Moody and A. Weiss. On Shelling E_{8} quasicrystals. Journal of Number Theory, 47(3):405–412, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710468.

Wang:1994:PS

[439] X. Y. Wang and A. Weiss. Permutation summands over Z. Journal of Number Theory, 47(3):413–434, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X8471047X.

Anonymous:1994:AIVb

[440] Anonymous. Author index for volume 47. Journal of Number Theory, 47(3):435–436, June 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710481.

Anonymous:1994:PJc

[441] Anonymous. Pages 1–111 (July 1994). Journal of Number Theory, 48(1): ??, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Sehmi:1994:PVN

[442] R. Sehmi and V. C. Dumir. Positive values of non-homogeneous indefinite quadratic forms of type (2, 5). Journal of Number Theory, 48(1):

1–35, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710493.

Jakubec:1994:CGP

[443] S. Jakubec. The congruence for Gauss period. Journal of Number Theory, 48(1):36–45, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X8471050X.

Oukhaba:1994:DFA

[444] H. Oukhaba. On discriminant functions associated to Drinfeld modules of rank 1. Journal of Number Theory, 48(1):46–61, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710511.

Chen:1994:NSD

[445] J. H. Chen. A new solution of the diophantine equation X^{2} + 1 = 2Y^{4}. Journal of Number Theory, 48(1):62–74, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710523.

Rynne:1994:HDS

[446] B. P. Rynne. The Hausdorff dimension of sets of points whose simultaneous rational approximation by sequences of integer vectors have errors with small product. Journal of Number Theory, 48(1):75– 79, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710535.

Myerson:1994:WLC

[447] G. Myerson. What the least common multiple divides. Journal of Number Theory, 48(1):80–87, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710547.

Jochner:1994:RPD

[448] M. Jochner and Y. Kitaoka. Representations of positive definite quadratic forms with congruence and primitive conditions. Journal of Number Theory, 48(1):88–101, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710559.

Lang:1994:VZF

[449] H. Lang. The values of the zeta-function of a class of ideals and the Ankeny, Artin and Chowla congruences for the class number of real quadratic number-fields. Journal of Number Theory, 48(1):102– 108, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710560.

Rousseau:1994:JS

[450] G. Rousseau. On the Jacobi symbol. Journal of Number Theory, 48(1): 109–111, July 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710572.

Anonymous:1994:PAb

[451] Anonymous. Pages 113–257 (August 1994). Journal of Number Theory, 48(2):??, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Markett:1994:TSR

[452] C. Markett. Triple sums and the Riemann zeta function. Journal of Number Theory, 48(2):113–132, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710584.

Shu:1994:CNF

[453] L. S. Shu. Class number formulas over global function fields. Journal of Number Theory, 48(2):133–161, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710596.

Sander:1994:PP

[454] J. W. Sander. A partition problem. Journal of Number Theory, 48(2):162–182, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710602.

Jia:1994:S

[455] X. D. Jia. On B_{2k}-sequences. Journal of Number Theory, 48(2):183– 196, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710614. See [488, 655].

Sarkozy:1994:FAT

[456] A. Sarkozy. Fine addition theorems, II. Journal of Number Theory, 48(2):197–218, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710626.

Geroldinger:1994:CNF

[457] A. Geroldinger. The cross number of finite abelian groups. Journal of Number Theory, 48(2):219–223, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710638.

Berndt:1994:NCT

[458] B. C. Berndt and L. C. Zhang. A new class of theta-function identities originating in Ramanujan’s notebooks. Journal of Number Theory, 48(2):224–242, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X8471064X.

Walling:1994:RDT

[459] L. H. Walling. A remark on differences of theta series. Journal of Number Theory, 48(2):243–251, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710651.

Ibukiyama:1994:FTZ

[460] T. Ibukiyama and H. Saito. On L-functions of ternary zero forms and exponential sums of Lee and Weintraub. Journal of Number Theory, 48(2):252–257, August 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710663.

Anonymous:1994:PS

[461] Anonymous. Pages 259–413 (September 1994). Journal of Number Theory, 48(3):??, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Cusick:1994:IDA

[462] T. W. Cusick, A. M. Rockett, and P. Szusz. On inhomogeneous Diophantine approximation. Journal of Number Theory, 48(3):259–283, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710675.

Wang:1994:JSF

[463] J. Wang. On the Jacobi sums for finite commutative rings with identity. Journal of Number Theory, 48(3):284–290, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710687.

Silverman:1994:VCHa

[464] J. H. Silverman. Variation of the canonical height on elliptic-surfaces II: local analyticity properties. Journal of Number Theory, 48(3):291–329, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710699.

Silverman:1994:VCHb

[465] J. H. Silverman. Variation of the canonical height on elliptic surfaces III: global boundedness properties. Journal of Number Theory, 48(3):330–352, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710705.

Davison:1994:LDP

[466] J. L. Davison. On the linear diophantine problem of Frobenius. Journal of Number Theory, 48(3):353–363, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710717.

Rangachari:1994:RRT

[467] S. S. Rangachari. On a result of Ramanujan on theta functions. Journal of Number Theory, 48(3):364–372, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710729.

Conway:1994:LET

[468] J. H. Conway and N. J. A. Sloane. On lattices equivalent to their duals. Journal of Number Theory, 48(3):373–382, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710730.

Tsumura:1994:EDS

[469] H. Tsumura. On evaluation of the Dirichlet series at positive integers by q-calculation. Journal of Number Theory, 48(3):383–391, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710742.

Wolfskill:1994:RAN

[470] J. Wolfskill. Reduced algebraic numbers in the complex plane. Journal of Number Theory, 48(3):392–412, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710754.

Anonymous:1994:AIVc

[471] Anonymous. Author index for volume 48. Journal of Number Theory, 48(3):413, September 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710766.

Anonymous:1994:PO

[472] Anonymous. Pages 1–132 (October 1994). Journal of Number Theory, 49(1):??, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Topfer:1994:ANM

[473] T. Topfer. An axiomatization of Nesterenko’s method and applications on Mahler functions. Journal of Number Theory, 49(1):1–26, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710778.

Baragar:1994:ISM

[474] A. Baragar. Integral solutions of Markov–Hurwitz equations. Journal of Number Theory, 49(1):27–44, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X8471078X.

Comalada:1994:TRE

[475] S. Comalada. Twists and reduction of an elliptic curve. Journal of Number Theory, 49(1):45–62, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710791.

Yang:1994:HAQ

[476] J. H. Yang. Harmonic analysis on the quotient spaces of Heisenberg groups, II. Journal of Number Theory, 49(1):63–72, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X84710808.

Heathbrown:1994:SC

[477] D. R. Heathbrown and K. Tsang. Sign changes of E(T), Δ(x), and P (x). Journal of Number Theory, 49(1):73–83, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X8471081X.

Ruzsa:1994:IET

[478] I. Z. Ruzsa. On an inequality of Erdős and Turán concerning uniform distribution modulo one, II. Journal of Number Theory, 49(1): 84–88, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710821.

Tong:1994:BAF

[479] J. C. Tong. The best approximation function to irrational numbers. Journal of Number Theory, 49(1):89–94, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710833.

Chung:1994:SGW

[480] F. R. K. Chung. Several generalizations of Weil sums. Journal of Number Theory, 49(1):95–106, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710845.

Helou:1994:PRC

[481] C. Helou. Power residue character of Jacobi sums. Journal of Number Theory, 49(1):107–117, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X84710857.

Gao:1994:DPI

[482] S. H. Gao and G. L. Mullen. Dickson polynomials and irreducible polynomials over finite fields. Journal of Number Theory, 49(1):118– 132, October 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710869.

Anonymous:1994:PN

[483] Anonymous. Pages 133–268 (November 1994). Journal of Number Theory, 49(2):??, November 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Louboutin:1994:ECG

[484] S. Louboutin. The exponent 2-class-group problem for non–Galois over Q quartic fields that are quadratic extensions of imaginary quadratic fields. Journal of Number Theory, 49(2):133–141, November 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710870.

Hensley:1994:NSE

[485] D. Hensley. The number of steps in the Euclidean Algorithm. Journal of Number Theory, 49(2):142–182, November 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710882.

Morton:1994:CCC

[486] P. Morton. Characterizing cyclic cubic extensions by automorphism polynomials. Journal of Number Theory, 49(2):183–208, November 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X84710894.

Farmer:1994:MVD

[487] D. W. Farmer. Mean value of Dirichlet series associated with holomorphic cusp forms. Journal of Number Theory, 49(2):209–245, November 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X84710900.

Helm:1994:RS

[488] M. Helm. A remark on B_{2k}-sequences. Journal of Number Theory, 49(2):246–249, November 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710912. See [455, 655].

Filaseta:1994:PVB

[489] M. Filaseta. Powerfree values of binary forms. Journal of Number Theory, 49(2):250–268, November 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84710924.

Anonymous:1994:PD

[490] Anonymous. Pages 269–392 (December 1994). Journal of Number Theory, 49(3):??, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Becker:1994:RPS

[491] P. G. Becker. k-regular power series and Mahler-type functional equations. Journal of Number Theory, 49(3):269–286, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X84710936.

Yu:1994:TIV

[492] K. T. Yu. On a trigonometric inequality of Vinogradov. Journal of Number Theory, 49(3):287–294, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710948.

Yanai:1994:DCT

[493] H. Yanai. On degenerate CM-types. Journal of Number Theory, 49 (3):295–303, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X8471095X.

Yamaguchi:1994:UAN

[494] I. Yamaguchi and H. Takeuchi. On the units of algebraic number fields. Journal of Number Theory, 49(3):304–307, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X84710961.

Lacroix:1994:NSR

[495] Y. Lacroix and A. Thomas. Number Systems and Repartition. Journal of Number Theory, 49(3):308–318, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710973.

Shu:1994:KCG

[496] L. H. Shu. Kummer’s criterion over global function fields. Journal of Number Theory, 49(3):319–359, December 1994. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710985.

Jager:1994:CDS

[497] H. Jager and J. Dejonge. The circular dispersion spectrum. Journal of Number Theory, 49(3):360–384, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X84710997.

Kozma:1994:BDF

[498] G. Kozma and A. Lev. On H-bases and H-decompositions of the finite solvable and alternating groups. Journal of Number Theory, 49 (3):385–391, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84711000.

Anonymous:1994:AIVd

[499] Anonymous. Author index for volume 48. Journal of Number Theory, 49(3):392, December 1994. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X84711012.

Anonymous:1995:PJa

[500] Anonymous. Pages 1–179 (January 1995). Journal of Number Theory, 50(1):??, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Guralnick:1995:SGE

[501] R. M. Guralnick and L. Stern. Solitary Galois extensions of algebraic number fields. Journal of Number Theory, 50(1):1–32, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710013.

Naranjani:1995:SRA

[502] A. M. Naranjani. Sequences of residues in algebraic number fields. Journal of Number Theory, 50(1):33–36, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710025.

Kim:1995:ES

[503] D. S. Kim. On Eisenstein series of GL_{n}. Journal of Number Theory, 50(1):37–65, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710037.

Fraenkel:1995:GPI

[504] A. S. Fraenkel and R. Holzman. Gap problems for integer part and fractional part sequences. Journal of Number Theory, 50(1):66–86, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710049.

Miki:1995:CCH

[505] H. Miki. On the calculation of certain Hilbert norm residue symbols and its application. Journal of Number Theory, 50(1):87–105, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710050.

Kaczorowski:1995:SRR

[506] J. Kaczorowski. On the Shanks–Rényi race problem mod5. Journal of Number Theory, 50(1):106–118, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710062.

Grundman:1995:SFU

[507] H. G. Grundman. Systems of fundamental units in cubic orders. Journal of Number Theory, 50(1):119–127, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710074.

Bowman:1995:AZ

[508] D. Bowman. Approximation of ⌊nα + s⌋ and the zero of nα + s. Journal of Number Theory, 50(1):128–144, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710086.

Jochner:1995:RPD

[509] M. Jochner. Representations of positive definite quadratic forms with congruence and primitive conditions, II. Journal of Number Theory, 50(1):145–153, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710098.

Sime:1995:HCF

[510] P. J. Sime. Hilbert class fields of real biquadratic fields. Journal of Number Theory, 50(1):154–166, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85710104.

Geijsberts:1995:GTK

[511] M. Geijsberts. On the generation of the tame kernel by Dennis– Stein symbols. Journal of Number Theory, 50(1):167–179, January 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85710116.

Anonymous:1995:PF

[512] Anonymous. Pages 181–377 (February 1995). Journal of Number Theory, 50(2):??, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Brown:1995:SFP

[513] T. C. Brown and P. J. S. Shiue. Sums of fractional parts of integer multiples of an irrational. Journal of Number Theory, 50(2):181–192, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710128.

Le:1995:NGR

[514] M. Le. A note on the generalized Ramanujan–Nagell equation. Journal of Number Theory, 50(2):193–201, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X8571013X.

Iovita:1995:CRV

[515] A. Iovita and A. Zaharescu. Completions of r.a.t.-valued fields of rational functions. Journal of Number Theory, 50(2):202–205, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85710141.

Hajir:1995:SEO

[516] F. Hajir. Some Ã_{n}-extensions obtained from generalized Laguerre polynomials. Journal of Number Theory, 50(2):206–212, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85710153.

Drmota:1995:UDM

[517] M. Drmota and R. Winkler. s(N) — uniform distribution modulo 1. Journal of Number Theory, 50(2):213–225, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710165.

Folk:1995:ECC

[518] D. Folk. Extra capitulation and central extensions. Journal of Number Theory, 50(2):226–232, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85710177.

Everest:1995:PAU

[519] G. R. Everest. On the proximity of algebraic units to divisors. Journal of Number Theory, 50(2):233–250, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710189.

Berger:1995:ICG

[520] R. I. Berger. Ideal class groups, some computations. Journal of Number Theory, 50(2):251–260, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85710190.

Baker:1995:PF

[521] R. C. Baker, G. Harman, and J. Rivat. Primes of the form [n^{c}]. Journal of Number Theory, 50(2):261–277, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710207.

Bremner:1995:E

[522] A. Bremner and J. W. Jones. On the equation x^{4} + mx^{2}y^{2} + y^{4} = z^{2}. Journal of Number Theory, 50(2):268–298, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710219.

Hazama:1995:RPC

[523] F. Hazama. Rational points on certain abelian varieties over function fields. Journal of Number Theory, 50(2):278–285, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85710220.

Cochrane:1995:SZQ

[524] T. Cochrane. Small zeros of quadratic congruences modulo pq, II. Journal of Number Theory, 50(2):299–308, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710232.

Bertin:1995:CTR

[525] M. J. Bertin and D. W. Boyd. A characterization of two related classes of Salem numbers. Journal of Number Theory, 50(2):309–317, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710244.

Zhu:1995:CPD

[526] F. Z. Zhu. On the construction of positive definite indecomposable unimodular even Hermitian forms. Journal of Number Theory, 50 (2):318–322, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710256.

Hutchinson:1995:NGG

[527] K. Hutchinson. Norm groups of global fields. Journal of Number Theory, 50(2):323–328, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710268.

Almkvist:1995:CHR

[528] G. Almkvist and H. S. Wilf. On the coefficients in the Hardy– Ramanujan–Rademacher formula for p(n). Journal of Number Theory, 50(2):329–334, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X8571027X.

Buck:1995:CFE

[529] M. W. Buck and D. P. Robbins. The continued fraction expansion of an algebraic power series satisfying a quartic equation. Journal of Number Theory, 50(2):335–344, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85710281.

Merel:1995:HOS

[530] L. Merel. Hecke operators and Γ(2) subgroups. Journal of Number Theory, 50(2):345–362, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710293.

Kapranov:1995:HDG

[531] M. Kapranov. A higher-dimensional generalization of the Goss zeta function. Journal of Number Theory, 50(2):363–375, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X8571030X.

Anonymous:1995:AIVa

[532] Anonymous. Author index for volume 50. Journal of Number Theory, 50(2):376–377, February 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710311.

Anonymous:1995:PMa

[533] Anonymous. Pages 1–146 (March 1995). Journal of Number Theory, 51 (1):??, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Zhang:1995:MVM

[534] W. B. Zhang. On the mean value of multiplicative functions. Journal of Number Theory, 51(1):1–15, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85710323.

Ivic:1995:FPM

[535] A. Ivic and Y. Motohashi. On the fourth power moment of the Riemann zeta-function. Journal of Number Theory, 51(1):16–45, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85710335.

Narkiewicz:1995:NEF

[536] W. Narkiewicz. A note on elasticity of factorizations. Journal of Number Theory, 51(1):46–47, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710347.

Muller:1995:RSM

[537] W. Muller. The Rankin–Selberg method for non-holomorphic automorphic forms. Journal of Number Theory, 51(1):48–86, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85710359.

Huard:1995:IBQ

[538] J. G. Huard, B. K. Spearman, and K. S. Williams. Integral bases for quartic fields with quadratic subfields. Journal of Number Theory, 51(1):87–102, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710360.

Jones:1995:IFE

[539] J. W. Jones. Iwasawa L-functions of elliptic curves with additive reduction. Journal of Number Theory, 51(1):103–117, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710372.

Morales:1995:TFS

[540] J. Morales. Trace forms and Stickelberger relations. Journal of Number Theory, 51(1):118–129, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710384.

Habsieger:1995:ACP

[541] L. Habsieger. On the additive completion of polynomial sets. Journal of Number Theory, 51(1):130–135, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710396.

Merklen:1995:WMA

[542] H. A. Merklen. Web modules and applications. Journal of Number Theory, 51(1):136–146, March 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710402.

Anonymous:1995:PAa

[543] Anonymous. Pages 147–306 (April 1995). Journal of Number Theory, 51(2):??, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Balakrishnan:1995:SDF

[544] U. Balakrishnan. On the sum of divisors function. Journal of Number Theory, 51(2):147–168, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710414.

Katsurada:1995:CFP

[545] H. Katsurada. A certain formal power series of several variables attached to local densities of quadratic forms, I. Journal of Number Theory, 51(2):169–209, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710426.

Arenas:1995:GRE

[546] A. Arenas. Genera and rational equivalence classes of integral quadratic forms. Journal of Number Theory, 51(2):210–218, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710438.

Schaefer:1995:DJH

[547] E. F. Schaefer. 2-descent on the Jacobians of hyperelliptic curves. Journal of Number Theory, 51(2):219–232, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X8571044X.

Kogiso:1995:SCR

[548] T. Kogiso. Simple calculation of the residues of the adelic zeta function associated with the space of binary cubic forms. Journal of Number Theory, 51(2):233–248, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710451.

Car:1995:QF

[549] M. Car. Quadratic forms on F_{2^{h}}[T]. Journal of Number Theory, 51(2): 249–274, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710463.

Ye:1995:LKS

[550] Y. B. Ye. The lifting of Kloosterman sums. Journal of Number Theory, 51(2):275–287, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710475.

Liverance:1995:FRN

[551] E. Liverance. A formula for the root number of a family of elliptic curves. Journal of Number Theory, 51(2):288–305, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710487.

Anonymous:1995:AIVb

[552] Anonymous. Author index for volume 51. Journal of Number Theory, 51 (2):306, April 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710499.

Anonymous:1995:PMb

[553] Anonymous. Pages 1–172 (May 1995). Journal of Number Theory, 52(1): ??, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Zhang:1995:DBI

[554] W. P. Zhang. On the difference between an integer and its inverse modulo n. Journal of Number Theory, 52(1):1–6, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710505.

Thong:1995:SNI

[555] N. Q. D. Thong and H. Opolka. Some numerical invariants related to central embedding problems. Journal of Number Theory, 52(1):7– 16, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710517.

Imamoglu:1995:KS

[556] O. Imamoglu. The Kubota symbol for Sp(2n, Q(i)). Journal of Number Theory, 52(1):17–34, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710529.

Jacroux:1995:CSI

[557] M. Jacroux. On the construction of sets of integers with equal power sums. Journal of Number Theory, 52(1):35–42, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710530.

Kucera:1995:PCN

[558] R. Kucera. On the parity of the class number of a biquadratic field. Journal of Number Theory, 52(1):43–52, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710542.

Schweizer:1995:DMP

[559] A. Schweizer. On the Drinfeld modular polynomial Φ_{T} (X, Y). Journal of Number Theory, 52(1):53–68, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85710554.

Clarke:1995:HLD

[560] F. Clarke. Hensel’s Lemma and the divisibility by primes of Stirling-like numbers. Journal of Number Theory, 52(1):69–84, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710566.

Sands:1995:DMA

[561] J. W. Sands and W. Schwarz. A Demjanenko matrix for abelian fields of prime power conductor. Journal of Number Theory, 52(1): 85–97, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710578.

Popescu:1995:SIP

[562] N. Popescu and A. Zaharescu. On the structure of the irreducible polynomials over local fields. Journal of Number Theory, 52(1):98– 118, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X8571058X.

Schultheis:1995:LGR

[563] F. Schultheis. Local and global residue symbols for algebraic functions fields. Journal of Number Theory, 52(1):119–124, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710591.

Young:1995:JSM

[564] P. T. Young. On Jacobi sums, multinomial coefficients, and padic hypergeometric functions. Journal of Number Theory, 52(1):125– 144, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710608.

Sander:1995:ICM

[565] J. W. Sander. Irrationality criteria for Mahler’s numbers. Journal of Number Theory, 52(1):145–156, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X8571061X.

Howard:1995:ARB

[566] F. T. Howard. Applications of a recurrence for the Bernoulli numbers. Journal of Number Theory, 52(1):157–172, May 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710621.

Anonymous:1995:PJb

[567] Anonymous. Pages 173–329 (June 1995). Journal of Number Theory, 52 (2):??, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Brocco:1995:NMG

[568] S. Brocco. A note on Mignosi’s generalization of the (3X + 1)-problem. Journal of Number Theory, 52(2):173–178, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710633.

Adler:1995:AFH

[569] A. Adler and L. C. Washington. p-adic L-functions and higherdimensional magic cubes. Journal of Number Theory, 52(2):179– 197, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710645.

Larcher:1995:DAC

[570] G. Larcher. On the distribution of an analog to classical Kroneckersequences. Journal of Number Theory, 52(2):198–215, June 1995. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0022314X85710657.

Bley:1995:GMS

[571] W. Bley. Galois module structure and elliptic functions. Journal of Number Theory, 52(2):216–242, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85710669.

Imamoglu:1995:PPT

[572] O. Imamoglu. The Plücker parametrization and the theta multiplier. Journal of Number Theory, 52(2):243–255, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710670.

Okada:1995:ABM

[573] T. Okada, T. Sekiguchi, and Y. Shiota. Applications of binomial measures to power sums of digital sums. Journal of Number Theory, 52(2):256–266, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710682.

Bayad:1995:GSR

[574] A. Bayad. Galois structure of rings of integers and elliptic curves without complex multiplication. Journal of Number Theory, 52(2):267– 279, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710694.

Mignotte:1995:CCE

[575] M. Mignotte. A criterion on Catalan’s equation. Journal of Number Theory, 52(2):280–283, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710700.

Berge:1995:MVP

[576] A. M. Berge. Minimal vectors of pairs of dual lattices. Journal of Number Theory, 52(2):284–298, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710712.

Albu:1995:KFE

[577] T. Albu and F. Nicolae. Kneser field extensions with Cogalois correspondence. Journal of Number Theory, 52(2):299–318, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710724.

Thunder:1995:TIC

[578] J. L. Thunder. On Thue inequalities and a conjecture of Schmidt. Journal of Number Theory, 52(2):319–328, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710736.

Anonymous:1995:AIVc

[579] Anonymous. Author index for volume 52. Journal of Number Theory, 52(2):329, June 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710748.

Anonymous:1995:PJc

[580] Anonymous. Pages 1–210 (July 1995). Journal of Number Theory, 53(1): ??, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kuhner:1995:FGC

[581] J. Kuhner. On a family of generalized continued fraction expansions with period length going to infinity. Journal of Number Theory, 53(1): 1–12, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X8571075X.

Petermann:1995:GSD

[582] Y. F. S. Petermann. On Golomb’s self describing sequence. Journal of Number Theory, 53(1):13–24, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85710761.

Chiu:1995:CHP

[583] P. Chiu. Covering with Hecke points. Journal of Number Theory, 53(1): 25–44, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710773.

Vangeemen:1995:GAS

[584] B. Vangeemen and N. O. Nygaard. On the geometry and arithmetic of some Siegel modular threefolds. Journal of Number Theory, 53(1):

45–87, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710785.

Mccall:1995:IBB

[585] T. M. Mccall, C. J. Parry, and R. R. Ranalli. Imaginary bicyclic biquadratic fields with cyclic 2-class group. Journal of Number Theory, 53(1):88–99, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710797.

Gaal:1995:RIF

[586] I. Gaal, A. Petho, and M. Pohst. On the resolution of index form equations in biquadratic number-fields: III. The bicyclic biquadratic case. Journal of Number Theory, 53(1):100–114, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710803.

Lochter:1995:NCK

[587] M. Lochter. New characterizations of Kronecker equivalence. Journal of Number Theory, 53(1):115–136, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710815.

Ward:1995:ARF

[588] T. Ward. Additive relations in fields: an entropy approach. Journal of Number Theory, 53(1):137–143, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710827.

Chan:1995:NPR

[589] H. H. Chan. New proofs of Ramanujan’s partition identities for moduli 5 and 7. Journal of Number Theory, 53(1):144–158, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710839.

Guerry:1995:CGC

[590] G. Guerry. On the 2-component of the group of classes of certain cyclic extensions of degree 2^{N}. Journal of Number Theory, 53(1):159– 172, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710840.

Tzermias:1995:GAF

[591] P. Tzermias. Group of automorphisms of the Fermat curve. Journal of Number Theory, 53(1):173–178, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710852.

Watt:1995:KSM

[592] N. Watt. Kloosterman sums and a mean value for Dirichlet polynomials. Journal of Number Theory, 53(1):179–210, July 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710864.

Anonymous:1995:PAb

[593] Anonymous. Pages 211–356 (August 1995). Journal of Number Theory, 53(2):??, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Pacheco:1995:UGC

[594] A. Pacheco. Unramified Galois coverings of algebraic curves. Journal of Number Theory, 53(2):211–228, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710876.

Bosma:1995:CST

[595] W. Bosma and H. W. Lenstra. Complete systems of two addition laws for elliptic curves. Journal of Number Theory, 53(2):229–240, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710888.

Gao:1995:ATF

[596] W. D. Gao. Addition theorems for finite abelian groups. Journal of Number Theory, 53(2):241–246, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X8571089X.

Voutier:1995:UBS

[597] P. M. Voutier. An upper bound for the size of integral solutions to Y^{m} = f(X). Journal of Number Theory, 53(2):247–271, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710906.

Berthe:1995:LCW

[598] V. Berthe. Linear combinations of ζ(s)/Π^{s} over F_{q}(x), where 1 ≤ s ≤ q − 2. Journal of Number Theory, 53(2):272–299, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85710918.

Stuart:1995:NCA

[599] D. Stuart and R. Perlis. A new characterization of arithmetic equivalence. Journal of Number Theory, 53(2):300–308, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X8571092X.

Slavutskii:1995:NBN

[600] I. S. Slavutskii. A note on Bernoulli numbers. Journal of Number Theory, 53(2):309–310, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710931.

Abercrombie:1995:HDS

[601] A. G. Abercrombie. The Hausdorff dimension of some exceptional sets of p-adic integer matrices. Journal of Number Theory, 53(2):311– 341, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710943.

Oda:1995:NRG

[602] T. Oda. A note on ramification of the Galois representation on the fundamental group of an algebraic curve, II. Journal of Number Theory, 53(2):342–355, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710955.

Anonymous:1995:AIVd

[603] Anonymous. Author index for volume 53. Journal of Number Theory, 53(2):356, August 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710967.

Anonymous:1995:PS

[604] Anonymous. Pages 1–171 (September 1995). Journal of Number Theory, 54(1):??, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Evans:1995:SJC

[605] R. J. Evans. Selberg–Jack character sums of dimension 2. Journal of Number Theory, 54(1):1–11, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85710979.

Malouf:1995:TPC

[606] J. L. Malouf. On a theorem of Plünnecke concerning the sum of a basis and a set of positive density. Journal of Number Theory, 54(1):12–22, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85710980.

Miki:1995:IPS

[607] H. Miki. On Ihara’s power series. Journal of Number Theory, 54(1):23– 38, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85710992.

Ling:1995:SSD

[608] S. Ling. Shimura subgroups and degeneracy maps. Journal of Number Theory, 54(1):39–59, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85711006.

Flammang:1995:LTP

[609] V. Flammang. The length of totally positive algebraic-integers. Journal of Number Theory, 54(1):60–72, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85711018.

Vulakh:1995:DAB

[610] L. Y. Vulakh. Diophantine approximation on Bianchi groups. Journal of Number Theory, 54(1):73–80, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X8571102X.

Hata:1995:RMT

[611] M. Hata. Remarks on Mahler’s transcendence measure for e. Journal of Number Theory, 54(1):81–92, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85711031.

Demathan:1995:IMT

[612] B. Demathan. Irrationality measures and transcendence in positive characteristic. Journal of Number Theory, 54(1):93–112, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85711043.

Beukers:1995:TFE

[613] F. Beukers. Ternary form equations. Journal of Number Theory, 54 (1):113–133, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85711055.

Buell:1995:PFQ

[614] D. A. Buell and V. Ennola. On a parameterized family of quadratic and cubic fields. Journal of Number Theory, 54(1):134–148, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85711067.

Gerlach:1995:NWP

[615] H. Gerlach. On the number of witnesses (a mod n) providing a proper divisor of n. Journal of Number Theory, 54(1):149–160, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85711079.

Yu:1995:IDM

[616] J. K. Yu. Isogenies of Drinfeld modules over finite fields. Journal of Number Theory, 54(1):161–171, September 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85711080.

Anonymous:1995:PO

[617] Anonymous. Pages 173–341 (October 1995). Journal of Number Theory, 54(2):??, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kim:1995:CWS

[618] J. M. Kim. Coates–Wiles series and Mirimanoff’s polynomial. Journal of Number Theory, 54(2):173–179, October 1995. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85711092.

Moran:1995:MRN

[619] W. Moran and A. D. Pollington. Metrical results on normality to distinct bases. Journal of Number Theory, 54(2):180–189, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85711109.

Quebbemann:1995:MLE

[620] H. G. Quebbemann. Modular lattices in Euclidean spaces. Journal of Number Theory, 54(2):190–202, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85711110.

Schmidt:1995:IMR

[621] T. A. Schmidt. Infinitely many real quadratic fields of class number one. Journal of Number Theory, 54(2):203–205, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85711122.

Guo:1995:Z

[622] C. R. Guo. On the zeros of ζ(s) and ζ′(s). Journal of Number Theory, 54(2):206–210, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85711134.

Wagner:1995:RNW

[623] G. Wagner. On rings of numbers which are normal to one base but non-normal to another. Journal of Number Theory, 54(2):211–231, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711146.

Uhrin:1995:IPL

[624] B. Uhrin. The index of a point lattice in a set. Journal of Number Theory, 54(2):232–247, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85711158.

Hegyvari:1995:APS

[625] N. Hegyvari. Additive properties of sequences of multiplicatively perturbed square values. Journal of Number Theory, 54(2):248–260, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X8571116X.

Fleckinger:1995:FNG

[626] V. Fleckinger and M. Verant. Families of non–Galois quartic fields. Journal of Number Theory, 54(2):261–273, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85711171.

Costa:1995:DCR

[627] A. Costa and F. Gerth. Densities for 4-class ranks of totally complex quadratic extensions of real quadratic fields. Journal of Number Theory, 54(2):274–286, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85711183.

Blanchard:1995:RLI

[628] J. L. Blanchard. A representation of large integers from combinatorial sieves. Journal of Number Theory, 54(2):287–296, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85711195.

Perelli:1995:HLN

[629] A. Perelli and J. Pintz. Hardy–Littlewood numbers in short intervals. Journal of Number Theory, 54(2):297–308, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85711201.

Clemens:1995:CFS

[630] L. E. Clemens, K. D. Merrill, and D. W. Roeder. Continued Fractions and Series. Journal of Number Theory, 54(2):309–317, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711213.

Yu:1995:CNR

[631] J. K. Yu. A class number relation over function fields. Journal of Number Theory, 54(2):318–340, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85711225.

Anonymous:1995:AIVe

[632] Anonymous. Author index for volume 54. Journal of Number Theory, 54(2):341, October 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85711237.

Anonymous:1995:PN

[633] Anonymous. Pages 1–143 (November 1995). Journal of Number Theory, 55(1):??, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Barthel:1995:MRL

[634] L. Barthel and R. Livne. Modular representations of GL_{2} of a localfield: the ordinary, unramified case. Journal of Number Theory, 55(1): 1–27, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711249.

Mignosi:1995:GP

[635] F. Mignosi. On a generalization of the 3x+1 problem. Journal of Number Theory, 55(1):28–45, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85711250.

Trifonov:1995:GBF

[636] O. Trifonov. On gaps between k-free numbers. Journal of Number Theory, 55(1):46–59, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85711262.

Delmer:1995:GFS

[637] F. Delmer and J. M. Deshouillers. On a generalization of Farey sequences, II. Journal of Number Theory, 55(1):60–67, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X85711274.

Paysantleroux:1995:FFU

[638] R. Paysantleroux. Families of fundamental units in real quadratic fields and pure cubic fields. Journal of Number Theory, 55(1):68–81, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711286.

Alon:1995:PET

[639] N. Alon and M. N. Kolountzakis. On a problem of Erdős and Turán and some related results. Journal of Number Theory, 55(1):82–93, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711298.

Baxa:1995:DS

[640] C. Baxa. On the discrepancy of the sequence (nα). Journal of Number Theory, 55(1):94–107, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85711304.

Joux:1995:CSL

[641] A. Joux and F. Morain. Character sums linked to elliptic-curves with complex multiplication. Journal of Number Theory, 55(1):108–128, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711316.

Beauzamy:1995:PMV

[642] B. Beauzamy. Products of many-variable polynomials: Pairs that are maximal in Bombieri’s norm. Journal of Number Theory, 55(1):129– 143, November 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711328.

Anonymous:1995:PD

[643] Anonymous. Pages 145–322 (December 1995). Journal of Number Theory, 55(2):??, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

VanDerVlugt:1995:HDC

[644] Marcel Van Der Vlugt. Hasse–Davenport curves, Gauss sums, and weight distributions of irreducible cyclic codes. Journal of Number Theory, 55(2):145–159, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X8571133X.

Sakmar:1995:GRT

[645] I. A. Sakmar. General Ramanujan-type diophantine equations and their complete solution. Journal of Number Theory, 55(2):160–169, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711341.

Ahlswede:1995:DIS

[646] Rudolf Ahlswede and Levon H. Khachatrian. Density Inequalities for Sets of Multiples. Journal of Number Theory, 55(2):170–180, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711353.

Poonen:1995:RSI

[647] Bjorn Poonen. Rigidity and semi-invariants in Drinfeld modules. Journal of Number Theory, 55(2):181–196, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85711365.

Okazaki:1995:EPT

[648] Ryotaro Okazaki. An elementary proof for a theorem of Thomas and Vasquez. Journal of Number Theory, 55(2):197–208, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X85711377.

Le:1995:SED

[649] Maohua Le. Some exponential diophantine equations. I. The equation D_{1}x^{2} − D_{2}y^{2} = λk^{z}. Journal of Number Theory, 55(2):209–221, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711389.

Sweet:1995:CGM

[650] W. Jay Sweet, Jr. A computation of the gamma matrix of a family of p-adic zeta integrals. Journal of Number Theory, 55(2):222–260, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X85711390.

Dumir:1995:PVN

[651] V. C. Dumir, R. J. Hans-Gill, and Ranjeet Sehmi. Positive values of nonhomogeneous indefinite quadratic forms of type (2, 4). Journal of Number Theory, 55(2):261–284, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X85711407.

Laurent:1995:FLD

[652] Michel Laurent, Maurice Mignotte, and Yuri Nesterenko. Formes linéaires en deux logarithmes et déterminants d’interpolation. (French) [Linear forms in two logarithms and interpolation determinants]. Journal of Number Theory, 55(2):285–321, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X85711419.

Anonymous:1995:AIVf

[653] Anonymous. Author index for volume 55. Journal of Number Theory, 55(2):322, December 1995. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X85711420.

Anonymous:1996:PJa

[654] Anonymous. Pages 1–210 (January 1996). Journal of Number Theory, 56(1):??, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Chen:1996:NS

[655] Sheng Chen. A note on B_{2k} sequences. Journal of Number Theory, 56(1):1–3, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900013.

Kolountzakis:1996:DSM

[656] Mihail N. Kolountzakis. The density of B_{h}[g] sequences and the minimum of dense cosine sums. Journal of Number Theory, 56(1):4–11, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900025.

Kiming:1996:LDP

[657] Ian Kiming. On the liftings of 2-dimensional projective Galois representations over Q. Journal of Number Theory, 56(1):12–35, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900037.

Bleicher:1996:PPS

[658] Michael N. Bleicher. On Prielipp’s problem on signed sums of k-th powers. Journal of Number Theory, 56(1):36–51, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900049.

Dummit:1996:SCM

[659] David S. Dummit and Wendy L. Miller. The Steinitz class of the Mordell–Weil group of some CM elliptic curves. Journal of Number Theory, 56(1):52–78, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900050.

Schaefer:1996:CGS

[660] Edward F. Schaefer. Class groups and Selmer groups. Journal of Number Theory, 56(1):79–114, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900062. See corrigendum [1255].

Schwartz:1996:MOC

[661] Anne R. Schwartz and Thomas Shemanske. Maximal orders in central simple algebras and Bruhat–Tits buildings. Journal of Number Theory, 56(1):115–138, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900074.

Kucera:1996:SIC

[662] Radan Kučera. On the Stickelberger ideal and circular units of a compositum of quadratic fields. Journal of Number Theory, 56(1):139– 166, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900086.

Bollobas:1996:PPA

[663] Béla Bollobás and Guoping Jin. Partition problems in additive number theory. Journal of Number Theory, 56(1):167–187, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96900098.

Kolitsch:1996:DWS

[664] Louis Worthy Kolitsch. A different way to sort M-order, colored generalized Frobenius partitions. Journal of Number Theory, 56(1):188– 193, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900104.

Rogers:1996:MAG

[665] Michael Rogers. On a multiplicative–additive Galois invariant and wildly ramified extensions. Journal of Number Theory, 56(1):194–204, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900116.

Melfi:1996:TCA

[666] Giuseppe Melfi. On two conjectures about practical numbers. Journal of Number Theory, 56(1):205–210, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900128.

Anonymous:1996:PF

[667] Anonymous. Pages 211–419 (February 1996). Journal of Number Theory, 56(2):??, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Gao:1996:TAT

[668] W. D. Gao. Two addition theorems on groups of prime order. Journal of Number Theory, 56(2):211–213, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X9690013X.

Conrey:1996:MVD

[669] J. B. Conrey, Eric Fransen, Robert Klein, and Clayton Scott. Mean values of Dedekind sums. Journal of Number Theory, 56(2):214–226, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900141.

Lenstra:1996:CMS

[670] H. W. Lenstra, Jr. Complex multiplication structure of elliptic curves. Journal of Number Theory, 56(2):227–241, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900153.

Krieg:1996:JFS

[671] Aloys Krieg. Jacobi forms of several variables and the Maaß space. Journal of Number Theory, 56(2):242–255, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900165.

Leopoldt:1996:AFG

[672] Heinrich-Wolfgang Leopoldt. Über die Automorphismengruppe des Fermatkörpers. (German) [On the automorphism group of the Fermat body]. Journal of Number Theory, 56(2):256–282, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900177.

Reinhart:1996:DPM

[673] Georg Martin Reinhart. Determinants of partition matrices. Journal of Number Theory, 56(2):283–297, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900189.

Jia:1996:SLU

[674] Xingde Jia. B_{h}[g]-sequences with large upper density. Journal of Number Theory, 56(2):298–308, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96900190.

Li:1996:APP

[675] Hua-Chieh Li. p-adic periodic points and Sen’s Theorem. Journal of Number Theory, 56(2):309–318, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900207.

Wingberg:1996:RHF

[676] Kay Wingberg. A Riemann–Hurwitz formula for the p-rank of ideal class groups of CM-fields. Journal of Number Theory, 56(2):319–328, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900219.

Jung:1996:RPW

[677] Soon-Mo Jung and Bodo Volkmann. Remarks on a Paper of Wagner. Journal of Number Theory, 56(2):329–335, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900220.

Tang:1996:IIR

[678] Shu-Leung Tang. Iwasawa invariants of real abelian number fields. Journal of Number Theory, 56(2):336–342, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900232.

Baker:1996:BTT

[679] R. C. Baker. The Brun–Titchmarsh Theorem. Journal of Number Theory, 56(2):343–365, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900244.

David:1996:ADS

[680] Chantal David. Average distribution of supersingular Drinfeld modules. Journal of Number Theory, 56(2):366–380, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900256.

Howe:1996:CDC

[681] Everett W. Howe. Constructing distinct curves with isomorphic Jacobians. Journal of Number Theory, 56(2):381–390, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96900268.

Wirsing:1996:CKA

[682] Eduard Wirsing, Tang Yuan-sheng, and Shao Pin-tsung. On a conjecture of Kátai for additive functions. Journal of Number Theory, 56(2):391–395, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9690027X.

Frisch:1996:SCS

[683] Sophie Frisch. Substitution and closure of sets under integer-valued polynomials. Journal of Number Theory, 56(2):396–403, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96900281.

Alon:1996:PMR

[684] Noga Alon, Melvyn B. Nathanson, and Imre Ruzsa. The polynomial method and restricted sums of congruence classes. Journal of Number Theory, 56(2):404–417, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96900293.

Anonymous:1996:AIVa

[685] Anonymous. Author index for volume 56. Journal of Number Theory, 56(2):418–419, February 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9690030X.

Anonymous:1996:PMa

[686] Anonymous. Pages 1–206 (March 1996). Journal of Number Theory, 57 (1):??, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Nakamula:1996:CQF

[687] Ken Nakamula. Certain quartic fields with small regulators. Journal of Number Theory, 57(1):1–21, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96900311.

Car:1996:GPG

[688] Mireille Car. The generalized polynomial Goldbach problem. Journal of Number Theory, 57(1):22–49, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900323.

Schmidt:1996:RPF

[689] Thomas A. Schmidt. Rational period functions and parabolic cohomology. Journal of Number Theory, 57(1):50–65, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900335.

Smati:1996:IDN

[690] A. Smati. Sur l’itération du nombre de diviseurs des entiers sans grand facteur premier. (French) [On the iteration of the number of divisors of integers without large prime factor]. Journal of Number Theory, 57(1):66–89, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900347.

Gaal:1996:SRI

[691] István Gaál, Attila Pethő, and Michael Pohst. Simultaneous representation of integers by a pair of ternary quadratic forms — with an application to index form equations in quartic number fields. Journal of Number Theory, 57(1):90–104, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900359.

Lucht:1996:RSA

[692] Lutz Lucht and Cordelia Methfessel. Recurrent sequences and affine functional equations. Journal of Number Theory, 57(1):105–113, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96900360.

Feng:1996:KES

[693] Keqin Feng and Fei Xu. Kolyvagin’s “Euler Systems” in cyclotomic function fields. Journal of Number Theory, 57(1):114–121, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900372.

Sitaraman:1996:VR

[694] Sankar Sitaraman. Vandiver Revisited. Journal of Number Theory, 57(1):122–129, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900384.

Buzzard:1996:EHO

[695] Kevin Buzzard. On the eigenvalues of the Hecke operator T_{2}. Journal of Number Theory, 57(1):130–132, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900396.

Remond:1996:ADL

[696] Gaël Rémond and Florent Urfels. Approximation diophantienne de logarithmes elliptiques p-adiques. (French) [Diophantine approximation of p-adic elliptic logarithms]. Journal of Number Theory, 57(1):133– 169, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900402.

Ono:1996:CFP

[697] Ken Ono. Congruences for Frobenius partitions. Journal of Number Theory, 57(1):170–180, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900414.

Lagarias:1996:HBA

[698] Jeffrey C. Lagarias and Yang Wang. Haar bases for L^{2}(R^{n}) and algebraic number theory. Journal of Number Theory, 57(1):181–197, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900426.

Ruppert:1996:RCP

[699] Wolfgang M. Ruppert. Reducibility of cubic polynomials modp. Journal of Number Theory, 57(1):198–206, March 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900438.

Anonymous:1996:PAa

[700] Anonymous. Pages 207–410 (April 1996). Journal of Number Theory, 57(2):??, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Pappalardi:1996:OFG

[701] Francesco Pappalardi. On the order of finitely generated subgroups of Q ∗ (modp) and divisors of p − 1. Journal of Number Theory, 57(2): 207–222, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9690044X.

Zheng:1996:PKI

[702] Zhiyong Zheng. The Petersson–Knopp identity for the homogeneous Dedekind sums. Journal of Number Theory, 57(2):223–230, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96900451.

Leep:1996:QFA

[703] David B. Leep and Charles C. Yeomans. Quintic Forms over p-adic Fields. Journal of Number Theory, 57(2):231–241, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900463.

Khan:1996:CPZ

[704] Mizan R. Khan. Computation of partial zeta values at s = 0 over a totally real cubic field. Journal of Number Theory, 57(2):242– 277, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900475.

Wildberger:1996:ASG

[705] N. J. Wildberger. On the algebraic structure of Gaussian periods. Journal of Number Theory, 57(2):278–291, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900487.

Arno:1996:DAN

[706] Steven Arno, M. L. Robinson, and Ferell S. Wheeler. On denominators of algebraic numbers and integer polynomials. Journal of Number Theory, 57(2):292–302, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900499.

Liu:1996:APS

[707] H.-Q. Liu. Almost primes in short intervals. Journal of Number Theory, 57(2):303–322, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900505.

Calkin:1996:CSI

[708] Neil J. Calkin and Angela C. Taylor. Counting sets of integers, no k of which sum to another. Journal of Number Theory, 57(2):323– 327, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900517.

Byeon:1996:CNC

[709] Dongho Byeon and Hyun Kwang Kim. Class number 1 criteria for real quadratic fields of Richaud–Degert type. Journal of Number Theory, 57(2):328–339, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900529.

Guo:1996:BKC

[710] Li Guo. On the Bloch–Kato Conjecture for Hecke L-functions. Journal of Number Theory, 57(2):340–365, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900530.

Cassou-Nogues:1996:PDN

[711] Philippe Cassou-Noguès and Arnaud Jehanne. Parité du nombre de classes des S_{4}-extensions de Q et courbes elliptiques. (French) [Parity of the number of classes of S_{4}-extensions of Q and elliptic curves]. Journal of Number Theory, 57(2):366–384, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900542.

Robinson:1996:ILZ

[712] Margaret M. Robinson. The Igusa local zeta function associated with the singular cases of the determinant and the pfaffian. Journal of Number Theory, 57(2):385–408, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900554.

Anonymous:1996:AIVb

[713] Anonymous. Author index for volume 57. Journal of Number Theory, 57(2):409–410, April 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900566.

Anonymous:1996:PMb

[714] Anonymous. Pages 1–212 (May 1996). Journal of Number Theory, 58(1): ??, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

James:1996:NEQ

[715] Donald G. James. The number of embeddings of quadratic Z-lattices. Journal of Number Theory, 58(1):1–8, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900578.

Hensley:1996:PTA

[716] Douglas Hensley. A polynomial time algorithm for the Hausdorff dimension of continued fraction Cantor sets. Journal of Number Theory, 58(1):9–45, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9690058X. See [768].

Muller:1996:QD

[717] Peter Müller and Helmut Völklein. On a Question of Davenport. Journal of Number Theory, 58(1):46–54, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900591.

Voloch:1996:TEM

[718] José Felipe Voloch. Transcendence of elliptic modular functions in characteristic p. Journal of Number Theory, 58(1):55–59, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900608.

Thakur:1996:ASP

[719] Dinesh S. Thakur. Automata-style proof of Voloch’s result on transcendence. Journal of Number Theory, 58(1):60–63, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9690061X.

Girstmair:1996:RNB

[720] Kurt Girstmair. A remark on normal bases. Journal of Number Theory, 58(1):64–65, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900621.

Nomura:1996:RBQ

[721] Akito Nomura. A remark on Boston’s question concerning the existence of unramified p-extensions. Journal of Number Theory, 58(1): 66–70, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900633.

Haugland:1996:AMO

[722] Jan Kristian Haugland. Advances in the minimum overlap problem. Journal of Number Theory, 58(1):71–78, May 1996. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900645.

Lev:1996:STM

[723] Vsevolod F. Lev. Structure theorem for multiple addition and the Frobenius Problem. Journal of Number Theory, 58(1):79–88, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96900657. See addendum [918].

Fletcher:1996:SAP

[724] Peter Fletcher, William Lindgren, and Carl Pomerance. Symmetric and asymmetric primes. Journal of Number Theory, 58(1):89– 99, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900669.

Gao:1996:CPF

[725] W. D. Gao. A combinatorial problem on finite abelian groups. Journal of Number Theory, 58(1):100–103, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900670.

Ribenboim:1996:STL

[726] Paulo Ribenboim and Wayne L. McDaniel. The square terms in Lucas sequences. Journal of Number Theory, 58(1):104–123, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900682.

Helm:1996:DS

[727] Martin Helm. On the distribution of B_{3}-sequences. Journal of Number Theory, 58(1):124–129, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900694.

Yuansheng:1996:RPA

[728] Tang Yuansheng. A reverse problem on arithmetic functions. Journal of Number Theory, 58(1):130–138, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900700.

Wang:1996:TSM

[729] Julie T.-Y. Wang. The truncated second main theorem of function fields. Journal of Number Theory, 58(1):139–157, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900712.

McIntosh:1996:AFB

[730] Richard J. McIntosh. An asymptotic formula for binomial sums. Journal of Number Theory, 58(1):158–172, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900724.

Skula:1996:SIC

[731] Ladislav Skula. On a special ideal contained in the Stickelberger ideal. Journal of Number Theory, 58(1):173–195, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900736.

Wan:1996:RHC

[732] Daqing Wan. On the Riemann hypothesis for the characteristic p zeta function. Journal of Number Theory, 58(1):196–212, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900748.

Anonymous:1996:PJb

[733] Anonymous. Pages 213–416 (June 1996). Journal of Number Theory, 58 (2):??, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Banaszak:1996:ESH

[734] Grzegorz Banaszak and Wojciech Gajda. Euler systems for higher K-theory of number fields. Journal of Number Theory, 58(2):213– 252, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9690075X.

Laubie:1996:SSF

[735] F. Laubie and A. Movahhedi. Substitutions de séries formelles et cyclotomie. (French) [Substitutions for formal series and cyclotomy]. Journal of Number Theory, 58(2):253–266, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900761.

Hsu:1996:LSI

[736] Chih-Nung Hsu. A large sieve inequality for rational function fields. Journal of Number Theory, 58(2):267–287, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900773.

Wooley:1996:NSC

[737] Trevor D. Wooley. A note on simultaneous congruences. Journal of Number Theory, 58(2):288–297, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96900785.

Dodson:1996:SDA

[738] M. M. Dodson, B. P. Rynne, and J. A. G. Vickers. Simultaneous Diophantine approximation and asymptotic formulae on manifolds. Journal of Number Theory, 58(2):298–316, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900797.

Shlapentokh:1996:DUA

[739] Alexandra Shlapentokh. Diophantine undecidability over algebraic function fields over finite fields of constants. Journal of Number Theory, 58(2):317–342, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900803.

Pacheco:1996:RPI

[740] Amílcar Pacheco. Rational points on Igusa curves and L-functions of symmetric representations. Journal of Number Theory, 58(2):343– 360, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900815.

Bley:1996:AAO

[741] W. Bley and D. Burns. Über arithmetische assoziierte Ordnungen. (German) [On arithmetic associated orders]. Journal of Number Theory, 58(2):361–387, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900827.

Ramharter:1996:MIP

[742] G. Ramharter. On Mordell’s inverse problem in dimension three. Journal of Number Theory, 58(2):388–415, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900839.

Anonymous:1996:AIVc

[743] Anonymous. Author index for volume 58. Journal of Number Theory, 58(2):416, June 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900840.

Anonymous:1996:PJc

[744] Anonymous. Pages 1–213 (July 1996). Journal of Number Theory, 59(1): ??, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Wu:1996:SEF

[745] J. Wu. A sharpening of effective formulas of Selberg–Delange type for some arithmetic functions on the semigroup G_{K}. Journal of Number Theory, 59(1):1–19, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900852.

vanderGeer:1996:QFG

[746] Gerard van der Geer and Marcel van der Vlugt. Quadratic forms, generalized Hamming weights of codes and curves with many points. Journal of Number Theory, 59(1):20–36, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96900864.

Goto:1996:AWD

[747] Yasuhiro Goto. Arithmetic of weighted diagonal surfaces over finite fields. Journal of Number Theory, 59(1):37–81, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900876.

Buchmann:1996:SIN

[748] Johannes A. Buchmann and Christine S. Hollinger. On smooth ideals in number fields. Journal of Number Theory, 59(1):82–87, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96900888.

Moree:1996:DPD

[749] Pieter Moree and Gary L. Mullen. Dickson polynomial discriminators. Journal of Number Theory, 59(1):88–105, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9690089X.

Zheng:1996:DZI

[750] Zhiyong Zheng. The distribution of zeros of an irreducible curve over a finite field. Journal of Number Theory, 59(1):106–118, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96900906.

Solomon:1996:ALS

[751] David Solomon. p-adic limits of Shintani generating functions for a real quadratic field. Journal of Number Theory, 59(1):119–158, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96900918.

Heiermann:1996:NIN

[752] Volker Heiermann. De nouveaux invariants numériques pour les extensions totalement ramifiées de corps locaux. (French) [New numerical invariants for fully branched extensions of local bodies]. Journal of Number Theory, 59(1):159–202, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9690092X.

Adongo:1996:DE

[753] Harun P. K. Adongo. The diophantine equation 3u^{2} − 2 = v^{6}. Journal of Number Theory, 59(1):203–208, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900931.

Wolke:1996:PEA

[754] Dieter Wolke. On a problem of Erdős in additive number theory. Journal of Number Theory, 59(1):209–213, July 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900943.

Anonymous:1996:PAb

[755] Anonymous. Pages 215–420 (August 1996). Journal of Number Theory, 59(2):??, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Idogawa:1996:WRD

[756] Tomoyuki Idogawa and Masayoshi Tsutsumi. A way to remove duplication from ∑_{i=1}^{N} 1/x_{i}. Journal of Number Theory, 59(2):215–224, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900955.

Saby:1996:TIG

[757] Nicolas Saby. Théorie d’Iwasawa géométrique: un théorème de comparaison. (French) [Iwasawa geometry theory: a comparison theorem]. Journal of Number Theory, 59(2):225–247, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900967.

Thakur:1996:ECF

[758] Dinesh S. Thakur. Exponential and continued fractions. Journal of Number Theory, 59(2):248–261, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96900979.

Green:1996:CVR

[759] Barry Green. On curves over valuation rings and morphisms to P^{1}. Journal of Number Theory, 59(2):262–290, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96900980.

Komatsu:1996:CPS

[760] Takao Komatsu. A certain power series and the inhomogeneous continued fraction expansions. Journal of Number Theory, 59(2):291– 312, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96900992.

Diaz-Vargas:1996:RH

[761] Javier Diaz-Vargas. Riemann Hypothesis for F_{p}[T]. Journal of Number Theory, 59(2):313–318, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96901006.

Ozluk:1996:APC

[762] Ali E. Özlük. On the q-analogue of the pair correlation conjecture. Journal of Number Theory, 59(2):319–351, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901018.

Dumir:1996:CGT

[763] V. C. Dumir, R. J. Hans-Gill, and J. B. Wilker. Contributions to a general theory of view-obstruction problems, II. Journal of Number Theory, 59(2):352–373, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9690102X.

Adelberg:1996:CAI

[764] Arnold Adelberg. Congruences of p-adic integer order Bernoulli numbers. Journal of Number Theory, 59(2):374–388, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901031.

Amou:1996:MRT

[765] Masaaki Amou. A metrical result on transcendence measures in certain fields. Journal of Number Theory, 59(2):389–397, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901043.

Ayyad:1996:CEM

[766] Anwar Ayyad, Todd Cochrane, and Zhiyong Zheng. The congruence x_{1}x_{2} ≡ x_{3}x_{4}(modp), the equation x_{1}x_{2} ≡ x_{3}x_{4}, and mean values of character sums. Journal of Number Theory, 59(2):398–413, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901055.

Joshi:1996:FEC

[767] Kirti Joshi. A family of étale coverings of the affine line. Journal of Number Theory, 59(2):414–418, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96901067.

Hensley:1996:EPA

[768] Douglas Hensley. Erratum: A Polynomial Algorithm for the Hausdorff Dimension of Continued Fraction Cantor Sets: Volume 58, Number 1 (1996), pages 9–45. Journal of Number Theory, 59(2):419, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901079. See [716].

Anonymous:1996:AIVd

[769] Anonymous. Author index for volume 59. Journal of Number Theory, 59(2):420, August 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901080.

Anonymous:1996:PS

[770] Anonymous. Pages 1–209 (September 1996). Journal of Number Theory, 60(1):??, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Feng:1996:SHA

[771] Keqin Feng and Wen-Ch’ing Winnie Li. Spectra of Hypergraphs and Applications. Journal of Number Theory, 60(1):1–22, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96901092.

Dilcher:1996:SPB

[772] Karl Dilcher. Sums of products of Bernoulli numbers. Journal of Number Theory, 60(1):23–41, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96901109.

Yamamura:1996:MUE

[773] Ken Yamamura. The maximal unramified extensions of the imaginary quadratic number fields with class number two. Journal of Number Theory, 60(1):42–50, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901110.

Hirschhorn:1996:SAF

[774] Michael D. Hirschhorn and James A. Sellers. Some amazing facts about 4cores. Journal of Number Theory, 60(1):51–69, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901122.

Tsumura:1996:DMM

[775] Hirofumi Tsumura. On Demjanenko’s matrix and Maillet’s determinant for imaginary abelian number fields. Journal of Number Theory, 60(1):70–79, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901134.

Amoroso:1996:ANC

[776] Francesco Amoroso. Algebraic numbers close to 1 and variants of Mahler’s measure. Journal of Number Theory, 60(1):80–96, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96901146.

Kiming:1996:NTG

[777] Ian Kiming. A Note on a Theorem of A. Granville and K. Ono. Journal of Number Theory, 60(1):97–102, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901158.

Barat:1996:DPA

[778] Guy Barat and Peter J. Grabner. Distribution properties of Gadditive functions. Journal of Number Theory, 60(1):103–123, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X9690116X.

Taguchi:1996:M

[779] Yuichiro Taguchi. On φ-modules. Journal of Number Theory, 60 (1):124–141, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901171.

Anderson:1996:ALI

[780] Greg W. Anderson. Another look at the index formulas of cyclotomic number theory. Journal of Number Theory, 60(1):142–164, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901183.

Anderson:1996:LAT

[781] Greg W. Anderson. Log-algebraicity of twisted A-harmonic series and special values of L-series in characteristic p. Journal of Number Theory, 60(1):165–209, September 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901195.

Anonymous:1996:PO

[782] Anonymous. Pages 211–418 (October 1996). Journal of Number Theory, 60(2):??, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Liedahl:1996:AWP

[783] Steven Liedahl. K-admissibility of wreath products of cyclic p-groups. Journal of Number Theory, 60(2):211–232, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901201.

Shparlinski:1996:ESS

[784] Igor Shparlinski. On exponential sums with sparse polynomials and rational functions. Journal of Number Theory, 60(2):233–244, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901213.

Kraus:1996:CES

[785] Alain Kraus. Courbes elliptiques semi-stables et corps quadratiques. (French) [Semi-stable elliptical curves and quadratic bodies]. Journal of Number Theory, 60(2):245–253, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901225.

Lassak:1996:FET

[786] Miroslav Laššák and Štefan Porubský. Fermat–Euler Theorem in algebraic number fields. Journal of Number Theory, 60(2):254–290, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901237.

Albu:1996:FRF

[787] Toma Albu and Florin Nicolae. Finite radical field extensions and crossed homomorphisms. Journal of Number Theory, 60(2):291–309, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901249.

Nebe:1996:NEL

[788] Gabriele Nebe and Boris B. Venkov. Nonexistence of extremal lattices in certain genera of modular lattices. Journal of Number Theory, 60(2):310–317, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901250.

Allouche:1996:TCG

[789] Jean-Paul Allouche. Transcendence of the Carlitz–Goss Gamma function at rational arguments. Journal of Number Theory, 60(2):318– 328, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901262.

Hoffman:1996:STH

[790] Michael E. Hoffman and Courtney Moen. Sums of triple harmonic series. Journal of Number Theory, 60(2):329–331, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901274.

Geyer:1996:PGG

[791] W.-D. Geyer and C. U. Jensen. Prodihedral groups as Galois groups over number fields. Journal of Number Theory, 60(2):332–372, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901286.

Bilu:1996:STE

[792] Yuri Bilu and Guillaume Hanrot. Solving Thue equations of high degree. Journal of Number Theory, 60(2):373–392, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901298.

Woodcock:1996:SAA

[793] C. F. Woodcock. Special p-adic analytic functions and Fourier transforms. Journal of Number Theory, 60(2):393–408, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901304.

Zhang:1996:EIB

[794] XianKe Zhang and FuHua Xu. Existence of integral bases for relative extensions of n-cyclic number fields. Journal of Number Theory, 60(2):409–416, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901316.

Anonymous:1996:AIVe

[795] Anonymous. Author index for volume 60. Journal of Number Theory, 60(2):417–418, October 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901328.

Anonymous:1996:PN

[796] Anonymous. Pages 1–208 (November 1996). Journal of Number Theory, 61(1):??, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Zhang:1996:MVQ

[797] Yuanli Zhang. Mean values of quadratic twists of automorphic L-functions. Journal of Number Theory, 61(1):1–24, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X9690133X.

Mauduit:1996:ASS

[798] Christian Mauduit and Andràs Sárközy. On the arithmetic structure of sets characterized by sum of digits properties. Journal of Number Theory, 61(1):25–38, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96901341.

Erdos:1996:SPN

[799] Paul Erdős and Anthony B. Evans. Sets of prime numbers satisfying a divisibility condition. Journal of Number Theory, 61(1):39–43, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901353.

Cram:1996:IRC

[800] G.-Martin Cram and Olaf Neiße. On integral representations over cyclotomic fields. Journal of Number Theory, 61(1):44–51, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96901365.

Masser:1996:SPE

[801] D. W. Masser and J. H. Rickert. Simultaneous Pell equations. Journal of Number Theory, 61(1):52–66, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901377.

Myerson:1996:WLC

[802] G. Myerson and J. W. Sander. What the least common multiple divides, II. Journal of Number Theory, 61(1):67–84, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901389.

Hsu:1996:DIP

[803] Chih-Nung Hsu. The distribution of irreducible polynomials in F_{q}[t]. Journal of Number Theory, 61(1):85–96, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901390.

Gao:1996:ZSS

[804] Weidong Gao. On zero-sum subsequences of restricted size. Journal of Number Theory, 61(1):97–102, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901407.

Esselmann:1996:MDL

[805] Frank Esselmann. Über die maximale Dimension von Lorentz-Gittern mit coendlicher Spiegelungsgruppe. (German) [On the maximum dimension of Lorentz gratings with a coincident reflection group]. Journal of Number Theory, 61(1):103–144, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901419.

Car:1996:QF

[806] Mireille Car. Quadratic forms on F_{q}[T]. Journal of Number Theory, 61(1):145–180, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901420.

Martin:1996:CTJ

[807] Yves Martin. A converse theorem for Jacobi forms. Journal of Number Theory, 61(1):181–193, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X96901432.

Burger:1996:UAN

[808] Edward B. Burger. Uniformly approximable numbers and the uniform approximation spectrum. Journal of Number Theory, 61(1):194–208, November 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901444.

Anonymous:1996:PD

[809] Anonymous. Pages 209–422 (December 1996). Journal of Number Theory, 61(2):??, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kargaev:1996:ARN

[810] Pavel Kargaev and Anatoly Zhigljavsky. Approximation of real numbers by rationals: some metric theorems. Journal of Number Theory, 61(2):209–225, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901456.

Cahen:1996:PC

[811] Paul-Jean Cahen. Polynomial Closure. Journal of Number Theory, 61(2):226–247, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901468.

Garcia:1996:ABS

[812] Arnaldo Garcia and Henning Stichtenoth. On the asymptotic behaviour of some towers of function fields over finite fields. Journal of Number Theory, 61(2):248–273, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X9690147X.

Padma:1996:ECC

[813] R. Padma and S. Venkataraman. Elliptic curves with complex multiplication and a character sum. Journal of Number Theory, 61(2):274–282, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901481.

Chapman:1996:APC

[814] Robin J. Chapman. Automorphism polynomials in cyclic cubic extensions. Journal of Number Theory, 61(2):283–291, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X96901493.

Brindza:1996:PVD

[815] B. Brindza, Á. Pintér, and J. Végső. On polynomial values of the discriminants of characteristic polynomials. Journal of Number Theory, 61(2):292–300, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9690150X.

Wenpeng:1996:DIM

[816] Zhang Wenpeng. On the distribution of inverses modulo n. Journal of Number Theory, 61(2):301–310, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901511.

Bugeaud:1996:MED

[817] Y. Bugeaud and M. Laurent. Minoration effective de la distance p-adique entre puissances de nombres algébriques. (French) [Effective minimization of the p-adic distance between powers of algebraic numbers]. Journal of Number Theory, 61(2):311–342, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901523. See

note [1254].

Dohmae:1996:BGC

[818] Kazuhiro Dohmae. On bases of groups of circular units of some imaginary abelian number fields. Journal of Number Theory, 61(2):343–364, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901535.

Dummigan:1996:SGL

[819] Neil Dummigan. Symplectic group lattices as Mordell–Weil sublattices. Journal of Number Theory, 61(2):365–387, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901547.

Graham:1996:SPR

[820] Ronald Graham, Vojtech Rödl, and Andrzej Ruciński. On Schur properties of random subsets of integers. Journal of Number Theory, 61 (2):388–408, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901559.

Ichikawa:1996:TCT

[821] Takashi Ichikawa. Theta constants and Teichmüller modular forms. Journal of Number Theory, 61(2):409–419, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X96901560.

Anonymous:1996:A

[822] Anonymous. Addendum. Journal of Number Theory, 61(2):420, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X96901572.

Anonymous:1996:AIVf

[823] Anonymous. Author index for volume 61. Journal of Number Theory, 61(2):421–422, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96901584.

Anonymous:1997:PJa

[824] Anonymous. Pages 1–220 (January 1997). Journal of Number Theory, 62(1):??, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Mauclaire:1997:SCR

[825] Jean-Loup Mauclaire. Some consequences of a result of Jean Coquet. Journal of Number Theory, 62(1):1–18, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920370.

Dumont:1997:GAP

[826] Jean Marie Dumont and Alain Thomas. Gaussian asymptotic properties of the sum-of-digits function. Journal of Number Theory, 62(1):

19–38, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920448.

Bremner:1997:SCS

[827] A. Bremner, R. J. Stroeker, and N. Tzanakis. On sums of consecutive squares. Journal of Number Theory, 62(1):39–70, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920400.

Hua:1997:CSD

[828] Chen Jian Hua and Paul Voutier. Complete solution of the diophantine equation X^{2} + 1 = dY^{4} and a related family of quartic Thue equations. Journal of Number Theory, 62(1):71–99, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920187.

Le:1997:NNS

[829] Maohua Le. A note on the number of solutions of the generalized Ramanujan–Nagell equation D_{1}x^{2} + D_{2} = 4p^{n}. Journal of Number Theory, 62(1):100–106, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920199.

Khare:1997:MBJ

[830] Chandrashekhar Khare. Maps between Jacobians of modular curves. Journal of Number Theory, 62(1):107–114, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9792028X.

Gupta:1997:RSM

[831] Shamita Dutta Gupta. On the Rankin–Selberg method for functions not of rapid decay on congruence subgroups. Journal of Number Theory, 62(1):115–126, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920357.

Lev:1997:ORS

[832] Vsevolod F. Lev. Optimal representations by sumsets and subset sums. Journal of Number Theory, 62(1):127–143, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920126.

Kwon:1997:TSE

[833] Soonhak Kwon. Torsion subgroups of elliptic curves over quadratic extensions. Journal of Number Theory, 62(1):144–162, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920369.

Saias:1997:EDD

[834] Eric Saias. Entiers à diviseurs denses 1. (French) [Whole numbers with dense divisors 1]. Journal of Number Theory, 62(1):163–191, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920576.

Komatsu:1997:ICF

[835] Takao Komatsu. On inhomogeneous continued fraction expansions and inhomogeneous diophantine approximation. Journal of Number Theory, 62(1):192–212, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920606.

Aiba:1997:CMG

[836] Akira Aiba. Carlitz modules and Galois module structure. Journal of Number Theory, 62(1):213–219, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920667.

Anonymous:1997:Ea

[837] Anonymous. Erratum. Journal of Number Theory, 62(1):220, January 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920977.

Anonymous:1997:PF

[838] Anonymous. Pages 221–439 (February 1997). Journal of Number Theory, 62(2):??, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Poe:1997:DSU

[839] Milja Poe. On distribution of solutions of S-unit equations. Journal of Number Theory, 62(2):221–241, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920461.

Bley:1997:CAO

[840] W. Bley. Computing associated orders and Galois generating elements of unit lattices. Journal of Number Theory, 62(2):242–256, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920503.

Byeon:1997:CNC

[841] Dongho Byeon and Hyun Kwang Kim. Class number 2 criteria for real quadratic fields of Richaud–Degert type. Journal of Number Theory, 62(2):257–272, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9792059X.

Spiess:1997:CFH

[842] Michael Spiess. Class formations and higher dimensional local class field theory. Journal of Number Theory, 62(2):273–283, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97920485.

Li:1997:IBD

[843] Hua-Chieh Li. Isogenies between dynamics of formal groups. Journal of Number Theory, 62(2):284–297, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920618.

Takagi:1997:PFC

[844] Tsuyoshi Takagi. Product formula of the cubic Gauss sum modulo the product of the primes. Journal of Number Theory, 62(2):298–306, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920643.

Zhao:1997:RNC

[845] Jianqiang Zhao. On root numbers connected with special values of L-functions over F_{q}(T). Journal of Number Theory, 62(2):307–321, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920564.

Grisel:1997:LPR

[846] Guillaume Grisel. Length of the powers of a rational fraction. Journal of Number Theory, 62(2):322–337, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920540.

Stern:1997:CEN

[847] Leonid Stern. Criterion for the equality of norm groups of Idele groups of algebraic number fields. Journal of Number Theory, 62(2):338–352, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920631.

Zhu:1997:CIP

[848] Fu-Zu Zhu. On the construction of indecomposable positive definite Hermitian forms over imaginary quadratic fields. Journal of Number Theory, 62(2):353–367, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920497.

Earnest:1997:RIP

[849] A. G. Earnest and Azar Khosravani. Representation of integers by positive definite binary Hermitian lattices over imaginary quadratic fields. Journal of Number Theory, 62(2):368–374, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920539.

Schneiders:1997:ERC

[850] Ursula Schneiders. Estimating the 2-rank of cubic fields by Selmer groups of elliptic curves. Journal of Number Theory, 62(2):375–396, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9792062X.

Ruzsa:1997:CTE

[851] Imre Z. Ruzsa. A Converse to a Theorem of Erdős and Fuchs. Journal of Number Theory, 62(2):397–402, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920679.

Gras:1997:PIP

[852] Georges Gras. Principalisation d’idéaux par extensions absolument abéliennes. Journal of Number Theory, 62(2):403–421, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97920680.

Degot:1997:FDM

[853] Jérome Dégot. Finite-dimensional Mahler measure of a polynomial and Szegő’s theorem. Journal of Number Theory, 62(2):422–427, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920527.

Ahlgren:1997:PEE

[854] Scott D. Ahlgren. Polynomial–exponential equations in two variables. Journal of Number Theory, 62(2):428–438, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920588.

Anonymous:1997:AIVa

[855] Anonymous. Author index for volume 62. Journal of Number Theory, 62(2):439, February 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921156.

Anonymous:1997:PMa

[856] Anonymous. Pages 1–190 (March 1997). Journal of Number Theory, 63 (1):??, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Chen:1997:NTI

[857] Zesen Chen and David R. Hayes. Norm and trace of the j-invariants of Drinfeld modules associated to hyperelliptic curves. Journal of Number Theory, 63(1):1–11, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920849.

Arledge:1997:UAG

[858] Jane Arledge. S-units attached to genus 3 hyperelliptic curves. Journal of Number Theory, 63(1):12–29, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920734.

Keune:1997:QNF

[859] Frans Keune. Quadratic number fields having {−1, −1} as a non-trivial element of their wild kernel. Journal of Number Theory, 63(1):30– 33, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920771.

Ahlswede:1997:NTC

[860] Rudolf Ahlswede and Levon H. Khachatrian. Number theoretic correlation inequalities for Dirichlet densities. Journal of Number Theory, 63(1):34–46, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920722.

Grundman:1997:HMV

[861] H. G. Grundman. Hilbert modular varieties of Galois quartic fields. Journal of Number Theory, 63(1):47–58, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920746.

Langevin:1997:CCS

[862] Philippe Langevin. Calculs de certaines sommes de Gauss. (French) [Calculations of certain sums of Gauss]. Journal of Number Theory, 63(1):59–64, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920783.

Sinha:1997:DRF

[863] Samarendra K. Sinha. Deligne’s reciprocity for function fields. Journal of Number Theory, 63(1):65–88, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920825.

Kim:1997:RPD

[864] Myung-Hwan Kim and Byeong-Kweon Oh. Representations of positive definite senary integral quadratic forms by a sum of squares. Journal of Number Theory, 63(1):89–100, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920692.

Gupta:1997:MVF

[865] Shamita Dutta Gupta. Mean values of L-functions over function fields. Journal of Number Theory, 63(1):101–131, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920758.

Meir:1997:PAC

[866] Ivan D. Meir. Pairs of additive congruences to a large prime modulus. Journal of Number Theory, 63(1):132–142, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920710.

Laubie:1997:RA

[867] F. Laubie and M. Saïne. Ramification of automorphisms of k((t)). Journal of Number Theory, 63(1):143–145, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920795.

Hsia:1997:DHZ

[868] Liang chung Hsia. On the dynamical height zeta functions. Journal of Number Theory, 63(1):146–169, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X9792076X.

Taguchi:1997:EFS

[869] Yuichiro Taguchi and Daqing Wan. Entireness of L-functions of φsheaves on affine complete intersections. Journal of Number Theory, 63(1):170–179, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920552.

Bertrand-Mathis:1997:ENR

[870] Anne Bertrand-Mathis. Ensembles normaux relatifs a des matrices non-commutantes. Journal of Number Theory, 63(1):180–190, March 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920837.

Anonymous:1997:PAa

[871] Anonymous. Pages 191–419 (April 1997). Journal of Number Theory, 63(2):??, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Sofer:1997:AAJ

[872] Adriana Sofer. p-adic aspects of Jacobi forms. Journal of Number Theory, 63(2):191–202, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920953.

Sandor:1997:PE

[873] Csaba Sándor. On a problem of Erdős. Journal of Number Theory, 63(2):203–210, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921132.

Call:1997:CHP

[874] Gregory S. Call and Susan W. Goldstine. Canonical heights on projective space. Journal of Number Theory, 63(2):211–243, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920990.

Bachman:1997:DPC

[875] Gennady Bachman. On divisibility properties of certain multinomial coefficients. Journal of Number Theory, 63(2):244–255, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97920989.

Raouj:1997:DCE

[876] A. Raouj. Sur la densité de certains ensembles de multiples, III. (French) [On the density of certain sets of multiples, III]. Journal of Number Theory, 63(2):256–266, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920473.

Cilleruelo:1997:LPH

[877] Javier Cilleruelo and Jorge Jiménez-Urroz. Lattice Points on Hyperbolas. Journal of Number Theory, 63(2):267–274, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920515.

Tomita:1997:ERF

[878] Koshi Tomita. Explicit representation of fundamental units of some real quadratic fields, II. Journal of Number Theory, 63(2):275–285, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920886.

Pintz:1997:VLG

[879] János Pintz. Very large gaps between consecutive primes. Journal of Number Theory, 63(2):286–301, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920813.

Yin:1997:ICU

[880] Linsheng Yin. On the index of cyclotomic units in characteristic p and its applications. Journal of Number Theory, 63(2):302– 324, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921119.

DeKoninck:1997:NCI

[881] Jean-Marie De Koninck, Imre Kátai, and Bui Minh Phong. A new characteristic of the identity function. Journal of Number Theory, 63(2): 325–338, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921089.

Gras:1997:CEP

[882] Georges Gras and Stéphane Jeannin. Critère effectif de puissance pième dans un corps de nombres galoisien. (French) [Effective criterion of p-th power in a field of Galois numbers]. Journal of Number Theory, 63(2):339–356, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920850.

Gupta:1997:IPF

[883] Rajiv Gupta and Keith Ramsay. Indivisible points on families of elliptic curves. Journal of Number Theory, 63(2):357–372, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97920928.

deWeger:1997:EBC

[884] Benjamin M. M. de Weger. Equal binomial coefficients: some elementary considerations. Journal of Number Theory, 63(2):373–386, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921090.

Khare:1997:BCL

[885] Chandrashekhar Khare. Base Change, Lifting, and Serre’s Conjecture. Journal of Number Theory, 63(2):387–395, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9792093X.

MendesFrance:1997:TCG

[886] Michel Mendès France and Jia yan Yao. Transcendence and the Carlitz–Goss gamma function. Journal of Number Theory, 63(2):396– 402, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921041.

Kucera:1997:NSD

[887] Radan Kučera. A note on Sinnott’s definition of circular units of an abelian field. Journal of Number Theory, 63(2):403–407, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97920941.

Washington:1997:FCF

[888] Lawrence C. Washington. A family of cubic fields and zeros of 3adic L-functions. Journal of Number Theory, 63(2):408–417, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97920965.

Anonymous:1997:AIVb

[889] Anonymous. Author index for volume 63. Journal of Number Theory, 63(2):418–419, April 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921168.

Anonymous:1997:PMb

[890] Anonymous. Pages 1–163 (May 1997). Journal of Number Theory, 64(1): ??, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Nakkajima:1997:ILD

[891] Yukiyoshi Nakkajima. On infinitesimal liftings and degenerations of Hodge–de Rham spectral sequences. Journal of Number Theory, 64(1):

1–12, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921120.

Berend:1997:PEF

[892] Daniel Berend. On the parity of exponents in the factorization of n!. Journal of Number Theory, 64(1):13–19, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921065.

Liu:1997:ETCa

[893] Jianya Liu. On an error term of Chowla, I. Journal of Number Theory, 64(1):20–35, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921016.

Liu:1997:ETCb

[894] Jianya Liu. On an error term of Chowla, II. Journal of Number Theory, 64(1):36–50, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921028.

Liu:1997:ETCc

[895] Jianya Liu. On an error term of Chowla, III. Journal of Number Theory, 64(1):51–58, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9792103X.

vanWamelen:1997:CCC

[896] Paul van Wamelen. On the CM character of the curves Y^{2} = x^{q} − 1. Journal of Number Theory, 64(1):59–83, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97920898.

Efrat:1997:PGG

[897] Ido Efrat. Pro-p Galois groups of algebraic extensions of Q. Journal of Number Theory, 64(1):84–99, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97920916.

Dabrowski:1997:NAF

[898] Andrzej Dabrowski. A note on p-adic q-ζ-functions. Journal of Number Theory, 64(1):100–103, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921077.

David:1997:PPH

[899] Sinnou David. Points de petite hauteur sur les courbes elliptiques. (French) [Points of small height on elliptical curves]. Journal of Number Theory, 64(1):104–129, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921004.

Kani:1997:ECG

[900] Ernst Kani. The existence of curves of genus two with elliptic differentials. Journal of Number Theory, 64(1):130–161, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921053.

Metsankyla:1997:LE

[901] Tauno Metsänkylä. Letter to the Editor. Journal of Number Theory, 64(1):162–163, May 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921399.

Anonymous:1997:PJb

[902] Anonymous. Pages 165–339 (June 1997). Journal of Number Theory, 64 (2):??, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Leprevost:1997:RCE

[903] F. Leprévost and F. Morain. Revêtement de courbes elliptiques à multiplication complexe par des courbes hyperelliptiques et sommes de caractères. (French) [Covering of elliptic curves with complex multiplication by hyperelliptic curves and sums of characters]. Journal of Number Theory, 64(2):165–182, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920709.

Katsurada:1997:RFL

[904] Hidenori Katsurada and Masaki Hisasue. A recursion formula for local densities. Journal of Number Theory, 64(2):183–210, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921272.

Ozaki:1997:CUGa

[905] Manabu Ozaki. On the cyclotomic unit group and the ideal class group of a real abelian number field. Journal of Number Theory, 64(2):211– 222, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97920862.

Ozaki:1997:CUGb

[906] Manabu Ozaki. On the cyclotomic unit group and the ideal class group of a real abelian number field ii. Journal of Number Theory, 64(2):223–232, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97920874.

Bilu:1997:ASI

[907] Yuri Bilu. Addition of sets of integers of positive density. Journal of Number Theory, 64(2):233–275, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921235.

Corrales-Rodriganez:1997:SPE

[908] Capi Corrales-Rodrigáñez and René Schoof. The support problem and its elliptic analogue. Journal of Number Theory, 64(2):276– 290, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921144.

Philippon:1997:AMP

[909] Patrice Philippon. Une approche méthodique pour la transcendance et l’indépendance algébrique de valeurs de fonctions analytiques. (French) [A methodical approach for transcendence and algebraic independence of values of analytical functions]. Journal of Number Theory, 64(2): 291–338, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921302.

Anonymous:1997:AIVc

[910] Anonymous. Author index for volume 64. Journal of Number Theory, 64(2):339, June 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9792117X.

Anonymous:1997:PJc

[911] Anonymous. Pages 1–181 (July 1997). Journal of Number Theory, 65(1): ??, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bayad:1997:VAR

[912] Abdelmejid Bayad. Valuation q-adique et relation de distribution additive pour certaines fonctions q-périodiques. (French) [q-adic valuation and additive distribution relation for certain q-periodic functions]. Journal of Number Theory, 65(1):1–22, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921430.

Fukuda:1997:GCR

[913] Takashi Fukuda. Greenberg’s Conjecture and relative unit groups for real quadratic fields. Journal of Number Theory, 65(1):23–39, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97921260.

Barbacioru:1997:IVP

[914] Cátálin Bárbácioru. Integer-valued polynomials on a subset. Journal of Number Theory, 65(1):40–47, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921247.

Pinner:1997:SFP

[915] Christopher Pinner. On sums of fractional parts nα + γ. Journal of Number Theory, 65(1):48–73, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97920801.

Rieger:1997:IPF

[916] G. J. Rieger. On the integer part function and uniform distribution mod1. Journal of Number Theory, 65(1):74–86, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921326.

Sodaigui:1997:CRP

[917] Bouchaïb Sodaïgui. Classes réalisables par des extensions métacycliques non abéliennes et éléments de Stickelberger. (French) [Classes achievable by non-abelian metacyclic extensions and Stickelberger elements].

Journal of Number Theory, 65(1):87–95, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921338.

Lev:1997:AST

[918] Vsevolod F. Lev. Addendum to “Structure theorem for multiple addition”. Journal of Number Theory, 65(1):96–100, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921284.

See [723].

Chamizo:1997:ODC

[919] Fernando Chamizo and Antonio Córdoba. One-dimensional crystals and quadratic residues. Journal of Number Theory, 65(1):101– 104, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921351.

Sharifi:1997:RGN

[920] Romyar T. Sharifi. Ramification groups of nonabelian Kummer extensions. Journal of Number Theory, 65(1):105–115, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921405.

Ding:1997:CC

[921] Ping Ding. On a conjecture of Chalk. Journal of Number Theory, 65(1):116–129, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921259.

Kargaev:1997:ADD

[922] Pavel Kargaev and Anatoly Zhigljavsky. Asymptotic distribution of the distance function to the Farey points. Journal of Number Theory, 65(1):130–149, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921107.

Johnson:1997:GCA

[923] David H. Johnson and Clifford S. Queen. Genus characters and the arithmetic of commutative monoids. Journal of Number Theory, 65(1): 150–174, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921508.

Byeon:1997:NCN

[924] Dongho Byeon. A note on class numbers of the simplest cubic fields. Journal of Number Theory, 65(1):175–178, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921442.

Anonymous:1997:Eb

[925] Anonymous. Erratum. Journal of Number Theory, 65(1):179, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97921569.

Cai:1997:OSA

[926] Yingchun Cai. “On Orders Solely of Abelian Groups, III” — a correction. Journal of Number Theory, 65(1):180–181, July 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921715. See

[176].

Anonymous:1997:PAb

[927] Anonymous. Pages 183–362 (August 1997). Journal of Number Theory, 65(2):??, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bushnell:1997:UBC

[928] C. J. Bushnell and G. Henniart. An upper bound on conductors for pairs. Journal of Number Theory, 65(2):183–196, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921429.

Girstmair:1997:DVD

[929] Kurt Girstmair. Digit variance and Dedekind sums. Journal of Number Theory, 65(2):197–205, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921491.

Lasjaunias:1997:DAC

[930] Alain Lasjaunias. Diophantine approximation and continued fraction expansions of algebraic power series in positive characteristic. Journal of Number Theory, 65(2):206–225, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921600.

Bassily:1997:NPD

[931] N. L. Bassily, I. Kátai, and M. Wijsmuller. Number of prime divisors of φ_{k}(n), where φ_{k} is the k-fold iterate of φ. Journal of Number Theory, 65(2):226–239, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921648.

Jaulent:1997:PEC

[932] J.-F. Jaulent and O. Sauzet. Pro-ℓ-extensions de corps de nombres lrationnels. (French) [Pro-ℓ-body extensions of l-rational numbers]. Journal of Number Theory, 65(2):240–267, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921582.

Kuhleitner:1997:OTG

[933] Manfred Kühleitner. An Omega theorem on the general asymmetric divisor problem. Journal of Number Theory, 65(2):268–278, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9792151X.

Feng:1997:ARN

[934] Keqin Feng. Anderson’s root numbers and Thakur’s Gauss sums. Journal of Number Theory, 65(2):279–294, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921363.

Yamasaki:1997:BGR

[935] Aiichi Yamasaki. The Brauer group of a rational function field over a perfect field. Journal of Number Theory, 65(2):295–304, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97921557.

Maeda:1997:AFS

[936] Hironobu Maeda. Auflösung der Flächensingularitäten und stabile Reduktion der Fermatkurve. (German) [Resolution of surface singularities and stable reduction of the Fermat curve]. Journal of Number Theory, 65(2):305–315, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921521.

Mao:1997:ASK

[937] Zhengyu Mao. Airy sums, Kloosterman sums, and Salié sums. Journal of Number Theory, 65(2):316–320, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921387.

Lindstrom:1997:BDP

[938] Bernt Lindström. On the binary digits of a power. Journal of Number Theory, 65(2):321–324, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921296.

Li:1997:PSN

[939] Xian-Jin Li. The positivity of a sequence of numbers and the Riemann Hypothesis. Journal of Number Theory, 65(2):325–333, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921375.

Zinoviev:1997:VCG

[940] Dmitrii Zinoviev. On Vinogradov’s constant in Goldbach’s ternary problem. Journal of Number Theory, 65(2):334–358, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921417.

Anonymous:1997:Ec

[941] Anonymous. Erratum. Journal of Number Theory, 65(2):359–360, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921727.

Anonymous:1997:AIVd

[942] Anonymous. Author index for volume 65. Journal of Number Theory, 65(2):361–362, August 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921181.

Anonymous:1997:PS

[943] Anonymous. Pages 1–200 (September 1997). Journal of Number Theory, 66(1):??, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Remy:1997:SAG

[944] Jean-Luc Rémy. Sur la suite autodécrite de Golomb. (French) [On the self-described sequel to Golomb]. Journal of Number Theory, 66(1):1–28, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921545.

Agoh:1997:FQC

[945] Takashi Agoh, Karl Dilcher, and Ladislav Skula. Fermat quotients for composite moduli. Journal of Number Theory, 66(1):29–50, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97921624.

Sander:1997:VDA

[946] J. W. Sander. On the value distribution of arithmetic functions. Journal of Number Theory, 66(1):51–69, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921594.

Wakabayashi:1997:FQT

[947] Isao Wakabayashi. On a family of quartic Thue inequalities I. Journal of Number Theory, 66(1):70–84, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921612.

Bae:1997:CDM

[948] Sunghan Bae and Seungjae Lee. On the coefficients of the Drinfeld modular equation. Journal of Number Theory, 66(1):85–101, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921636.

Barre:1997:MAS

[949] Katia Barré. Mesure d’approximation simultanée de q et J(q). (French) [Simultaneous approximation measure of q and J(q)]. Journal of Number Theory, 66(1):102–128, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921223.

Thakur:1997:PCF

[950] Dinesh S. Thakur. Patterns of continued fractions for the analogues of e and related numbers in the function field case. Journal of Number Theory, 66(1):129–147, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X9792134X.

Car:1997:PT

[951] Mireille Car. Pòlya’s Theorem for F_{q}[T]. Journal of Number Theory, 66(1):148–171, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921454.

Barbolosi:1997:ADT

[952] D. Barbolosi. Une application du théorème ergodique sous-additif à la théorie métrique des fractions continues. (French) [An application of the subadditive ergodic theorem to the metric theory of continuous fractions]. Journal of Number Theory, 66(1):172–182, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921211.

Denis:1997:IAC

[953] Laurent Denis. Indépendance algébrique en caractéristique deux. (French) [Algebraic independence in characteristic two]. Journal of Number Theory, 66(1):183–200, September 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921478.

Anonymous:1997:PO

[954] Anonymous. Pages 201–331 (October 1997). Journal of Number Theory, 66(2):??, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kida:1997:NEC

[955] Masanari Kida and Takaaki Kagawa. Nonexistence of elliptic curves with good reduction everywhere over real quadratic fields. Journal of Number Theory, 66(2):201–210, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921776.

Luczak:1997:ISF

[956] Tomasz Łuczak and Tomasz Schoen. On infinite sum-free sets of natural numbers. Journal of Number Theory, 66(2):211–224, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921697.

Sohne:1997:UBH

[957] Peter Söhne. An upper bound for Hecke zeta-functions with Groessencharacters. Journal of Number Theory, 66(2):225–250, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97921673.

Johansson:1997:TAF

[958] Stefan Johansson. Traces in arithmetic Fuchsian groups. Journal of Number Theory, 66(2):251–270, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921685.

Poulakis:1997:PBS

[959] Dimitrios Poulakis. Polynomial bounds for the solutions of a class of diophantine equations. Journal of Number Theory, 66(2):271–281, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921533.

Bezivin:1997:SRC

[960] Jean-Paul Bezivin. Sur les suites récurrentes à coefficients polynômes. (French) [On recurrent sequences with polynomial coefficients]. Journal of Number Theory, 66(2):282–290, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921570.

Boltje:1997:CGR

[961] Robert Boltje. Class group relations from Burnside ring idempotents. Journal of Number Theory, 66(2):291–305, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9792165X.

Philibert:1997:LZM

[962] Georges Philibert. Un lemme de zéros modulaire. (French) [A modular zeros lemma]. Journal of Number Theory, 66(2):306–313, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921314.

Miyata:1997:IFK

[963] Yoshimasa Miyata. On the invariant factors of Kummer orders in the rings of integers of p-adic number fields of degree p^{2}. Journal of Number Theory, 66(2):314–330, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921703.

Anonymous:1997:AIVe

[964] Anonymous. Author index for volume 66. Journal of Number Theory, 66(2):331, October 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921193.

Anonymous:1997:PN

[965] Anonymous. Pages 1–137 (November 1997). Journal of Number Theory, 67(1):??, November 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Louboutin:1997:CFC

[966] Stéphane Louboutin. CM-fields with cyclic ideal class groups of 2-power orders. Journal of Number Theory, 67(1):1–10, November 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9792179X.

Vulakh:1997:MST

[967] L. Ya Vulakh. The Markov spectra for triangle groups. Journal of Number Theory, 67(1):11–28, November 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X97921818.

Fuhrmann:1997:MC

[968] Rainer Fuhrmann, Arnaldo Garcia, and Fernando Torres. On maximal curves. Journal of Number Theory, 67(1):29–51, November 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X9792148X.

Diamond:1997:SPL

[969] Harold G. Diamond and H. Halberstam. On the sieve parameters α_{κ} and β_{κ} for large κ. Journal of Number Theory, 67(1):52–84, November 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921661.

deDiego:1997:PRF

[970] Teresa de Diego. Points rationnels sur les familles de courbes de genre au moins 2. (French) [Rational points on families of curves of genus at least 2]. Journal of Number Theory, 67(1):85–114, November 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921466.

Yukie:1997:PVS

[971] Akihiko Yukie. Prehomogeneous vector spaces and field extensions III. Journal of Number Theory, 67(1):115–137, November 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9792182X.

Anonymous:1997:PD

[972] Anonymous. Pages 139–291 (December 1997). Journal of Number Theory, 67(2):??, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Srivastav:1997:UQE

[973] A. Srivastav and S. Venkataraman. Unramified quadratic extensions of real quadratic fields, normal integral bases, and 2-adic L-functions. Journal of Number Theory, 67(2):139–145, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921806.

Ferenczi:1997:TNL

[974] Sébastien Ferenczi and Christian Mauduit. Transcendence of numbers with a low complexity expansion. Journal of Number Theory, 67(2):146–161, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921752.

Yokota:1997:NIR

[975] Hisashi Yokota. On number of integers representable as a sum of unit fractions, II. Journal of Number Theory, 67(2):162–169, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921879.

Jeong:1997:RPA

[976] Sangtae Jeong. Rational points on algebraic curves that change genus. Journal of Number Theory, 67(2):170–181, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921764.

Kohnen:1997:NHF

[977] Winfried Kohnen. Nonvanishing of Hecke L-functions associated to cusp forms inside the critical strip. Journal of Number Theory, 67 (2):182–189, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921788.

Baruch:1997:BDA

[978] Ehud Moshe Baruch. On Bessel distributions for GL_{2} over a p-adic field. Journal of Number Theory, 67(2):190–202, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921739.

Agoh:1997:SSP

[979] Takashi Agoh. Stickelberger subideals for a prime modulus related to Kummer-type congruences. Journal of Number Theory, 67(2):203–214, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921831.

Cornelissen:1997:DMF

[980] Gunther Cornelissen. Drinfeld modular forms of weight one. Journal of Number Theory, 67(2):215–228, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921855.

Benjamin:1997:IQF

[981] E. Benjamin, F. Lemmermeyer, and C. Snyder. Imaginary quadratic fields k with cyclic Cl_{2}(k^{1}). Journal of Number Theory, 67(2):229–245, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921740.

Bennett:1997:SFS

[982] Michael A. Bennett. Solving families of simultaneous Pell equations. Journal of Number Theory, 67(2):246–251, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921880.

Cornacchia:1997:AMC

[983] Pietro Cornacchia. Anderson’s module for cyclotomic fields of prime conductor. Journal of Number Theory, 67(2):252–276, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97921843.

Bremner:1997:SIC

[984] Andrew Bremner. Some interesting curves of genus 2 to 7. Journal of Number Theory, 67(2):277–290, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921892.

Anonymous:1997:AIVf

[985] Anonymous. Author index for volume 67. Journal of Number Theory, 67(2):291, December 1997. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9792120X.

Anonymous:1998:PJa

[986] Anonymous. Pages 1–130 (January 1998). Journal of Number Theory, 68(1):??, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Godinho:1998:AZA

[987] Hemar Godinho. On p-adic zeros of additive forms of even degree. Journal of Number Theory, 68(1):1–20, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97922046.

Faivre:1998:RCA

[988] C. Faivre. The rate of convergence of approximations of a continued fraction. Journal of Number Theory, 68(1):21–28, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921867.

Aiba:1998:CMG

[989] Akira Aiba. Carlitz modules and Galois module structure II. Journal of Number Theory, 68(1):29–35, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921971.

Ichimura:1998:LUM

[990] Humio Ichimura. Local units modulo Gauss sums. Journal of Number Theory, 68(1):36–56, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9792206X.

Alzer:1998:RA

[991] Horst Alzer. On rational approximation to e. Journal of Number Theory, 68(1):57–62, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921995.

Ruzsa:1998:ISS

[992] Imre Z. Ruzsa. An infinite Sidon sequence. Journal of Number Theory, 68(1):63–71, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97921922.

Bulant:1998:PCN

[993] Michal Bulant. On the parity of the class number of the field Q(p, q, r). Journal of Number Theory, 68(1):72–86, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921909.

Warlimont:1998:NNS

[994] Richard Warlimont. On natural numbers as sums of consecutive h-th powers. Journal of Number Theory, 68(1):87–98, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97921946.

Blackburn:1998:OSP

[995] Simon R. Blackburn. Orthogonal sequences of polynomials over arbitrary fields. Journal of Number Theory, 68(1):99–111, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97922010.

Yangbo:1998:EST

[996] Ye Yangbo. Exponential sums for GL(n) and their applications to base change. Journal of Number Theory, 68(1):112–130, January 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X9792196X.

Anonymous:1998:PF

[997] Anonymous. Pages 131–243 (February 1998). Journal of Number Theory, 68(2):??, February 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Alexandru:1998:CS

[998] Victor Alexandru, Nicolae Popescu, and Alexandru Zaharescu. On the closed subfields of C_{p}. Journal of Number Theory, 68(2):131–150, February 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97921983.

Calkin:1998:CGS

[999] Neil J. Calkin and Jan McDonald Thomson. Counting generalized sum-free sets. Journal of Number Theory, 68(2):151–159, February 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97921910.

Choie:1998:RCO

[1000] YoungJu Choie and Wolfgang Eholzer. Rankin–Cohen operators for Jacobi and Siegel forms. Journal of Number Theory, 68(2):160–177, February 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922034.

Huang:1998:MRG

[1001] Sen-Shan Huang. On modular relations for the Göllnitz–Gordon functions with applications to partitions. Journal of Number Theory, 68 (2):178–216, February 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97922058.

Cerveau:1998:TFS

[1002] D. Cerveau and F. Loray. Un théorème de Frobenius singulier via l’arithmétique élémentaire. (French) [A singular Frobenius theorem via elementary arithmetic]. Journal of Number Theory, 68(2):217–228, February 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922022.

Hazama:1998:DFC

[1003] Fumio Hazama. Determinantal formula for the cuspidal class number of the modular curve X_{1}(m). Journal of Number Theory, 68(2):229– 242, February 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922101.

Anonymous:1998:AIVa

[1004] Anonymous. Author index for volume 68. Journal of Number Theory, 68(2):243, February 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922600.

Anonymous:1998:PMa

[1005] Anonymous. Pages 1–124 (March 1998). Journal of Number Theory, 69 (1):??, March 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Greither:1998:EVI

[1006] Cornelius Greither. An equivariant version of an index formula of Fröhlich and McCulloh. Journal of Number Theory, 69(1):1–15, March 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922095.

Belliard:1998:SGU

[1007] Jean-Robert Belliard. Sur la structure galoisienne des unités circulaires dans les Z_{p}-extensions. (French) [On the Galois structure of circular units in the Z_{p}-extensions]. Journal of Number Theory, 69(1): 16–49, March 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922009.

Washington:1998:AFS

[1008] Lawrence C. Washington. p-adic L-functions and sums of powers. Journal of Number Theory, 69(1):50–61, March 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921958.

Stevenson:1998:CRP

[1009] Katherine F. Stevenson. Conditions related to π_{1} of projective curves. Journal of Number Theory, 69(1):62–79, March 1998. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97921934.

Hsia:1998:SMD

[1010] Liang-Chung Hsia and Jing Yu. On singular moduli of Drinfeld modules in characteristic two. Journal of Number Theory, 69(1):80–97, March 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922113.

Robertson:1998:PBC

[1011] Leanne Robertson. Power bases for cyclotomic integer rings. Journal of Number Theory, 69(1):98–118, March 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97922149.

Allouche:1998:AIB

[1012] Jean-Paul Allouche and Luca Q. Zamboni. Algebraic irrational binary numbers cannot be fixed points of non-trivial constant length or primitive morphisms. Journal of Number Theory, 69(1):119–124, March 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922071.

Anonymous:1998:PAa

[1013] Anonymous. Pages 125–231 (April 1998). Journal of Number Theory, 69(2):??, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Shen:1998:IRB

[1014] Li-Chien Shen. On an identity of Ramanujan based on the hypergeometric series _{2}F_{1}(13, 23; 12; x). Journal of Number Theory, 69(2):125– 134, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922125.

Hwang:1998:RVF

[1015] Hsien-Kuei Hwang. Sur la répartition des valeurs des fonctions arithmétiques. Le nombre de facteurs premiers d’un entier. (French) [On the distribution of the values of the arithmetic functions. The number of prime factors of an integer]. Journal of Number Theory, 69(2):135–152, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X97922162.

Alladi:1998:PTG

[1016] Krishnaswami Alladi. On a partition theorem of Göllnitz and quartic transformations. Journal of Number Theory, 69(2):153–180, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X97922150.

Wang:1998:CNR

[1017] Julie T.-Y Wang and Jing Yu. On class number relations over function fields. Journal of Number Theory, 69(2):181–196, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97922137.

Lindstrom:1998:WDS

[1018] Bernt Lindström. Well distribution of Sidon sets in residue classes. Journal of Number Theory, 69(2):197–200, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X97922174.

Ritter:1998:LDP

[1019] Stefan Matthias Ritter. On a linear diophantine problem of Frobenius: extending the basis. Journal of Number Theory, 69(2):201– 212, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922198.

Daberkow:1998:CHC

[1020] M. Daberkow and M. Pohst. On the computation of Hilbert class fields. Journal of Number Theory, 69(2):213–230, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X97922083.

Anonymous:1998:AIVb

[1021] Anonymous. Author index for volume 69. Journal of Number Theory, 69 (2):231, April 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922617.

Anonymous:1998:PMb

[1022] Anonymous. Pages 1–98 (May 1998). Journal of Number Theory, 70(1): ??, May 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Balakrishnan:1998:AEC

[1023] U. Balakrishnan and Y.-F. S. Pétermann. Asymptotic estimates for a class of summatory functions. Journal of Number Theory, 70(1): 1–36, May 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X97922186.

Ose:1998:DFF

[1024] David T. Ose. Deformations for function fields. Journal of Number Theory, 70(1):37–61, May 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922247.

Cochrane:1998:SSL

[1025] Todd Cochrane and Patrick Mitchell. Small solutions of the Legendre equation. Journal of Number Theory, 70(1):62–66, May 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922211.

Jiang:1998:FTI

[1026] Dihua Jiang. The first term identities for Eisenstein series. Journal of Number Theory, 70(1):67–98, May 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X98922235.

Anonymous:1998:PJb

[1027] Anonymous. Pages 99–191 (June 1998). Journal of Number Theory, 70 (2):??, June 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Erdos:1998:API

[1028] Paul Erdős, Christian Mauduit, and András Sárközy. On arithmetic properties of integers with missing digits I: distribution in residue classes. Journal of Number Theory, 70(2):99–120, June 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922296.

Sze:1998:CMS

[1029] Lawrence Sze. Calculating p(n) modulo small primes using quadratic forms. Journal of Number Theory, 70(2):121–126, June 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922302.

Bluher:1998:RBC

[1030] Antonia Wilson Bluher. Relations between certain power sums of elliptic modular forms in characteristic two. Journal of Number Theory, 70(2):127–145, June 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922260.

Angles:1998:CGP

[1031] Bruno Angles. On the class group problem for function fields. Journal of Number Theory, 70(2):146–159, June 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922259.

Yukie:1998:PVS

[1032] Akihiko Yukie. Prehomogeneous vector spaces and ergodic theory III. Journal of Number Theory, 70(2):160–183, June 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922314.

Kuhnlein:1998:TCC

[1033] Stefan Kühnlein. Torsion classes in cohomology and Galois representations. Journal of Number Theory, 70(2):184–190, June 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922284.

Anonymous:1998:AIVc

[1034] Anonymous. Author index for volume 70. Journal of Number Theory, 70(2):191, June 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922624.

Anonymous:1998:PJc

[1035] Anonymous. Pages 1–157 (July 1998). Journal of Number Theory, 71(1): ??, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Ward:1998:FMS

[1036] Thomas Ward. A family of Markov shifts (almost) classified by periodic points. Journal of Number Theory, 71(1):1–11, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922429.

Eliahou:1998:SVS

[1037] Shalom Eliahou and Michel Kervaire. Sumsets in vector spaces over finite fields. Journal of Number Theory, 71(1):12–39, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922351.

Hironaka:1998:LZF

[1038] Yumiko Hironaka. Local zeta functions on Hermitian forms and its application to local densities. Journal of Number Theory, 71(1):40– 64, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922375.

Levesley:1998:GIJ

[1039] J. Levesley. A general inhomogeneous Jarnik–Besicovitch Theorem. Journal of Number Theory, 71(1):65–80, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922399.

Chan:1998:RCD

[1040] Wai Kiu Chan, Dennis R. Estes, and M. Jöchner. Representations of codimension ≥ 3 by definite quadratic forms. Journal of Number Theory, 71(1):81–85, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922338.

Dummigan:1998:CCT

[1041] Neil Dummigan and Pham Huu Tiep. Congruences for certain theta series. Journal of Number Theory, 71(1):86–105, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9892234X.

Voloch:1998:DOI

[1042] José Felipe Voloch. Differential operators and interpolation series in power series fields. Journal of Number Theory, 71(1):106–108, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98922417.

Evans:1998:CJS

[1043] Ronald Evans. Congruences for Jacobi sums. Journal of Number Theory, 71(1):109–120, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922363.

Sheats:1998:RHG

[1044] Jeffrey T. Sheats. The Riemann Hypothesis for the Goss zeta function for F_{q}[T]. Journal of Number Theory, 71(1):121–157, July 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922326.

Anonymous:1998:PAb

[1045] Anonymous. Pages 159–312 (August 1998). Journal of Number Theory, 71(2):??, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Gupta:1998:RSM

[1046] Shamita Dutta Gupta and Xiaotie She. The GL(2) Rankin–Selberg method for functions not of rapid decay. Journal of Number Theory, 71(2):159–165, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922491.

Rynne:1998:HDS

[1047] Bryan P. Rynne. The Hausdorff dimension of sets arising from diophantine approximation with a general error function. Journal of Number Theory, 71(2):166–171, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X98922533.

Stohr:1998:LGZ

[1048] Karl-Otto Stöhr. Local and global zeta-functions of singular algebraic curves. Journal of Number Theory, 71(2):172–202, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922405.

Gouvea:1998:OFM

[1049] Fernando Q. Gouvêa and Noriko Yui. Observations on Fermat motives of K3-type. Journal of Number Theory, 71(2):203–226, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922442.

Bugeaud:1998:BEP

[1050] Yann Bugeaud. Bornes effectives pour les solutions des équations en S-unités et des équations de Thue–Mahler. (French) [Effective bounds for solutions of equations in S-units and Thue-Mahler equations]. Journal of Number Theory, 71(2):227–244, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922454.

Helou:1998:PRB

[1051] Charles Helou. Power reciprocity for binomial cyclotomic integers. Journal of Number Theory, 71(2):245–256, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922508.

Hsu:1998:CP

[1052] Chih-Nung Hsu. On Carmichael polynomials. Journal of Number Theory, 71(2):257–274, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922272.

Didier:1998:CRF

[1053] Gilles Didier. Codages de rotations et fractions continues. (French) [Coding of rotations and continuous fractions]. Journal of Number Theory, 71(2):275–306, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922466.

Grekos:1998:OMA

[1054] Georges Grekos. On the order of a minimal additive basis. Journal of Number Theory, 71(2):307–311, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X9892248X.

Anonymous:1998:AIVd

[1055] Anonymous. Author index for volume 71. Journal of Number Theory, 71(2):312, August 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922636.

Anonymous:1998:PS

[1056] Anonymous. Pages 1–152 (September 1998). Journal of Number Theory, 72(1):??, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Zannier:1998:NTR

[1057] Umberto Zannier. On the number of times a root of f(n, x) = 0 generates a field containing a given number field. Journal of Number Theory, 72(1):1–12, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922430.

Gonzalez:1998:REP

[1058] Josep González and Joan-C. Lario. Rational and elliptic parametrizations of Q-curves. Journal of Number Theory, 72(1):13–31, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98922594.

Grantham:1998:PPT

[1059] Jon Grantham. A probable prime test with high confidence. Journal of Number Theory, 72(1):32–47, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922478.

Stolin:1998:PGI

[1060] Alexander Stolin. On the Picard group of the integer group ring of the cyclic p-group and certain Galois groups. Journal of Number Theory, 72(1):48–66, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922545.

Bhargava:1998:GFF

[1061] Manjul Bhargava. Generalized factorials and fixed divisors over subsets of a Dedekind domain. Journal of Number Theory, 72(1):67–75, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9892220X.

Kim:1998:HFN

[1062] Jun Kyo Kim. On highly factorable numbers. Journal of Number Theory, 72(1):76–91, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922387.

Ramakrishna:1998:DER

[1063] Ravi Ramakrishna. Deforming an even representation II, raising the level. Journal of Number Theory, 72(1):92–109, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922521.

Kontogeorgis:1998:GAF

[1064] Aristides I. Kontogeorgis. The group of automorphisms of the function fields of the curve x^{n} + y^{m} + 1 = 0. Journal of Number Theory, 72(1):110–136, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9892251X.

Sury:1998:CCA

[1065] B. Sury. On a conjecture of Chowla et al. Journal of Number Theory, 72(1):137–139, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922557.

Ichimura:1998:CNM

[1066] Humio Ichimura. On the class numbers of the maximal real subfields of cyclotomic function fields, II. Journal of Number Theory, 72 (1):140–149, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922727.

Yokota:1998:C

[1067] Hisashi Yokota. Corrigendum. Journal of Number Theory, 72(1):150, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922934.

Hazama:1998:C

[1068] Fumio Hazama. Corrigendum. Journal of Number Theory, 72(1):151– 152, September 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922971.

Anonymous:1998:PO

[1069] Anonymous. Pages 153–377 (October 1998). Journal of Number Theory, 72(2):??, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Caldeira:1998:PTR

[1070] Cristina Caldeira and J. A. Dias da Silva. A Pollard type result for restricted sums. Journal of Number Theory, 72(2):153–173, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922697.

Laubie:1998:RSA

[1071] F. Laubie and M. Saïne. Ramification of some automorphisms of local fields. Journal of Number Theory, 72(2):174–182, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922739.

Vallee:1998:DFC

[1072] Brigitte Vallée. Dynamique des fractions continues à contraintes périodiques. Journal of Number Theory, 72(2):183–235, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98922764.

Hsia:1998:CMP

[1073] Liang chung Hsia. On the coefficients of modular polynomials for Drinfeld modules. Journal of Number Theory, 72(2):236–256, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922715.

vanderGeer:1998:GRM

[1074] G. van der Geer and M. van der Vlugt. Generalized Reed–Muller codes and curves with many points. Journal of Number Theory, 72(2):257– 268, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922776.

Schmidt:1998:CIP

[1075] Bernhard Schmidt. Cyclotomic integers of prescribed absolute value and the class group. Journal of Number Theory, 72(2):269–281, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922752.

Cohen:1998:MIF

[1076] Arjeh M. Cohen, Gabriele Nebe, and Wilhelm Plesken. Maximal integral forms of the algebraic group G_{2} defined by finite subgroups. Journal of Number Theory, 72(2):282–308, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922685.

Yang:1998:EFL

[1077] Tonghai Yang. An explicit formula for local densities of quadratic forms. Journal of Number Theory, 72(2):309–356, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922582.

Conway:1998:NOU

[1078] J. H. Conway and N. J. A. Sloane. A note on optimal unimodular lattices. Journal of Number Theory, 72(2):357–362, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922570.

Czogala:1998:HDH

[1079] Alfred Czogała and Andrzej Sładek. Higher degree Hilbert symbol equivalence of algebraic number fields, II. Journal of Number Theory, 72(2):363–376, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922661.

Anonymous:1998:AIVe

[1080] Anonymous. Author index for volume 72. Journal of Number Theory, 72(2):377, October 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922648.

Anonymous:1998:PN

[1081] Anonymous. Pages 1–138 (November 1998). Journal of Number Theory, 73(1):??, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Elsner:1998:EHI

[1082] C. Elsner and J. W. Sander. On the exact height of integerdetecting sequences. Journal of Number Theory, 73(1):1–13, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X9892282X.

McDaniel:1998:SCL

[1083] Wayne L. McDaniel and Paulo Ribenboim. Square-classes in Lucas sequences having odd parameters. Journal of Number Theory, 73(1):14–27, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922806.

Barbacioru:1998:EFI

[1084] Catalin Barbacioru. Extension fields and integer-valued polynomials. Journal of Number Theory, 73(1):28–33, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922673.

Goto:1998:TAV

[1085] Keiji Goto. A twisted adjoint l-value of an elliptic modular form. Journal of Number Theory, 73(1):34–46, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922703.

Adelberg:1998:ACN

[1086] Arnold Adelberg. 2-adic congruences of Nörlund numbers and of Bernoulli numbers of the second kind. Journal of Number Theory, 73(1):47–58, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922788.

Bhowmik:1998:TKI

[1087] Gautami Bhowmik and Olivier Ramaré. A Turán–Kubilius inequality for integer matrices. Journal of Number Theory, 73(1):59–71, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922818.

Bachman:1998:SRN

[1088] Gennady Bachman. Some remarks on nonnegative multiplicative functions on arithmetic progressions. Journal of Number Theory, 73(1):72– 91, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9892279X.

Liardet:1998:ACC

[1089] Pierre Liardet and Pierre Stambul. Algebraic computations with continued fractions. Journal of Number Theory, 73(1):92–121, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98922740.

Zieve:1998:ND

[1090] Michael Zieve. A Note on the Discriminator. Journal of Number Theory, 73(1):122–138, November 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922569.

Anonymous:1998:PD

[1091] Anonymous. Pages 139–534 (December 1998). Journal of Number Theory, 73(2):??, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Prevost:1998:I

[1092] M. Prevost. On the irrationality of ∑ t^{n}Aα^{n} + Bβ^{n}. Journal of Number Theory, 73(2):139–161, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X98923046.

Mokrane:1998:QRO

[1093] A. Mokrane. Quelques remarques sur l’ordinarité. (French) [Some remarks on ordinariness]. Journal of Number Theory, 73(2):162–181, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922879.

Benjamin:1998:RQF

[1094] Elliot Benjamin, Franz Lemmermeyer, and C. Snyder. Real quadratic fields with abelian 2-class field tower. Journal of Number Theory, 73(2):182–194, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922910.

Voloch:1998:EC

[1095] José Felipe Voloch. The equation ax + by = 1 in characteristic p. Journal of Number Theory, 73(2):195–200, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922958.

Queffelec:1998:TFC

[1096] M. Queffélec. Transcendance des fractions continues de Thue–Morse. (French) [Transcendence of Thue–Morse continuous fractions]. Journal of Number Theory, 73(2):201–211, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923058.

Viader:1998:NLM

[1097] Pelegrí Viader, Jaume Paradís, and Lluís Bibiloni. A new light on Minkowski’s ?(x) function. Journal of Number Theory, 73(2):212–227, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922946.

Dabrowski:1998:KSR

[1098] Romuald Dabrowski and Mark Reeder. Kloosterman sets in reductive groups. Journal of Number Theory, 73(2):228–255, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923010.

Mauduit:1998:FPB

[1099] Christian Mauduit and András Sárközy. On finite pseudorandom binary sequences: II. The Champernowne, Rudin–Shapiro, and Thue– Morse sequences, a further construction. Journal of Number Theory, 73(2):256–276, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922867.

Treatman:1998:ES

[1100] Stefan G. Treatman. Euclidean systems. Journal of Number Theory, 73(2):277–291, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922843.

Nicolas:1998:PAR

[1101] J.-L Nicolas, I. Z. Ruzsa, and A. Sárközy. On the parity of additive representation functions. Journal of Number Theory, 73(2):292–317, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922880.

Chang:1998:CNP

[1102] Ku-Young Chang and Soun-Hi Kwon. Class number problem for imaginary cyclic number fields. Journal of Number Theory, 73(2):318–338, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98922983.

Yanagisawa:1998:AFC

[1103] Naoki Yanagisawa. An asymptotic formula for a certain mean value in a divisor problem. Journal of Number Theory, 73(2):339–358, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9892296X.

Rains:1998:STM

[1104] E. M. Rains and N. J. A. Sloane. The shadow theory of modular and unimodular lattices. Journal of Number Theory, 73(2):359–389, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9892306X.

Lesage:1998:DPC

[1105] Jean-Luc Lesage. Différence entre puissances et carrés d’entiers. (French) [Difference between powers and squares of integers]. Journal of Number Theory, 73(2):390–425, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X98923034.

DiPippo:1998:RPA

[1106] Stephen A. DiPippo and Everett W. Howe. Real polynomials with all roots on the unit circle and abelian varieties over finite fields. Journal of Number Theory, 73(2):426–450, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923022. See

corrigendum [1252].

Yang:1998:LAF

[1107] Zifeng Yang. Locally analytic functions over completions of F_{r}[U]. Journal of Number Theory, 73(2):451–458, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98923083.

Cornacchia:1998:FIC

[1108] Pietro Cornacchia and Cornelius Greither. Fitting ideals of class groups of real fields with prime power conductor. Journal of Number Theory, 73(2):459–471, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923009.

Summerer:1998:CFS

[1109] Leonhard Summerer. Cubic forms as sum of cubes of linear forms. Journal of Number Theory, 73(2):472–517, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98922922.

Sands:1998:BCH

[1110] Jonathan W. Sands. Base change for higher Stickelberger ideals. Journal of Number Theory, 73(2):518–526, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98922831.

James:1998:NIH

[1111] Kevin James and Ken Ono. A note on the irreducibility of Hecke polynomials. Journal of Number Theory, 73(2):527–532, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98922855.

Anonymous:1998:AIVf

[1112] Anonymous. Author index for volume 73. Journal of Number Theory, 73(2):533–534, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9892265X.

Anonymous:1999:PJa

[1113] Anonymous. Pages 1–171 (January 1999). Journal of Number Theory, 74(1):??, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Wong:1999:TGR

[1114] Siman Wong. Twists of Galois representations and projective automorphisms. Journal of Number Theory, 74(1):1–18, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98923071.

Ribenboim:1999:ADP

[1115] Paulo Ribenboim. An algorithm to determine the points with integral coordinates in certain elliptic curves. Journal of Number Theory, 74(1):19–38, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922909.

Ohno:1999:GDS

[1116] Yasuo Ohno. A generalization of the duality and sum formulas on the multiple zeta values. Journal of Number Theory, 74(1):39–43, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923149.

Ducos:1999:RRE

[1117] Lionel Ducos. Réalisation régulière explicite des groupes abéliens finis comme groupes de Galois. (French) [Explicit regular realization of finite abbreviated groups as Galois groups]. Journal of Number Theory, 74(1):44–55, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923101.

Lauter:1999:FCC

[1118] Kristin Lauter. A formula for constructing curves over finite fields with many rational points. Journal of Number Theory, 74(1):56–72, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923125.

Prasad:1999:SRR

[1119] Dipendra Prasad. Some remarks on representations of a division algebra and of the Galois group of a local field. Journal of Number Theory, 74(1):73–97, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922892.

Mollin:1999:PAI

[1120] R. A. Mollin and K. Cheng. Palindromy and ambiguous ideals revisited. Journal of Number Theory, 74(1):98–110, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98923137.

Thuswaldner:1999:SDF

[1121] Jörg M. Thuswaldner. The sum of digits function in number fields: distribution in residue classes. Journal of Number Theory, 74(1):111– 125, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923174.

Chen:1999:VOP

[1122] Yong-Gao Chen and T. W. Cusick. The view-obstruction problem for n-dimensional cubes. Journal of Number Theory, 74(1):126–133, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923095.

Ribenboim:1999:ACP

[1123] Paulo Ribenboim and Gary Walsh. The ABC Conjecture and the powerful part of terms in binary recurring sequences. Journal of Number Theory, 74(1):134–147, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923150.

Hosokawa:1999:ILZ

[1124] Hiroshi Hosokawa. Igusa local zeta functions and parabolic castling transformation of prehomogeneous vector spaces. Journal of Number Theory, 74(1):148–171, January 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923113.

Anonymous:1999:PF

[1125] Anonymous. Pages 173–350 (February 1999). Journal of Number Theory, 74(2):??, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Satoh:1999:SPT

[1126] Junya Satoh. Sums of products of two q-Bernoulli numbers. Journal of Number Theory, 74(2):173–180, February 1999. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923319.

Li:1999:CSA

[1127] Wen-Ching Winnie Li. Character Sums over p-adic Fields. Journal of Number Theory, 74(2):181–229, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923289.

Wielonsky:1999:HPA

[1128] Franck Wielonsky. Hermite–Padé approximants to exponential functions and an inequality of Mahler. Journal of Number Theory, 74 (2):230–249, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923344.

Coleman:1999:HHC

[1129] M. D. Coleman. The Hooley–Huxley contour method for problems in number fields: I. Arithmetic functions. Journal of Number Theory, 74(2):250–277, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98922995.

Chen:1999:SMC

[1130] Imin Chen. On Siegel’s modular curve of level 5 and the class number one problem. Journal of Number Theory, 74(2):278–297, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923204.

Lopez:1999:EPF

[1131] Bartolomé López and Bodo Wack. On the expansion of a product formula related with the Drinfeld discriminant function. Journal of Number Theory, 74(2):298–306, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923290.

Kim:1999:JDA

[1132] Dong-Hyun Kim. On the joint distribution of q-additive functions in residue classes. Journal of Number Theory, 74(2):307–336, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923277.

Taguchi:1999:FIC

[1133] Yuichiro Taguchi. Finiteness of an isogeny class of Drinfeld modules — correction to a previous paper. Journal of Number Theory, 74(2):337–348, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923162. See [354].

Anonymous:1999:E

[1134] Anonymous. Erratum. Journal of Number Theory, 74(2):349, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X9892371X.

Anonymous:1999:AIVa

[1135] Anonymous. Author index for volume 74. Journal of Number Theory, 74(2):350, February 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99923727.

Anonymous:1999:PMa

[1136] Anonymous. Pages 1–159 (March 1999). Journal of Number Theory, 75 (1):??, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Erdos:1999:SNM

[1137] Paul Erdős and Hugh L. Montgomery. Sums of numbers with many divisors. Journal of Number Theory, 75(1):1–6, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9892323X.

Godbole:1999:RSS

[1138] Anant P. Godbole, Svante Janson, Nicholas W. Locantore, and Rebecca Rapoport. Random Sidon sequences. Journal of Number Theory, 75(1):7–22, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923253.

Tateyama:1999:CFD

[1139] Koichi Tateyama. Continuous functions on discrete valuation rings. Journal of Number Theory, 75(1):23–33, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923332.

Choudhry:1999:QDE

[1140] Ajai Choudhry. On the quartic diophantine equation f(x, y) = f(u, v). Journal of Number Theory, 75(1):34–40, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923228.

Luo:1999:BIH

[1141] Wenzhi Luo. Bounds for incomplete hyper-Kloosterman sums. Journal of Number Theory, 75(1):41–46, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X9892340X.

Brueggeman:1999:NCG

[1142] Sharon Brueggeman. The nonexistence of certain Galois extensions unramified outside 5. Journal of Number Theory, 75(1):47–52, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923186.

Solomon:1999:SCI

[1143] David Solomon. The Shintani cocycle: II. Partial ζ-functions, cohomologous cocycles and p-adic interpolation. Journal of Number Theory, 75(1):53–108, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923320.

Mazur:1999:FAS

[1144] Marcin Mazur. Finite arithmetic subgroups of GL_{n}. Journal of Number Theory, 75(1):109–119, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923307.

Cangelmi:1999:RAC

[1145] Leonardo Cangelmi and Francesco Pappalardi. On the r-rank Artin Conjecture, II. Journal of Number Theory, 75(1):120–132, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923198.

Escassut:1999:UUN

[1146] Alain Escassut, Labib Haddad, and Robert Vidal. Urs, Ursim, and Non–Urs for p-adic Functions and Polynomials. Journal of Number Theory, 75(1):133–144, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923241.

Chou:1999:PDS

[1147] Wun-Seng Chou and Stephen D. Cohen. Polynomial distribution and sequences of irreducible polynomials over finite fields. Journal of Number Theory, 75(1):145–159, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923216.

Anonymous:1999:PAa

[1148] Anonymous. Pages 161–334 (April 1999). Journal of Number Theory, 75(2):??, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Boston:1999:SCF

[1149] Nigel Boston. Some cases of the Fontaine–Mazur conjecture, II. Journal of Number Theory, 75(2):161–169, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X9892337X.

Taya:1999:AFE

[1150] Hisao Taya. On p-adic L-functions and Z_{p}-extensions of certain real Abelian number fields. Journal of Number Theory, 75(2):170– 184, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923265.

Doud:1999:ERO

[1151] Darrin Doud. S_{4} and S_{4} extensions of Q ramified at only one prime. Journal of Number Theory, 75(2):185–197, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923459.

Gunnells:1999:MSR

[1152] Paul E. Gunnells. Modular symbols for Q-rank one groups and Voronoı reduction. Journal of Number Theory, 75(2):198–219, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923472.

Alladi:1999:QKI

[1153] Krishnaswami Alladi and George E. Andrews. A quartic key identity for a partition theorem of Göllnitz. Journal of Number Theory, 75(2):220–236, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923356.

Stern:1999:OHN

[1154] Leonid Stern. On an obstruction to the Hasse norm principle and the equality of norm groups of algebraic number fields. Journal of Number Theory, 75(2):237–261, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923423.

Matsumoto:1999:ZHD

[1155] Kohji Matsumoto and Yoshio Tanigawa. On the zeros of higher derivatives of Hardy’s Z-function. Journal of Number Theory, 75(2):262– 278, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923411.

Bosser:1999:MFL

[1156] Vincent Bosser. Minorations de formes linéaires de logarithmes pour les modules de Drinfeld. (French) [Minimization of linear forms of logarithms for Drinfeld modules]. Journal of Number Theory, 75(2):279– 323, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923368.

Hazama:1999:DFS

[1157] Fumio Hazama. Determinantal formula for the special values of the Dedekind zeta function of the cyclotomic field. Journal of Number Theory, 75(2):324–333, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923381.

Anonymous:1999:AIVb

[1158] Anonymous. Author index for volume 75. Journal of Number Theory, 75 (2):334, April 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99921169.

Anonymous:1999:PMb

[1159] Anonymous. Pages 1–165 (May 1999). Journal of Number Theory, 76(1): ??, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

DiazyDiaz:1999:AAD

[1160] F. Diaz y Diaz and F. Soriano. Approche algorithmique du groupe des classes logarithmiques. (French) [Algorithmic approach to the group of logarithmic classes]. Journal of Number Theory, 76(1):1– 15, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923447.

James:1999:ECS

[1161] Kevin James. Elliptic curves satisfying the Birch and Swinnerton-Dyer Conjecture mod3. Journal of Number Theory, 76(1):16–21, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923496.

Kuhleitner:1999:DTP

[1162] M. Kühleitner. On differences of two k-th powers: an asymptotic formula for the mean-square of the error term. Journal of Number Theory, 76(1):22–44, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923526.

Heuberger:1999:FPT

[1163] Clemens Heuberger. On families of parametrized Thue equations. Journal of Number Theory, 76(1):45–61, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923484.

Ono:1999:CSE

[1164] Ken Ono. On the circular summation of the eleventh powers of Ramanujan’s theta function. Journal of Number Theory, 76(1):62– 65, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9892354X.

Raouj:1999:PDE

[1165] A. Raouj and A. Stef. Sur la proximité des diviseurs des entiers. (French) [On the proximity of integer divisors]. Journal of Number Theory, 76(1):66–93, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923551.

Sert:1999:VED

[1166] Alain Sert. Une version effective du théorème de Lindemann–Weierstrass par les déterminants d’interpolation. (French) [An effective version of the Lindemann–Weierstrass theorem by interpolation determinants]. Journal of Number Theory, 76(1):94–119, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923563.

Kable:1999:PBD

[1167] Anthony C. Kable. Power bases in dihedral quartic fields. Journal of Number Theory, 76(1):120–129, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923502.

Kulesz:1999:AMD

[1168] Leopoldo Kulesz. Application de la méthode de Dem’janenko–Manin à certaines familles de courbes de genre 2 et 3. (French) [Application of the Dem’janenko–Manin method to certain families of genus curves 2 and 3]. Journal of Number Theory, 76(1):130–146, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98923393.

Kolountzakis:1999:UDR

[1169] Mihail N. Kolountzakis. On the uniform distribution in residue classes of dense sets of integers with distinct sums. Journal of Number Theory, 76(1):147–153, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923514.

Bars:1999:BMC

[1170] Francesc Bars. Bielliptic modular curves. Journal of Number Theory, 76(1):154–165, May 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923435.

Anonymous:1999:PJb

[1171] Anonymous. Pages 167–337 (June 1999). Journal of Number Theory, 76 (2):??, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Benjamin:1999:RLB

[1172] Elliot Benjamin and Charles J. Parry. Refined lower bounds on the 2-class number of the Hilbert 2-class field of imaginary quadratic number fields with elementary 2-class group of rank 3. Journal of Number Theory, 76(2):167–177, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923617.

Huang:1999:BNP

[1173] I-Chiau Huang and Su-Yun Huang. Bernoulli numbers and polynomials via residues. Journal of Number Theory, 76(2):178–193, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923642.

Bocherer:1999:SAF

[1174] Siegfried Böcherer and Rainer Schulze-Pillot. Squares of automorphic forms on quaternion algebras and central values of L-functions of modular forms. Journal of Number Theory, 76(2):194–205, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98923629.

Yokota:1999:LIE

[1175] Hisashi Yokota. The largest integer expressible as a sum of reciprocal of integers. Journal of Number Theory, 76(2):206–216, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923599. See erratum [1253].

Lubinsky:1999:SUC

[1176] D. S. Lubinsky. The size of (q; q)_{n} for q on the unit circle. Journal of Number Theory, 76(2):217–247, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923654.

Vazzana:1999:RRI

[1177] A. Vazzana. 8-ranks of K_{2} of rings of integers in quadratic number fields. Journal of Number Theory, 76(2):248–264, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X98923587.

Chan:1999:QEP

[1178] Wai Kiu Chan. Quaternary even positive definite quadratic forms of discriminant 4p. Journal of Number Theory, 76(2):265–280, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923630.

Kochubei:1999:LCF

[1179] Anatoly N. Kochubei. F_{q}-linear calculus over function fields. Journal of Number Theory, 76(2):281–300, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923605.

Thakur:1999:AAS

[1180] Dinesh S. Thakur. An alternate approach to solitons for F_{q}[t]. Journal of Number Theory, 76(2):301–319, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923575.

Kim:1999:AAL

[1181] Taekyun Kim. On a q-analogue of the p-adic log gamma functions and related integrals. Journal of Number Theory, 76(2):320– 329, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99923739.

Lagarias:1999:HBA

[1182] Jeffrey C. Lagarias and Yang Wang. Haar bases for L^{2}(R^{n}) and algebraic number theory. Journal of Number Theory, 76(2):330– 336, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X98923538.

Anonymous:1999:AIVc

[1183] Anonymous. Author index for volume 76. Journal of Number Theory, 76(2):337, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99922606.

Anonymous:1999:PJc

[1184] Anonymous. Pages 1–169 (July 1999). Journal of Number Theory, 77(1): ??, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Grabitz:1999:CHO

[1185] Martin Grabitz. Continuation of hereditary orders in local central simple algebras. Journal of Number Theory, 77(1):1–26, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99923740.

Perret:1999:ICG

[1186] Marc Perret. On the ideal class group problem for global fields. Journal of Number Theory, 77(1):27–35, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X98923666.

Aubry:1999:IBB

[1187] Yves Aubry and Dominique Le Brigand. Imaginary bicyclic biquadratic function fields in characteristic two. Journal of Number Theory, 77(1): 36–50, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99923776.

Zink:1999:MEL

[1188] Ernst Wilhelm Zink. More on embeddings of local fields in simple algebras. Journal of Number Theory, 77(1):51–61, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99923752.

Ruppert:1999:RPM

[1189] Wolfgang M. Ruppert. Reducibility of polynomials f(x, y) modulo p. Journal of Number Theory, 77(1):62–70, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99923818.

Sadi:1999:DED

[1190] Bounab Sadi. Descente effective du corps de définition des revêtements galoisiens. (French) [Effective descent of the body defining the Galois coatings]. Journal of Number Theory, 77(1):71–82, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9892368X.

Efrat:1999:FGP

[1191] Ido Efrat. Finitely generated pro-p absolute Galois groups over global fields. Journal of Number Theory, 77(1):83–96, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9992379X.

Schertz:1999:ER

[1192] Reinhard Schertz. An elliptic resolvent. Journal of Number Theory, 77(1):97–121, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923691.

Peter:1999:APN

[1193] Manfred Peter. Almost periodicity of the normalized representation numbers associated to positive definite ternary quadratic forms. Journal of Number Theory, 77(1):122–144, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923678.

Harcos:1999:SFS

[1194] Gergely Harcos. On sums of four smooth squares. Journal of Number Theory, 77(1):145–154, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99923806.

Amou:1999:IRV

[1195] Masaaki Amou and Masanori Katsurada. Irrationality results for values of generalized Tschakaloff series. Journal of Number Theory, 77(1): 155–169, July 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99923764.

Anonymous:1999:PAb

[1196] Anonymous. Pages 171–347 (August 1999). Journal of Number Theory, 77(2):??, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Tipp:1999:GFD

[1197] Ulrich Tipp. Green’s functions for Drinfeld modular curves. Journal of Number Theory, 77(2):171–199, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X98923708.

Biro:1999:NPH

[1198] András Biró. Notes on a problem of H. Cohn. Journal of Number Theory, 77(2):200–208, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99923788.

Armon:1999:ARC

[1199] M. V. Armon. Averages of real character sums. Journal of Number Theory, 77(2):209–226, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99923831.

Grant:1999:UTJ

[1200] David Grant. Units from 5-torsion on the Jacobian of y^{2} = x^{5} + 1/4 and the conjectures of Stark and Rubin. Journal of Number Theory, 77(2):227–251, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X98923460.

Byott:1999:IGM

[1201] Nigel P. Byott. Integral Galois module structure of some Lubin– Tate extensions. Journal of Number Theory, 77(2):252–273, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X99923855.

Bombieri:1999:CLC

[1202] Enrico Bombieri and Jeffrey C. Lagarias. Complements to Li’s criterion for the Riemann Hypothesis. Journal of Number Theory, 77(2):274– 287, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99923922.

Anuradha:1999:NPP

[1203] N. Anuradha and S. A. Katre. Number of points on the projective curves aY^{l} = bX^{l}+cZ^{l} and aY^{2l} = bX^{2l}+cZ^{2l} defined over finite fields, l an odd prime. Journal of Number Theory, 77(2):288–313, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9992382X.

Gekeler:1999:GOC

[1204] Ernst-Ulrich Gekeler. Growth order and congruences of coefficients of the Drinfeld discriminant function. Journal of Number Theory, 77(2):314–325, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99923879.

Hegyvari:1999:DHC

[1205] Norbert Hegyvári. On the dimension of the Hilbert cubes. Journal of Number Theory, 77(2):326–330, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99923892.

Jimenez-Urroz:1999:NTZ

[1206] Jorge Jimenez-Urroz. Non-trivial zeroes for quadratic twists of Hasse– Weil L-functions. Journal of Number Theory, 77(2):331–335, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X99923910.

Wang:1999:IPP

[1207] Julie Tzu-Yueh Wang. Integral points of projective spaces omitting hyperplanes over function fields of positive characteristic. Journal of Number Theory, 77(2):336–346, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924009.

Anonymous:1999:AIVd

[1208] Anonymous. Author index for volume 77. Journal of Number Theory, 77(2):347, August 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924502.

Anonymous:1999:PS

[1209] Anonymous. Pages 1–165 (September 1999). Journal of Number Theory, 78(1):??, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Tsuji:1999:SLU

[1210] Takae Tsuji. Semi-local units modulo cyclotomic units. Journal of Number Theory, 78(1):1–26, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99923983.

Bollobas:1999:NSM

[1211] Béla Bollobás and Imre Leader. The number of k-sums modulo k. Journal of Number Theory, 78(1):27–35, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924058.

Girstmair:1999:ACN

[1212] Kurt Girstmair. An algorithm for the construction of a normal basis. Journal of Number Theory, 78(1):36–45, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99923880.

Hsu:1999:DIP

[1213] Chih-Nung Hsu. Diophantine inequalities for polynomial rings. Journal of Number Theory, 78(1):46–61, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99923909.

Schinzel:1999:CNS

[1214] A. Schinzel, J. Urbanowicz, and P. Van Wamelen. Class numbers and short sums of Kronecker symbols. Journal of Number Theory, 78(1):62–84, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9992396X.

Moree:1999:PAP

[1215] Pieter Moree. On primes in arithmetic progression having a prescribed primitive root. Journal of Number Theory, 78(1):85–98, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924095.

Szymiczek:1999:RCG

[1216] Kazimierz Szymiczek. p-ranks of class groups of Witt equivalent number fields. Journal of Number Theory, 78(1):99–106, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X99923971.

Uchida:1999:AIP

[1217] Yoshihisa Uchida. Algebraic independence of the power series defined by blocks of digits. Journal of Number Theory, 78(1):107–118, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99923995.

Bocherer:1999:TOT

[1218] Siegfried Böcherer, Jens Funke, and Rainer Schulze-Pillot. Trace operator and theta series. Journal of Number Theory, 78(1):119–139, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924046.

Jimenez-Urroz:1999:NCE

[1219] Jorge Jiménez-Urroz. A Note on a conjecture of Erdős and Rosenfeld. Journal of Number Theory, 78(1):140–143, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924071.

Reznick:1999:N

[1220] Bruce Reznick. On the nonmonotonicity of (|ℑ(z^{n})|). Journal of Number Theory, 78(1):144–148, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99923946.

Leprevost:1999:CSE

[1221] Franck Leprévost. Sur certaines surfaces elliptiques et courbes elliptiques de Mordell de rang ≥ 1 associées à des discriminants de polynômes cubiques ou quartiques. (French) [On certain elliptic surfaces and elliptic curves of Mordell of rank ≥ 1 associated with discriminants of cubic or quartic polynomials]. Journal of Number Theory, 78(1):149–165, September 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99923934.

Anonymous:1999:PO

[1222] Anonymous. Pages 167–331 (October 1999). Journal of Number Theory, 78(2):??, October 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bockle:1999:PAP

[1223] Gebhard Böckle and Ariane Mézard. The prime-to-adjoint principle and unobstructed Galois deformations in the Borel case. Journal of Number Theory, 78(2):167–203, October 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99923843.

Young:1999:CBE

[1224] Paul Thomas Young. Congruences for Bernoulli, Euler, and Stirling Numbers. Journal of Number Theory, 78(2):204–227, October 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X99924010.

Angles:1999:ERL

[1225] Bruno Anglès. On explicit reciprocity laws for the local Carlitz– Kummer symbols. Journal of Number Theory, 78(2):228–252, October 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924034.

Conrad:1999:RMR

[1226] Brian Conrad and Siman Wong. Remarks on mod-l^{n} representations, l = 3, 5. Journal of Number Theory, 78(2):253–270, October 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X9992406X.

Lan:1999:LF

[1227] Yizhong Lan. A limit formula for ζ(2k + 1). Journal of Number Theory, 78(2):271–286, October 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924083.

Remond:1999:HTC

[1228] Gaël Rémond. Hauteurs thêta et construction de Kodaira. (French) [Theta heights and Kodaira construction]. Journal of Number Theory, 78(2):287–311, October 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924113.

Heuberger:1999:GRN

[1229] Clemens Heuberger and Maohua Le. On the generalized Ramanujan– Nagell equation x^{2} + D = p^{z}. Journal of Number Theory, 78(2):312– 331, October 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924150.

Anonymous:1999:PN

[1230] Anonymous. Pages 1–173 (November 1999). Journal of Number Theory, 79(1):??, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Dou:1999:TCR

[1231] Ze-Li Dou. On a theta correspondence with respect to a quadratic extension. Journal of Number Theory, 79(1):1–18, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924162.

Khare:1999:DRB

[1232] Chandrashekhar Khare. Deformation rings and base change. Journal of Number Theory, 79(1):19–28, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924186.

Schmidt:1999:ONJ

[1233] Ralf Schmidt. On old and new Jacobi forms. Journal of Number Theory, 79(1):29–57, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9992423X.

Chapman:1999:FFC

[1234] Scott T. Chapman, Jeremy Herr, and Natalie Rooney. A factorization formula for class number two. Journal of Number Theory, 79(1):58–66, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924265.

Hsu:1999:BTT

[1235] Chih-Nung Hsu. The Brun–Titchmarsh Theorem in function fields. Journal of Number Theory, 79(1):67–82, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924174.

Kleinbock:1999:BAS

[1236] Dmitry Kleinbock. Badly approximable systems of affine forms. Journal of Number Theory, 79(1):83–102, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924198.

Papadopoulos:1999:CEA

[1237] Ioannis Papadopoulos. Courbes elliptiques ayant même 6-torsion qu’une courbe elliptique donnée. (French) [Elliptical curves having the same 6-torsion as a given elliptical curve]. Journal of Number Theory, 79 (1):103–114, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924101.

Ruzsa:1999:EI

[1238] Imre Z. Ruzsa. Erdős and the integers. Journal of Number Theory, 79(1):115–163, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99923958.

Greither:1999:SMH

[1239] Cornelius Greither, Daniel R. Replogle, Karl Rubin, and Anupam Srivastav. Swan modules and Hilbert–Speiser number fields. Journal of Number Theory, 79(1):164–173, November 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924253.

Anonymous:1999:PD

[1240] Anonymous. Pages 175–345 (December 1999). Journal of Number Theory, 79(2):??, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Yang:1999:NNQ

[1241] Hee-Sun Yang and Soun-Hi Kwon. The non-normal quartic CM-fields and the octic dihedral CM-fields with relative class number two. Journal of Number Theory, 79(2):175–193, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924216.

Hoit:1999:DGS

[1242] Abigail Hoit. The distribution of generalized sum-of-digits functions in residue classes. Journal of Number Theory, 79(2):194–216, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924241.

Ota:1999:SDC

[1243] Kaori Ota. On saturated distinguished chains over a local field. Journal of Number Theory, 79(2):217–248, December 1999. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X9992437X.

Byeon:1999:CNI

[1244] Dongho Byeon. Class numbers and Iwasawa invariants of certain totally real number fields. Journal of Number Theory, 79(2):249–257, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924332.

Angles:1999:OCU

[1245] Bruno Anglès. On the orthogonal of cyclotomic units in positive characteristic. Journal of Number Theory, 79(2):258–283, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924289.

Thakur:1999:DAE

[1246] Dinesh S. Thakur. Diophantine approximation exponents and continued fractions for algebraic power series. Journal of Number Theory, 79(2):284–291, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924137.

Zhai:1999:ACP

[1247] Wenguang Zhai. The additive completion of k-th powers. Journal of Number Theory, 79(2):292–300, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924411.

Hoffstein:1999:AFS

[1248] Jeffrey Hoffstein, Kathy D. Merrill, and Lynne H. Walling. Automorphic forms and sums of squares over function fields. Journal of Number Theory, 79(2):301–329, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9992454X.

Ahlswede:1999:CFP

[1249] Rudolf Ahlswede, Levon H. Khachatrian, and András Sárközy. On the counting function of primitive sets of integers. Journal of Number Theory, 79(2):330–344, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924277.

Anonymous:1999:AIVe

[1250] Anonymous. Author index for volume 79. Journal of Number Theory, 79(2):345, December 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924861.

Escassut:2000:CUU

[1251] Alain Escassut, Labib Haddad, and Robert Vidal. Corrigendum: Urs, Ursim, and Non-Urs for p-adic Functions and Polynomials: Volume 75, Number 1 (1999), pages 133–144. Journal of Number Theory, 80(2):316–317, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924666. See [1146].

Anonymous:2000:CVN

[1252] Anonymous. Corrigendum: Volume 73, number 2 (1998), in the article “Real Polynomials with All Roots on the Unit Circle and Abelian Varieties over Finite Fields,” by Stephen A. DiPippo and Everett W. Howe, pages 426–450. Journal of Number Theory, 83(1): 182, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925654. See [1106].

Anonymous:2000:EVN

[1253] Anonymous. Erratum: Volume 76, number 2 (1999), in the article “The Largest Integer Expressible as a Sum of Reciprocal of Integers,” by Hisashi Yokota, pages 206–216. Journal of Number Theory, 83(1):183–184, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925666. See [1175].

Yamada:2010:NPB

[1254] Tomohiro Yamada. A note on the paper by Bugeaud and Laurent “Minoration effective de la distance p-adique entre puissances de nombres algébriques”. Journal of Number Theory, 130(9):1889–1897, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000983. See [817].

Schaefer:2014:CCG

[1255] Edward F. Schaefer. Corrigendum to “Class groups and Selmer groups” [J. Number Theory 56 (1) (1996) 79–114]. Journal of Number Theory, 135(??):390, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13002370. See [660].